### v1

# Quantum information with quantum-like bits

Abstract
In previous work we have proposed a construction of quantum-like bits that could endow a large, complex classical system, for example of oscillators, with quantum-like function that is not compromised by decoherence. In the present paper we investigate further this platform of quantum-like states. Firstly, we discuss a general protocol on how to construct synchronizing networks that allow for emergent states. We then study how gates can be implemented on those states. This suggests the possibility of quantum-like computing on specially-constructed classical networks. Finally, we define a notion of measurement that allows for non-Kolmogorov interference, a feature that separates our model from a classical probabilistic system. This paper aims to explore the mathematical structure of quantum-like resources, and shows how arbitrary gates can be implemented by manipulating emergent states in those systems.
I Introduction
The present noisy-intermediate scale quantum (NISQ) era of quantum devices is witnessing the successful implementation and test of diverse quantum algorithms on medium-size quantum hardware, from search on unsorted databases, Zhang et al. (2021) to variational quantum algorithms, Cerezo et al. (2021) to simulations of open quantum dynamics. Sun, Shih, and Cheng (2024) However, hardware of at least hundreds of thousands of physical qubits, way beyond the current engineering capabilities, is needed to obtain practical quantum advantage. This involves storing and processing the information volume required by the most challenging and relevant computational problems. Beverland et al. (2022) The critical issue limiting the scalability of quantum resources is the buildup of uncontrolled noise and dissipation, leading to loss of coherence and fidelity. Preskill (2018); Lau et al. (2022); Gill et al. (2022)
In parallel with quantum engineering advancements, ongoing research investigates whether classical computing resources can continue to play a key role in information processing. For example, tensor network simulations have been shown to outperform the most advanced quantum processors in selected algorithms. Tindall et al. (2024) Efficient quantum-inspired protocols have been successfully developed, leveraging advancements in information processing derived from quantum logic. Gharehchopogh (2023)
In the spirit of quantum-inspired computing, we showed in recent work that a large class of synchronizing classical systems can be utilized as a resource for information processing. Scholes (2020, 2023, 2024); Scholes and Amati (2024) The robustness of these systems to noise and stochastic disorder improves with size, following a somehow opposite and favorable trend compared to genuine quantum resources. This well-known behavior is related to the occurrence of classical emergent phenomena and synchronization. Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) Graph theory is the natural mathematical framework used to describe synchronizing many-body systems. Scholes (2020, 2023); Kassabov, Strogatz, and Townsend (2022); Townsend, Stillman, and Strogatz (2020); Rodrigues et al. (2016) In particular, emergent states can be clearly distinguished from a broad band of random states in the spectrum of the adjacency matrices of Erdös-Renyi random graphs and expander families, such as the $k$-regular graphs. Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) In Ref. Scholes, 2024, we mapped emergent states onto a computational basis of quantum-like (QL) bits, which can be easily identified from sharp and isolated peaks in the spectra of regular graphs. This feature implies that QL states are easily distinguishable and measurable even in presence of noise and dissipation, that can usually broaden the spectral lines of the network. In Ref. Scholes and Amati, 2024 we built on these results and proposed a general structure for the resources of multiple QL bits, based on the Cartesian products of one-QL-bit graphs.
Our approach is derived consistently with a QL probabilistic framework proposed by Khrennikov. Khrennikov (2007a, 2019, 2010) This theory shows that genuine quantum systems are not strictly required to process information in a QL way. Instead, to observe QL phenomena it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and their measurement.
In the present paper we extend the foundational work of Ref. Scholes, 2024; Scholes and Amati, 2024; Scholes, 2023 to formalize the notion of QL states and gates. Furthermore, consistently with Khrennikov’s approach, Khrennikov (2010) we list a set of measurement axioms which allows to define a generalized notion of nonclassical interference between observable operators.
The present paper is structured as follows. In ?? we summarize the formalism describing how to construct a single-QL-bit Hilbert space. In ?? we discuss its generalization to higher dimensions, and we show explicitly in ?? how to construct Bell states for two QL bits. In ?? we study the action of a complete set of QL gates, and we provide concrete examples of their action. Finally, in ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach. In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models. Khrennikov (2010)
In conjunction with Refs. Scholes, 2023; Scholes and Amati, 2024; Scholes, 2024, this work provides the mathematical foundation for a rigorous approach to QL information processing from classical synchronizing systems.
II Quantum-like bits
In this section we derive a mapping a from classical network to the Hilbert space of interacting QL bits. While the concept has been already elucidated in Ref. Scholes, 2024, here we lay out a more formal description. Later in the paper we use this formalism to study the notion of gates to perform quantum information processing, and we provide concrete examples.
II.1 Isolated system
Let us consider two classical many-body systems with pairwise interactions between the elementary constituents. Those systems can be mapped onto two graphs $G_{1}$ and $G_{2}$. We assume here that both graphs involve the same number of vertices ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices. For simplicity we focus here on undirected simple graphs, without loops. The adjacency matrix for this class of graphs is symmetric with entries either 00 or $1$ and null diagonal. Additionally, we assume an isotropic structure for all two-body interactions, such that the valence of the graphs is constant and equal to $k$. Diestel (2017) In particular,
$\sum_{i=1}^{N_{G}}[A_{n}]_{ij^{\prime}}=\sum_{j=1}^{N_{G}}[A_{n}]_{i^{\prime}j}=k,\hskip 22.76219pt\forall i^{\prime},j^{\prime}=1,\cdots,{N_{G}},$ (II.1)
where $n=1,2$. We highlight that, as proven in recent work, Scholes (2023); Scholes and Amati (2024) the present formalism is robust even in the presence of disorder in the network, which may violate the condition of exact regularity. Consequently, the theory discussed in the paper is also applicable to quasi-regular resources, provided analytical results on matrix algebra are replaced by their more general numerical analogs. The direct sum of the two adjacency matrices is
$A_{1}\oplus A_{2}=\begin{bmatrix}A_{1}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&A_{2}\end{bmatrix},$ (II.2)
where we labeled by $\mathbb{0}_{N_{G}}\in M_{N_{G}}(\mathbb{N})$ the null matrix on the off-diagonal blocks. As proposed in Ref. Scholes, 2024, to define QL bits we switch on an interaction between the two subgraphs, and define the resource
$\mathcal{R}=\begin{bmatrix}A_{1}&C\\ C^{\top}&A_{2}\end{bmatrix},$ (II.3)
involving a coupling matrix $C$. Note that $C$ does not represent a simple graph by itself, i.e. it is in general non-symmetric, and it can admit diagonal entries mapping corresponding vertices in $A_{1}$ and $A_{2}$. Conversely, ??, being overall symmetric and loop-free, describes a simple undirected graph. We sample $C$ by randomly assigning $1$ to $N_{\mathrm{C}}\leq{N_{G}}^{2}$ of its entries. This matrix is chosen to be regular with valency $l=N_{\mathrm{C}}/{N_{G}}$ [similarly to ??]. ?? represents the fundamental resource used to build an isolated QL bit. Scholes (2024)
To better describe the structure of the graphs that are suitable resources for our QL platform, we show in ?? four instances of ?? for increasing values of $l$, as specified in the captions.
$l=0$
$l=k/4$
$l=k/2$
$l=k$
Figure 1: Graphs associated to the resource ??, for increasing values of the valency $l$ of the coupling matrix $C$, as specified in the captions. $l$ is given here in units of the valency $k$ of the subgraphs $A_{1}$ (light green/light) and $A_{2}$ (dark orange/dark). The other parameters are fixed to ${N_{G}}=50$ and $k=20$.
The case $l=0$ corresponds to uncoupled regular graphs, as in ??. By increasing $l$, we strengthen the coupling between the two subgraphs. For $l=k$, ?? turns into a regular matrix with total valency $2k$, with the same density of both intra- and cross- subgraphs edges.
As discussed in Refs. Scholes, 2023, 2024, resources with structure as in ?? are particularly appealing for information processing. In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states. By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information logic that can be controlled by network manipulation. We will elaborate on this important concept further in the manuscript.
We study the spectral properties of the resources by plotting in ?? the distribution $\rho(\lambda)$ of the eigenvalues of ?? for increasing values of the valency $l$ of $C$.
Figure 2: Main panel: distribution of the eigenvalues $\lambda$ of the QL-resource ??, for increasing values of the order $l$ of the coupling matrix $C$, from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light). Inset: largest and second largest eigenvalues of the spectrum as a function of $l$, with the same color code as in the main panel. We fix here ${N_{G}}=100$ and $k=25$.
Results are shown from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light). The other parameters are fixed to ${N_{G}}=100$ and $k=25$. The emergent states can be identified for each $l$ from the isolated sharp peaks in the spectrum. For $l=0$ we observe a single isolated line associated to the two-fold degenerate eigenvalue $\lambda=k=25$ (red markers). As the valency $l$ increases, this emergent eigenvalue split into two, and the peaks shift to $\lambda_{\pm}=k\pm l$. We track the “ballistic” evolutions of the eigenvectors as a function of the coupling valency $l$ in the inset of ??. Here, the largest and second largest eigenvalues are shown as squared and dotted markers, respectively. While the value of the largest eigenvalue increases linearly for arbitrary $l$, we observe that the second largest eigenvalue plateaus at values close to $\tilde{\lambda}=2\sqrt{k-1}$. This barrier is defined by the radius of the Wigner semicircle distribution for a graph of order $k$, shown as a black dashed line in the figure. McKay (1981); Farkas et al. (2001) This portion of the spectrum is populated by random low-frequency states, whose high density makes them hardly distinguishable. Scholes (2020, 2023, 2024) From the present analysis we infer that, by fixing $l=(k-\tilde{\lambda})/2$, both the distance between the two emergent eigenvalues and between those and the dark states are maximized. As elucidated in Ref. Scholes, 2023, 2024, this analysis offers valuable insight for creating well-resolved and easily distinguishable emergent states for quantum information purposes.
We discuss now how to construct the Hilbert space $\mathcal{H}^{(1)}$ of an isolated QL qubit from coupled regular graphs, with adjacency matrices with structure as in ??. In particular, we define the mapping
$\psi_{+}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\mapsto\ket{+},\hskip 17.07164pt\psi_{-}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ -a_{2}\end{pmatrix}\mapsto\ket{-},$ (II.4)
where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$. The vectors $\psi_{\sigma}$, $\sigma\in\{+,-\}$, are eigenvectors of ?? with eigenvalues $k\pm l$. In the following we show that the mapping in ?? leads to a well-defined Hilbert space. First of all, it is straightforward to show that the $\psi_{\sigma}$’s are mutually orthogonal and normalized:
$\braket{\psi_{\sigma},\psi_{\sigma^{\prime}}}$ $=\sum_{i=1}^{2{N_{G}}}\psi_{\sigma,i}^{*}\psi_{\sigma^{\prime},i}=\frac{1}{2}\left(|a_{1}|^{2}+(2\delta_{\sigma\sigma^{\prime}}-1)|a_{2}|^{2}\right)$
$=\delta_{\sigma\sigma^{\prime}}\mapsto\braket{{\sigma}{\sigma^{\prime}}}{,}$ (II.5)
where $\sigma,\sigma^{\prime}\in\{+,-\}\equiv\mathcal{S}$. Operators acting on ?? are expressed in terms of the matrices in $M_{2{N_{G}}}(\mathbb{R})$
$\Psi_{\sigma\sigma^{\prime}}=\psi_{\sigma}\otimes\psi_{\sigma^{\prime}}^{\dagger}\mapsto\ket{\sigma}\bra{\sigma^{\prime}}.$ | . (II.6)
As it would be expected in an abstract Hilbert space, the maps in ?? combine according to the rule
$\Psi_{\sigma\sigma^{\prime}}\Psi_{\pi\pi^{\prime}}=\Psi_{\sigma\pi^{\prime}}\braket{\psi_{\sigma^{\prime}},\psi_{\pi}},$ (II.7)
and they transform under Hermitian conjugation as $\Psi_{\sigma\sigma^{\prime}}^{\dagger}=\Psi_{\sigma^{\prime}\sigma}$.
In this section we summarized a list of prescriptions for defining a mapping from a bipartite graph to the Hilbert space of a single qubit, and related properties. Scholes and Amati (2024); Scholes (2024) In the next ?? we expand on an idea proposed in Ref. Scholes and Amati, 2024, to represent entanglement and superposition by coupling together multiple graph resources. Formal derivations on how to implement information processing and define a QL probabilistic model in this framework, follow in the later ????, respectively.
II.2 Many-body system
II.2.1 Formalism
Various representations of the quantum states are possible with respect to classical network structures. For a small number of QL bits the representation defined in Ref. Scholes, 2024 is useful. In that work the Bell states, involving two QL bits, are designed to give insight on proof-of-principle examples of entanglement. In Ref. Scholes and Amati, 2024 we propose a formal representation of states applicable to an arbitrary number of QL bits. This is based on the Cartesian product of single-QL-bit graphs. Consistently with that work, we define here as a simple generalization of ?? the resource for a system of ${N_{\mathrm{QL}}}$ bits
$\mathcal{R}=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\mathbb{1}_{2{N_{G}}}\right]\otimes\mathcal{R}^{(q)}\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}\right],$ (II.8)
where each $\mathcal{R}^{(q)}$ has the structure shown in ??. The tensors
$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}},$ (II.9)
are eigenvectors of ??. Here, different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\cdots,\sigma^{(N_{\mathrm{QL}})}\right\}$, while the components of ?? are expressed via the multi-index $\bm{i}=\left\{i^{(1)},\cdots,i^{(N_{\mathrm{QL}})}\right\}$ as
$[\psi_{\bm{\sigma}}]_{\bm{i}}$ $=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}.$ (II.10)
As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\{+,-\}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$. Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$. The action of $\mathcal{R}$ on $\psi_{\bm{\sigma}}$ is then given by
$\mathcal{R}\psi_{\bm{\sigma}}$ $=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\psi_{\sigma^{(p)}}\right]\otimes\left[\mathcal{R}^{(q)}\psi_{\sigma^{(q)}}\right]\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(p^{\prime})}}\right]$
$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.11)
where
$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.12)
???? tell us that the spectrum of the resource $\mathcal{R}$ of ${N_{\mathrm{QL}}}$ QL bits will involve peaks corresponding to sums of the emergent eigenvalues of each $\mathcal{R}^{(q)}$. Scholes and Amati (2024)
In analogy to ?? for ${N_{\mathrm{QL}}}=1$, the states in ?? define an orthonormal basis for $\mathcal{H}$ according to the scalar product defined by
$\braket{\psi_{\bm{\sigma}},\psi_{\bm{\sigma}^{\prime}}}$ $=\frac{1}{2^{N_{\mathrm{QL}}}}\prod_{q=1}^{{N_{\mathrm{QL}}}}\left[|a_{1}^{(q)}|^{2}+(2\delta_{\sigma^{(q)},\sigma^{\prime(q)}}-1)|a_{2}^{(q)}|^{2}\right]$
$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.13)
Similarly, we extend the mapping of operators in ?? to
$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.14)
with components
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}\right]_{\bm{i,j}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}\psi^{*}_{\sigma^{\prime(q)},j^{(q)}}.$ (II.15)
The Hermitian conjugate of ?? is defined such that
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.16)
and we generalize ?? to
$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.17)
The mapping discussed here provides a general description of state vectors, entanglement and operators constructed from the emergent states of an arbitrary number of QL bits. In the next ?? we show explicitly, as an example, how to construct maximally entangled Bell states for ${N_{\mathrm{QL}}}=2$.
II.2.2 Example: Bell states
Here, we study the spectral properties of two-QL-bit resources, and we show how to construct entangled Bell states with them. ?? in this case simplifies to
$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.18)
The spectrum $\rho(\lambda)$ of a resource with structure as in ?? is shown in ?? for the system parameters ${N_{G}}=16$, $k=6$ and $l=1$. To improve the smoothness of all distributions, we average over a number of $N_{\mathrm{samp}}=10^{3}$ realizations of random regular graphs. The black curve corresponds to the direct (dir.) calculation from the eigenvalues of ??. An alternative approach to calculate the spectrum is possible, by observing that each eigenvalue $\lambda_{i}$ of ?? can be expressed as a sum $\lambda_{i}=\lambda^{(1)}_{j}+\lambda^{(2)}_{k}$, where the $\lambda^{(q)}_{m}$’s are eigenvalues of $\mathcal{R}^{(q)}$, $q=1,2$.Scholes and Amati (2024) We can therefore calculate the spectrum of ?? via the convolution
$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.19)
where $\rho^{(q)}(x)$ is the spectrum of a single QL bit, analogously to what shown in ??. The generalization of ?? provides a simple recipe to construct the spectrum of a system of an arbitrary number of QL bits starting from the spectra for ${N_{\mathrm{QL}}}=1$. The distribution calculated from the convolution (conv.) ??, is shown as a green (light) dashed curve in ??, and it agrees satisfactorily with the direct calculation. We mark in red on the picture the isolated peaks corresponding to the emergent eigenvalues of the two-QL-bit systems. These are $\lambda_{++}=2k+2l$ (circle), $\lambda_{+-}=\lambda_{-+}=2k$ (cross) and $\lambda_{--}=2k-2l$ (square). These eigenvalues are associated to the corresponding eigenvectors $\{\psi_{\sigma\sigma^{\prime}}\}_{\sigma,\sigma^{\prime}\in\mathcal{S}}$, defining an orthonormal basis for the Hilbert space of two QL bits.
Figure 3: Distribution of the eigenvalues of ??, for $N_{\mathrm{s}amp}=10^{3}$ random samples of the graphs resources. The parameters of the graph are fixed here to ${N_{G}}=16$, $k=6$, $l=1$. The solid black curve corresponds to the direct sampling of the eigenvalues of ??, while the green dashed curve is obtained from the convolution of two identical one-QL-bit distributions [??].
To obtain the explicit expression of the eigenvectors, it is convenient to label the eigenvectors of each QL bit as
$\psi^{(1)}_{\pm}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\pm}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.20)
where $a_{i}$’s and $b_{i}$’s are defined as in ?? for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively. With ?? we can express
$\psi_{+\pm}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\pm a_{2}\otimes b_{2}\end{pmatrix},$ (II.21a)
$\psi_{-\pm}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}&\mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.21b)
We can finally construct maximally entangled Bell states via the linear combinations
$\Phi^{\mathrm{Bell}}_{\pm}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{++}\pm\psi_{--}\right),$ (II.22a)
$\Psi^{\mathrm{Bell}}_{\pm}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{+-}\pm\psi_{-+}\right),$ (II.22b)
that is,
$\Phi^{\mathrm{Bell}}_{+}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&a_{2}\otimes b_{2}\end{pmatrix},$ (II.23a)
$\Phi^{\mathrm{Bell}}_{-}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix},$ (II.23b)
$\Psi^{\mathrm{Bell}}_{+}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&-a_{2}\otimes b_{2}\end{pmatrix},$ (II.23c)
$\Psi^{\mathrm{Bell}}_{-}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&-a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix}.$ (II.23d)
By applying the mapping from emergent states to the Hilbert space of QL bits, as discussed here, we can introduce the concept of QL gates for information processing. We propose an approach to implement these gates in the next ??.
III Quantum-like gates
In this section, we discuss the fundamental principles of QL information processing. Specifically, we demonstrate how the emergent states of QL resources can be transformed consistently to any arbitrary quantum gate. We then outline a perspectives on how to translate these transformations onto operations on the edges of the graph resources. This is particularly relevant for the future computational and experimental implementation of QL algorithms.
To connect the usual definition of gates acting on an abstract Hilbert space to our QL formalism, it is convenient to introduce a map $U_{\mathrm{cb}}$ transforming the emergent eigenvectors onto the standard factorized computational basis. This map is defined by
$U_{\mathrm{cb}}=[V_{\mathrm{H}}\otimes\mathbb{1}_{N_{G}}]^{\otimes{N_{\mathrm{QL}}}},$ (III.1)
where
$V_{\mathrm{H}}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (III.2)
denotes the standard Hadamard (H) gate. ?? acts on the factorized states as
$U_{\mathrm{cb}}\psi_{\bm{\sigma}}$ $=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\begin{pmatrix}\frac{1}{2}[1+\sigma^{(q)}]1_{N_{G}}\\ \frac{1}{2}[1-\sigma^{(q)}]1_{N_{G}}\end{pmatrix},$ (III.3)
where
$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (III.4)
The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$, where
$U_{\mathrm{cb}}\psi_{+}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{-}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix}.$ (III.5)
We show in ?? how to construct the transformation $U_{\mathrm{cb}}$ in the more general case of non-regular graphs. The QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ is defined by the matrix transformation
$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (III.6)
Let us consider a few examples of ??. The QL maps of the identity operator and the three Pauli gates are, in the limit of exactly regular graphs,
$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.7a)
$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.7b)
$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.7c)
$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.7d)
respectively. As expected, the transformations in ?? act as the standard Pauli gates:
$U_{\mathrm{X}}\psi_{\pm}=\psi_{\mp},\hskip 14.22636ptU_{\mathrm{Y}}\psi_{\pm}=\pm i\psi_{\mp},\hskip 14.22636ptU_{\mathrm{Z}}\psi_{\pm}=\pm\psi_{\pm}.$ (III.8)
As another example, the QL map of the Hadamard gate entangles the states of the factorized basis:
$U_{\mathrm{H}}\psi_{\pm}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\pm}=\frac{1}{\sqrt{2}}(\psi_{+}\pm\psi_{-}).$ (III.9)
[…]
$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}},$ (III.10)
transforming elements of the factorized basis as follows:
$U_{\mathrm{CN}}\psi_{+\pm}=\psi_{+\pm},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{-\pm}=\psi_{-\mp}.$ (III.11)
Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis. We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??. In future work, we aim to design a bijective map $\psi\mapsto\mathcal{R}_{\psi}$ between QL states and resources. This approach enables the mapping of any gate transformation on states, $\psi\mapsto U_{\mathrm{g}}\psi=\psi_{\mathrm{g}}$, onto operations on cluster of edges of the corresponding resources, $\mathcal{R}_{\psi}\mapsto\mathcal{R}_{\psi_{\mathrm{g}}}$. This provides a clear numerical and experimental framework for QL computing.
In the next ?? we shift the discussion towards the benefits provided by a suitable measure theory designed to add “quantumness” to an otherwise classical network. This framework fulfills premises which can be key for a future quantification of QL advantage, as originally proposed in Ref. Khrennikov, 2019.
IV Quantum-like measurements
In this section, we discuss criteria that allows to introduce a notion of QL interference of probabilities in our approach. This highlights that genuine quantum mechanical systems are not strictly required to build QL stochastic models provided the notions of states, observables and measurement are defined according to a given set of prescriptions.
IV.1 Axiomatic definition
In the following we present a list of axioms, originally proposed in Ref. Khrennikov, 2010, to equip our classical network resources with a stochastic theory which mirrors the probability laws of genuine quantum mechanical systems.
Axiom 1.
QL states are represented by normalized vectors in a Hilbert space $\mathcal{H}$.
States can be expressed by linear combinations of elements of the factorized basis ??:
$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (IV.1)
Pure QL states are by definition normalized, and we establish the equivalence class $\beta\psi\sim\psi$ for $\beta\in\mathbb{C}$, $|\beta|=1$.
Axiom 2.
QL observables are represented by self-adjointed operators on $\mathcal{H}$. Their action on QL states, expressed by linear maps, defines a QL measurement.
Observables are expressed in term of linear combinations with structure
$\mathcal{F}$ $=\sum_{{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}}\left(\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}+\gamma^{*}_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right),\hskip 14.22636pt\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\in\mathbb{C},$ (IV.2)
and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$, as defined in ??. The QL measurement of $\mathcal{F}$ on a state $\psi$ is defined by the map $\psi\mapsto\mathcal{F}\psi$.
Axiom 3.
The spectrum of a QL observable is a measurable quantity.
An arbitrary observable $\mathcal{F}$, as defined in ??, is diagonalizable. Its normalized eigenvectors $\{f_{\bm{\sigma}}\}$σ∈S define an orthonormal basis of $\mathcal{H}$, while the corresponding eigenvalues are real and distinct. Proofs of these standard properties for the specific structure of our QL Hilbert space are given in ??.
Axiom 4.
A QL probability measure defined by Born’s rule is assigned to the spectrum of physical observables.
Specifically, let us consider a generic state $\psi\in\mathcal{H}$. We define by
$P_{\psi}\left(\lambda_{\bm{\sigma}}^{\mathcal{F}}\right)=\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}=|\braket{f_{\bm{\sigma}},\psi}|^{2},$ (IV.3)
the QL probability of collapsing onto the subspace spanned by the eigenvector $f_{\bm{\sigma}}$, with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$. ?? is expressed in terms of the projector
$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}\otimes f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (IV.4)
Axiom 5.
QL conditional probabilities are defined via consecutive projections onto the eigenstates of physical observables.
In particular, let us consider two generic QL observable operators $\mathcal{F}$ and $\mathcal{G}$ whose spectra are defined by
$\mathcal{F}f_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}},\hskip 28.45274pt\mathcal{G}g_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}.$ (IV.5)
For a nonzero state $\psi$ we denote by
$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)=\frac{\left\lVert\Pi^{\mathcal{G}}_{\bm{\sigma}^{\prime}}\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}{\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}$ (IV.6)
the QL conditional probabilities associated to the consecutive measures of $\mathcal{F}$ and $\mathcal{G}$ on a state $\psi$.
We show in the next ?? that the QL probability model defined via Axioms 1 to 5 violates by construction Kolmogorovian classical statistics. This feature allows to introduce a notion of QL interference in our approach. Khrennikov (2003, 2019)
IV.2 Quantum probabilities in a QL model
Before introducing a formula for the update of QL probabilities, it is instructive to briefly recall its classical correspondent, to assess analogies and differences between the two. A classical probability space is defined by the triplet $\{\Omega,\Sigma,P^{(\mathrm{cl})}\}$, where $\Omega$ is a space of experimental outcomes, $\Sigma$ the $\sigma$-algebra of its subsets, and $P^{(\mathrm{cl})}$ a probability measure on $\Sigma$.Shiryaev and Wilson (2013) Be $\{\mathcal{A}_{j}\}_{j=1}^{J}$ a partition of $\Omega$, i.e. a collection of $J$ sets defined such that $A_{j}\neq\emptyset\;\forall j$, $\mathcal{A}_{j}\cap A_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ and $\cup_{j=1}^{J}\mathcal{A}_{j}=\Omega$. $P^{(\mathrm{cl})}$ is a positive measure on the partition elements, that is $P^{(\mathrm{cl})}(\mathcal{A}_{j})>0\;\forall j$, and it is normalized: $\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{A}_{j})=1$. For every $\mathcal{B}\in\Sigma$ the classical formula of total probabilities holds: Kolmogoroff (1933)
$P^{(\mathrm{cl})}(\mathcal{B})$ $=P^{(\mathrm{cl})}\left(\mathcal{B}\cap\cup_{j=1}^{J}A_{j}\right)$
$=\sum_{j=1}^{J}P^{(\mathrm{cl})}\left(\mathcal{B}\cap A_{j}\right)=\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})P^{(\mathrm{cl})}(\mathcal{A}_{j}),$ (IV.7)
where we introduced the classical conditional probability
$P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})=\frac{P^{(\mathrm{cl})}(\mathcal{B}\cap\mathcal{A}_{j})}{P^{(\mathrm{cl})}(\mathcal{A}_{j})}.$ (IV.8)
We discuss now the correspondent of ?? for the present QL framework. Let $\mathcal{F}$ and $\mathcal{G}$ be two nondegenerate QL observbles, whose eigenvalue problems are defined as in ??. In this case, the QL conditional probabilities defined in ?? lose their dependence on the initial state $\psi$:
$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)$ $\mapsto|\braket{g_{\bm{\sigma}^{\prime}},f_{\bm{\sigma}}}|^{2}\equiv P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right).$ (IV.9)
Note that the probabilities in ?? are symmetric, differently from their classical correspondents in ??. Observables $\mathcal{F}$ and $\mathcal{G}$ obeying this property for all $\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}$ are called symmetrically conditioned. Khrennikov (2010) This feature is a signature of Bohr’s principle of contextuality, Bohr (1987); Khrennikov and Kozyrev (2005); Khrennikov (2022) expressing the impossibility of strictly separating physical observables and the experimental context under which those are measured. The role of contextuality in a QL probabilitistic framework can be highlighted by rewriting total and conditional probabilities [????, respectively], on an equal footing:
$P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}}|\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}\right)=\left.P_{\psi}(\lambda^{\mathcal{G}}_{\bm{\sigma}})\right|_{\psi=f_{\bm{\sigma}^{\prime}}}.$ (IV.10)
If ?? is fulfilled, Axioms 1 to 5 lead to the QL formula for probability updates Khrennikov (2007b, 2010); Bulinski and Khrennikov (2004)
$P_{\psi}\left(\lambda_{\bm{\sigma}}^{\mathcal{G}}\right)$ $=\sum_{\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime}}^{\mathcal{F}}\right)P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)$
$\quad+\sum_{\begin{subarray}{c}\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}\in\bm{\mathcal{S}}\\ \bm{\sigma}^{\prime}\neq\bm{\sigma}^{\prime\prime}\end{subarray}}\Lambda^{\psi}_{\bm{\sigma},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}}\Big{[}P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime}}^{\mathcal{F}}\right)P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)$
×$\quad\times P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime\prime}}^{\mathcal{F}}\right)P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)\Big{]}^{1/2}.$ (IV.11)
An extended derivation of ?? is provided in ??, where we give the explicit expression of the nonclassical interference coefficient $\Lambda^{\psi}_{\bm{\sigma},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}}$ [??]. Information processing can be designed in a QL way, to enhance or suppress some spectral measurements compared to classical computing protocols. Khrennikov (2021)
As discussed in Ref. Khrennikov, 2021, an important feature of the QL probability formula ?? is that it does not require the calculation of exponentially time-consuming joint probability distributions. Instead, these are necessary for the classical updates described by ??. This observation is connected to the foundation of quantum advantage. We aim to explore this relevant property in future test of QL algorithms.
V Conclusions
In this paper, we expanded upon the general approach derived in Refs. Scholes, 2023, 2024; Scholes and Amati, 2024 to define and manipulate QL states. In ?? we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs. We introduced the Hilbert space for an arbitrary number of QL bits, and we derived explicitly the Bell states in dimension two. We then proposed in ?? a notion of quantum gates applicable to QL states, and we provided concrete examples of their action. Finally, we discussed in ?? a list of measurement axioms which allow to observe QL interference effects in our approach.
Several open questions remain, presenting diverse opportunities for future research.
In the paper we introduced notions of entanglement, interference, and QL probabilities which avoid the expensive calculation of joint distributions. In future work we plan to investigate how each of these features contributes to achieving a “QL advantage” over strictly classical information processing. The ultimate test of the approach proposed in this work will be to implement benchmark algorithms, and evaluate their time efficiency and resource cost. Bharti et al. (2022)
In the last decades, advanced error correction techniques have been developed to minimize loss of information and decoherence in quantum processors. Cai et al. (2023); Campbell (2024) Similarly, we can ask how to prevent and minimize errors in our framework, e.g. in presence of impurities and irregularities in the structure of the graphs. In this regard, we note that a foundational property of synchronizing graphs is that their robustness against uncontrolled noise and dissipation increases with size. Scholes (2023); Scholes and Amati (2024); Anwar et al. (2023) Therefore, we conjecture that error correction is achievable in our approach by simply increasing the number of classical nodes of the graphs. We aim to further explore this concept in future work, to provide bounds of fidelity of gate transformations as a function of the size of the system and strength of external noise.
It can be insightful to directly relate our work to the rigorous approach to quantum non-Markovianity proposed in Refs. Pollock et al., 2018a; White et al., 2020; Berk et al., 2023. In that framework a notion of conditional probabilities is defined, that accounts explicitly for the role of measurement devices on the statistics. An important feature of that formalism is that it provides an operational definition of non-Markovianity which, differently from other strategies, converges to the correct Kolmogorov conditions in the classical limit. Pollock et al. (2018b)
A broad class of computational approaches to open quantum systems and quantum chemistry involves a perspective somehow complementary to our approach. In particular, quantum nonadiabatic systems are often mapped onto quasiclassical models to minimize computational complexity and inexpensively approximate long-time dynamics. Accurate quasiclassical methods have been derived, by requiring that Kolmogorov probabilities [fulfilling ??] obey by construction quantum detailed balance at long times. Amati, Mannouch, and Richardson (2023); Amati, Runeson, and Richardson (2023); Runeson et al. (2022); Amati et al. (2022); Mannouch and Richardson (2023) It can be insightful to assess whether even more accurate quasiclassical approaches can be derived from a QL stochastic theory which approximates ?? beyond the first classical term.
From an experimental standpoint, different options can be explored to implement synchronizing resources for QL information processing. Insulator-to-metal-transition devices and spin-torque oscillators are viable candidates, as they allow for both distributed and parallel computing, while ensuring good robustness to noise and fault tolerance. Raychowdhury et al. (2019) In future work, we aim to further analyze platforms to experimentally implement QL resources, and the properties required by different approaches for optimal network implementation and manipulation.
The environmental implications of the large-scale industrial production of quantum processors are topics of ongoing debate. Ezratty (2023); Ikonen, Salmilehto, and Möttönen (2017); Paler and Basmadjian (2022) For instance, superconducting devices require significant energy due to the cryogenic cooling of the constituent superconducting circuits. Similarly, quantum computers based on arrays of Rydberg atoms demand intensive energy resources. Meyer (2000); Bernien et al. (2017) Implementing energy-efficient classical circuits for QL computing can contribute to sustainable quantum-processing engineering. Additionally, our method leverages classical synchronization, which scales more efficiently with resources compared to its quantum-mechanical counterpart. Aifer, Thingna, and Deffner (2024)
Finally, it can be a fascinating topic for future research to connect QL information processing introduced here to other information theories on human decision-making and psychology derived in the framework of contextual probability models. This idea has been extensively discussed in the literature of QL probabilistic models.Khrennikov (2016); Ozawa and Khrennikov (2020); Scholes (2024); Widdows, Rani, and Pothos (2023) Implementing our formalism on brain-like neuromorphic circuits Pal et al. (2024) can foster theoretical and experimental investigations in this framework.
VI Acknowledgments
This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
Appendix A Transformation of the emergent states to the computational basis
In ?? we discussed how a QL representation of quantum gates can be straightforwardly defined once the emergent states are mapped onto the computational basis. In general, this transformation cannot be expressed in analytic form as soon as the subgraphs defining the resources are not exactly regular. However, we show here that the problem of determining the map can be rewritten as a matrix inversion problem, easily solvable with standard numeric algebra tools.
Firstly, it is convenient to parametrize the generalization of ?? as follows:
$U_{\mathrm{cb}}=\bigotimes_{q=1}^{N_{\mathrm{QL}}}\begin{pmatrix}X^{(q)}&Y^{(q)}\\ W^{(q)}&Z^{(q)}\end{pmatrix},$ (A.1)
where $X^{(q)}\in\mathbb{R}^{{N_{G}}}$ is diagonal, with components $X_{ij}^{(q)}=\delta_{ij}x_{j}^{(q)}$, and similarly for the other three blocks in ??. ?? can be rewritten as the linear problem
$x_{i}^{(q)}a_{1,i}^{(q)}+y_{i}^{(q)}a_{2,i}^{(q)}$ $={N_{G}}^{-1/2},$ (A.2a)
$w_{i}^{(q)}a_{1,i}^{(q)}+z_{i}^{(q)}a_{2,i}^{(q)}$ $=0,$ (A.2b)
$x_{i}^{(q)}a_{1,i}^{(q)}-y_{i}^{(q)}a_{2,i}^{(q)}$ $=0,$ (A.2c)
$w_{i}^{(q)}a_{1,i}^{(q)}-z_{i}^{(q)}a_{2,i}^{(q)}$ $={N_{G}}^{-1/2},$ (A.2d)
where $q=1,\cdots,{N_{\mathrm{QL}}}$, $i=1,\cdots,{N_{G}}$, and where we made use of the generalization of ?? to arbitrary values of ${N_{\mathrm{QL}}}$:
$\psi_{\pm}^{(q)}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}^{(q)}\\ a_{2}^{(q)}\end{pmatrix}.$ (A.3)
?? is expressed in matrix form as
$\begin{pmatrix}D_{1}^{(q)}&D_{2}^{(q)}&\mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}&D_{1}^{(q)}&D_{2}^{(q)}\\ D_{1}^{(q)}&-D_{2}^{(q)}&\mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}&D_{1}^{(q)}&-D_{2}^{(q)}\end{pmatrix}\begin{pmatrix}x^{(q)}\\ y^{(q)}\\ w^{(q)}\\ z^{(q)}\end{pmatrix}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\\ 0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (A.4)
where $\left[D_{n}^{(q)}\right]_{ij}=\delta_{ij}a_{n,j}^{(q)}$, $n=1,2$. ?? can be inverted to determine a solution for $\left(x^{(q)},y^{(q)},w^{(q)},z^{(q)}\right)$ $\forall q=1,\cdots{N_{\mathrm{QL}}}$, corresponding to a unique expression for $U_{\mathrm{cb}}$.
Appendix B Derivation of the quantum-like formula of total probabilities
In this appendix we derive ??, describing the update of probabilities for two QL observables $\mathcal{F}$ and $\mathcal{G}$, with eigenvalue problems defined as in ??.
Any state vector $\psi\in\mathcal{H}$ can be expanded as
$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}G^{\psi}_{\bm{\sigma}}g_{\bm{\sigma}}=\sum_{\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}F^{\psi}_{\bm{\sigma}^{\prime}}f_{\bm{\sigma}^{\prime}},$ (B.1)
where $F^{\psi}_{\bm{\sigma}^{\prime}}=\braket{f_{\bm{\sigma}^{\prime}},\psi}$ and $G^{\psi}_{\bm{\sigma}}=\braket{g_{\bm{\sigma}},\psi}$. We can connect the probabilities of the two observables by expressing
$\psi$ $=\sum_{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}F^{\psi}_{\bm{\sigma}^{\prime}}u_{\bm{\sigma}^{\prime},\bm{\sigma}}g_{\bm{\sigma}},$ (B.2a)
$u_{\bm{\sigma}^{\prime},\bm{\sigma}}$ $=\braket{g_{\bm{\sigma}},f_{\bm{\sigma}^{\prime}}}.$ (B.2b)
which leads to the following mapping between the coefficients of $\psi$ in the two bases:
$G^{\psi}_{\bm{\sigma}}=\sum_{\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}F^{\psi}_{\bm{\sigma}^{\prime}}u_{\bm{\sigma}^{\prime},\bm{\sigma}}.$ (B.3)
To relate ?? to QL probabilities, we define
$F^{\psi}_{\bm{\sigma}^{\prime}}=\,\mathrm{e}^{i\gamma_{\bm{\sigma}^{\prime}}^{\psi}}\sqrt{P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime}}^{\mathcal{F}}\right)},\hskip 11.38109ptu_{\bm{\sigma}^{\prime},\bm{\sigma}}=\,\mathrm{e}^{i\Gamma_{\bm{\sigma}^{\prime},\bm{\sigma}}}\sqrt{P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)},$ (B.4)
where $\gamma^{\psi}_{\bm{\sigma}^{\prime}},\Gamma_{\bm{\sigma}^{\prime},\bm{\sigma}}\in[0,2\pi)$. By expanding the squared modulus of ??, we finally obtain ??. This involves the phase factor
$\Lambda^{\psi}_{\bm{\sigma},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}}=\,\mathrm{e}^{i\left(\gamma^{\psi}_{\bm{\sigma}^{\prime}}-\gamma^{\psi}_{\bm{\sigma}^{\prime\prime}}\right)}\,\mathrm{e}^{i\left(\Gamma_{\bm{\sigma}^{\prime},\bm{\sigma}}-\Gamma_{\bm{\sigma}^{\prime\prime},\bm{\sigma}}\right)},$ (B.5)
which, being any other term of ?? real, can be replaced by its real part:
$\Lambda^{\psi}_{\bm{\sigma},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}}\mapsto\cos\left[(\gamma^{\psi}_{\bm{\sigma}^{\prime}}-\gamma^{\psi}_{\bm{\sigma}^{\prime\prime}})(\Gamma_{\bm{\sigma}^{\prime},\bm{\sigma}}-\Gamma_{\bm{\sigma}^{\prime\prime},\bm{\sigma}})\right].$ (B.6)
Appendix C Spectral properties of QL observables
In this appendix we study the spectral properties of QL observables [??], that are Hermitian operators according to the definition ??. In particular, we show that (i) their eigenvalues are real, (ii) eigenvectors corresponding to different eigenvalues are orthogonal, and (iii) there exists an orthonormal basis of $\mathcal{H}$ consisting of eigenvectors. We refer here to the eigenvalues and eigenvectors of $\mathcal{F}$ as $\lambda_{\bm{\sigma}}$’s and $f_{\bm{\sigma}}$’s, respectively.
  1. (i)
    For an arbitrary $f_{\bm{\sigma}}$
    $\braket{f_{\bm{\sigma}},\mathcal{F}f_{\bm{\sigma}}}=\lambda_{\bm{\sigma}}\left\lVert f_{\bm{\sigma}}\right\rVert^{2}=\braket{\mathcal{F}f_{\bm{\sigma}},f_{\bm{\sigma}}}=\lambda_{\bm{\sigma}}^{*}\left\lVert f_{\bm{\sigma}}\right\rVert^{2},$
    which implies that $\lambda_{\bm{\sigma}}\in\mathbb{R}$.
  2. (ii)
    Suppose that $\lambda_{\bm{\sigma}}\neq\lambda_{\bm{\sigma}^{\prime}}$ for $\bm{\sigma}\neq\bm{\sigma}^{\prime}$. By multiplying
    $\mathcal{F}f_{\bm{\sigma}}=\lambda_{\bm{\sigma}}f_{\bm{\sigma}},$ (C.1)
    by $f_{\bm{\sigma}^{\prime}}$, we obtain
    $\langle f_{\bm{\sigma}^{\prime}},\mathcal{F}f_{\bm{\sigma}}\rangle$ $=\lambda_{\bm{\sigma}}\langle f_{\bm{\sigma}^{\prime}},f_{\bm{\sigma}}\rangle$
[…]
    which implies that $\langle f_{\bm{\sigma}^{\prime}},f_{\bm{\sigma}}\rangle=0$.
  3. (iii)
    Let us order the eigenvalues and eigenvectors of $\mathcal{F}$ in increasing order, via the map
    $\bm{\sigma}\mapsto n=\sum_{q=1}^{N_{\mathrm{QL}}}2^{{N_{\mathrm{QL}}}-q}\left(\frac{1-\sigma^{(q)}}{2}\right)+1.$ (C.3)
    Let us define by $V_{1}$ the set of vectors orthogonal to the eigenvector $f_{1}$ of $\mathcal{F}$ with eigenvalue $\lambda_{1}$. Note that $\mathcal{F}$ maps $V_{1}$ onto itself, i.e., $\mathcal{F}\psi\in V_{1}$ for $\psi\in V_{1}$. In fact,
    $\braket{f_{1},\mathcal{F}\psi}=\braket{\mathcal{F}f_{1},\psi}=\lambda_{1}\braket{f_{1},\psi}=0.$ (C.4)
    The linear operator $L(\phi)=\mathcal{F}\phi$, when restricted to $V_{1}$, is also Hermitian, and it admits an eigenvalue $\lambda_{2}$ with eigenvector $f_{2}\in V_{1}$. By construction, $f_{2}$ is orthogonal to $f_{1}$. We then define by $V_{2}$ the orthogonal complement to $\mathrm{span}\{f_{1},f_{2}\}$. Again, $\mathcal{F}$ maps $V_{2}$ onto itself. By proceeding this way, we find a sequence $\{\lambda_{n},f_{n},V_{n}\}_{n}$, where the subspace $V_{n}$ is orthogonal to $\mathrm{span}\{f_{1},\cdots,f_{n}\}$. The sequence ends at the step of order $2^{N_{\mathrm{QL}}}$, given that $\dim(V_{n})=2^{N_{\mathrm{QL}}}-n$. This process allows us to define a complete set of mutually orthogonal eigenvectors.

### v4

@@
 In the spirit of quantum-inspired computing, we showed in recent work that a large class of synchronizing classical systems can be utilized as a resource for information processing.
-Scholes (2020, 2023, 2024); Scholes and Amati (2024) The robustness of these systems to noise and stochastic disorder improves with size, following a somehow opposite and favorable trend compared to genuine quantum resources.
-This well-known behavior is related to the occurrence of classical emergent phenomena and synchronization.
-Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) Graph theory is the natural mathematical framework used to describe synchronizing many-body systems.
-Scholes (2020, 2023); Kassabov, Strogatz, and Townsend (2022); Townsend, Stillman, and Strogatz (2020); Rodrigues et al.
-(2016) In particular, emergent states can be clearly distinguished from a broad band of random states in the spectrum of the adjacency matrices of Erdös-Renyi random graphs and expander families, such as the $k$-regular graphs.
-Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) In Ref.
-Scholes, 2024, we mapped emergent states onto a computational basis of quantum-like (QL) bits, which can be easily identified from sharp and isolated peaks in the spectra of regular graphs.
-This feature implies that QL states are easily distinguishable and measurable even in presence of noise and dissipation, that can usually broaden the spectral lines of the network.
+Scholes (2020, 2023, 2024); Scholes and Amati (2024) We focused, in particular, on many body systems whose correlations obey the graph structure shown in panel a) of ??.
+Here, two densely connected regions (in blue) of a network are mutually interacting via sparse links (in red).
+A special feature of this class of networks is that they produce pairs of emergent states, obtained by the direct sum of the emergent states of the blue subgraphs, with identical and opposite phases.
+The key requirement of the blue subgraphs, then, is that they exhibit an emergent state—a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs.
+Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) In our framework, emergent states can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??).
+Their robustness against noise and stochastic disorder increases with system size, exhibiting a favorable trend in contrast to the typical behavior of genuine quantum resources.
+Scholes (2023, 2020) This intriguing feature is connected to the occurrence of classical synchronization phenomena.
+Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) We showed in Ref.
+Scholes, 2024 that the emergent states of connected expander subgraphs are isomorphic to the Hilbert space of a qubit.
+This led us to introducing the notion of quantum-like (QL) bits within the context of classically-synchronizing networks.
 In Ref.
-Scholes and Amati, 2024 we built on these results and proposed a general structure for the resources of multiple QL bits, based on the Cartesian products of one-QL-bit graphs.
-Our approach is derived consistently with a QL probabilistic framework proposed by Khrennikov.
-Khrennikov (2007a, 2019, 2010) This theory shows that genuine quantum systems are not strictly required to process information in a QL way.
+Scholes and Amati, 2024 we built on these results and proposed a general framework to construct resources for multiple QL bits.
+In particular, we showed that the Cartesian product ($\square$) of ${N_{\mathrm{QL}}}$ bipartite graphs obeys the same dimensionality of a register of ${N_{\mathrm{QL}}}$ qubits [see panel b) of ??].
+In the present paper, we build on the foundational work of Ref.
+Scholes, 2024; Scholes and Amati, 2024; Scholes, 2023 to formalize the concept of QL states and gates.
+We present a computational framework for implementing information processing within the QL formalism through appropriate manipulations of the emergent states of network resources.
+As depicted in panel c) of ??, we show that an arbitrary gate transformation $V_{\mathrm{g}}$ can be expressed both as graph transformation, and as a map of the emergent states of the graph.
+Finally, we propose a measurement approach upon information processing that aligns with the QL probabilistic framework developed by Khrennikov Khrennikov (2007a, 2019, 2010).
+This theory shows that genuine quantum systems are not strictly required to process information in a QL way.
 Instead, to observe QL phenomena it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and their measurement.
-In the present paper we extend the foundational work of Ref.
-Scholes, 2024; Scholes and Amati, 2024; Scholes, 2023 to formalize the notion of QL states and gates.
-Furthermore, consistently with Khrennikov’s approach, Khrennikov (2010) we list a set of measurement axioms which allows to define a generalized notion of nonclassical interference between observable operators.
+Consistently with Khrennikov’s approach, Khrennikov (2010) we list a set of measurement axioms which allows to define a generalized notion of nonclassical interference between observable operators.
 The present paper is structured as follows.
@@
 we study the action of a complete set of QL gates, and we provide concrete examples of their action.
-Finally, in ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach.
+In ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach.
 In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models.
-Khrennikov (2010)
+Khrennikov (2010) Finally, in ??, we outline the complete workflow for QL information processing and discuss its efficiency and accuracy.
 In conjunction with Refs.
 Scholes, 2023; Scholes and Amati, 2024; Scholes, 2024, this work provides the mathematical foundation for a rigorous approach to QL information processing from classical synchronizing systems.
+Figure 1: Illustration of the main concepts presented in this paper and in prior work by the authors.
+Scholes (2024); Scholes and Amati (2024) a) Many body systems consisting of networks of synchronizing dipoles can easily exhibit classical synchronization phenomena.
+By structuring two-body correlations into two densely connected regions (blue) weakly interacting through sparse links (red), synchronization gives rise to pairs of emergent states.
+These states are well-distinguishable from a large band of random states at lower energies (dashed line in the sketch of the spectrum).
+In Ref.
+Scholes, 2024 we proposed a map from emergent states to the computational basis of a qubit.
+b) In Ref.
+Scholes and Amati, 2024, we demonstrated that the emergent states of the Cartesian product of a set of synchronizing graphs scale with the same dimensionality of a qubit system.
+In this paper, we expand over these concepts by providing further mathematical details on the formalism for this many-body problem.
+c) Additionally, we elaborate on the concept of information processing within the QL framework.
+We show that operations on graphs, implemented through suitable transformations of their edges, are equivalent to gate transformations $V_{\mathrm{g}}$ applied directly to the emergent states.
 II Quantum-like bits
-In this section we derive a mapping a from classical network to the Hilbert space of interacting QL bits.
+In this section we derive a mapping from a classical network to the Hilbert space of interacting QL bits.
 While the concept has been already elucidated in Ref.
@@
 Those systems can be mapped onto two graphs $G_{1}$ and $G_{2}$.
-We assume here that both graphs involve the same number of vertices ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices.
+We assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices.
 For simplicity we focus here on undirected simple graphs, without loops.
@@
 Note that $C$ does not represent a simple graph by itself, i.e.
-it is in general non-symmetric, and it can admit diagonal entries mapping corresponding vertices in $A_{1}$ and $A_{2}$.
+it is in general non-symmetric, and it can admit diagonal entries connecting corresponding vertices in $A_{1}$ and $A_{2}$.
 Conversely, ??, being overall symmetric and loop-free, describes a simple undirected graph.
@@
 In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states.
-By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information logic that can be controlled by network manipulation.
+By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information processing that can be controlled by network manipulation.
 We will elaborate on this important concept further in the manuscript.
@@
 In particular, we define the mapping
-$\psi_{+}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\mapsto\ket{+},\hskip 17.07164pt\psi_{-}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ -a_{2}\end{pmatrix}\mapsto\ket{-},$ (II.4)
+$\psi_{\downarrow}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ -a_{2}\end{pmatrix}\mapsto\ket{\downarrow},\hskip 17.07164pt\psi_{\uparrow}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\mapsto\ket{\uparrow}$ (II.4)
 where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$.
-The vectors $\psi_{\sigma}$, $\sigma\in\{+,-\}$, are eigenvectors of ??
-with eigenvalues $k\pm l$.
+The vectors $\psi_{\sigma}$, $\sigma\in\{\downarrow,\uparrow\}\equiv\mathcal{S}$, are eigenvectors of ??
+with eigenvalues $\{k-l,k+l\}$, respectively.
+The current arrow notation for the computational basis helps prevent any ambiguity with additional numeric subscripts and superscripts, which are used here and in the following sections for different purposes.
+In prior work we used the intention of the QL bit construction to represent explicitly the basis of the the subgraphs, so that the emergent state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and similarly $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2024) ??
+describes another basis choice, equally valid.
+The present approach establishes a direct connection between the computational basis and the emergent states of the composite graph described by ??.
 In the following we show that the mapping in ??
@@
 $=\delta_{\sigma\sigma^{\prime}}\mapsto\braket{{\sigma}{\sigma^{\prime}}}{,}$ (II.5)
-where $\sigma,\sigma^{\prime}\in\{+,-\}\equiv\mathcal{S}$. Operators acting on ?? are expressed in terms of the matrices in $M_{2{N_{G}}}(\mathbb{R})$
+where $\sigma,\sigma^{\prime}\in\mathcal{S}$. Operators acting on ?? are expressed in terms of the matrices in $M_{2{N_{G}}}(\mathbb{R})$
 $\Psi_{\sigma\sigma^{\prime}}=\psi_{\sigma}\otimes\psi_{\sigma^{\prime}}^{\dagger}\mapsto\ket{\sigma}\bra{\sigma^{\prime}}.$ | . (II.6)
@@
 $\mathcal{R}=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\mathbb{1}_{2{N_{G}}}\right]\otimes\mathcal{R}^{(q)}\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}\right],$ (II.8)
-where each $\mathcal{R}^{(q)}$ has the structure shown in ??. The tensors
-$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}},$ (II.9)
+where each $\mathcal{R}^{(q)}$ has the structure shown in ??. ?? is the adjacency matrix of the Cartesian product graph Scholes and Amati (2024)
+[…]
+Here, $\mathcal{G}^{(q)}$ is the single-QL-bit graph with adjacency matrix $\mathcal{R}^{(q)}$. The tensors
+$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}},$ (II.10)
 are eigenvectors of ??.
@@
 are expressed via the multi-index $\bm{i}=\left\{i^{(1)},\cdots,i^{(N_{\mathrm{QL}})}\right\}$ as
-$[\psi_{\bm{\sigma}}]_{\bm{i}}$ $=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}.$ (II.10)
-As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\{+,-\}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$.
+$[\psi_{\bm{\sigma}}]_{\bm{i}}$ $=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}.$ (II.11)
+As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\{0,1\}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$.
 Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$.
@@
 $\mathcal{R}\psi_{\bm{\sigma}}$ $=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\psi_{\sigma^{(p)}}\right]\otimes\left[\mathcal{R}^{(q)}\psi_{\sigma^{(q)}}\right]\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(p^{\prime})}}\right]$
-$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.11)
+$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.12)
 where
-$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.12)
+$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.13)
 ????
@@
 $\braket{\psi_{\bm{\sigma}},\psi_{\bm{\sigma}^{\prime}}}$ $=\frac{1}{2^{N_{\mathrm{QL}}}}\prod_{q=1}^{{N_{\mathrm{QL}}}}\left[|a_{1}^{(q)}|^{2}+(2\delta_{\sigma^{(q)},\sigma^{\prime(q)}}-1)|a_{2}^{(q)}|^{2}\right]$
-$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.13)
+$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.14)
 Similarly, we extend the mapping of operators in ?? to
-$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.14)
+$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.15)
 with components
-$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}\right]_{\bm{i,j}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}\psi^{*}_{\sigma^{\prime(q)},j^{(q)}}.$ (II.15)
+$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}\right]_{\bm{i,j}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}\psi^{*}_{\sigma^{\prime(q)},j^{(q)}}.$ (II.16)
 The Hermitian conjugate of ?? is defined such that
-$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.16)
+$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.17)
 and we generalize ?? to
-$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.17)
+$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.18)
 The mapping discussed here provides a general description of state vectors, entanglement and operators constructed from the emergent states of an arbitrary number of QL bits.
@@
 Here, we study the spectral properties of two-QL-bit resources, and we show how to construct entangled Bell states with them. ?? in this case simplifies to
-$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.18)
+$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.19)
 The spectrum $\rho(\lambda)$ of a resource with structure as in ??
@@
 via the convolution
-$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.19)
-where $\rho^{(q)}(x)$ is the spectrum of a single QL bit, analogously to what shown in ??.
+$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.20)
+where $\rho^{(q)}(x)$ is the spectrum of a single QL bit, shown in ??.
 The generalization of ??
@@
 We mark in red on the picture the isolated peaks corresponding to the emergent eigenvalues of the two-QL-bit systems.
-These are $\lambda_{++}=2k+2l$ (circle), $\lambda_{+-}=\lambda_{-+}=2k$ (cross) and $\lambda_{--}=2k-2l$ (square).
+These are $\lambda_{11}=2k+2l$ (circle), $\lambda_{10}=\lambda_{01}=2k$ (cross) and $\lambda_{00}=2k-2l$ (square).
 These eigenvalues are associated to the corresponding eigenvectors $\{\psi_{\sigma\sigma^{\prime}}\}_{\sigma,\sigma^{\prime}\in\mathcal{S}}$, defining an orthonormal basis for the Hilbert space of two QL bits.
@@
 To obtain the explicit expression of the eigenvectors, it is convenient to label the eigenvectors of each QL bit as
-$\psi^{(1)}_{\pm}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\pm}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.20)
-where $a_{i}$’s and $b_{i}$’s are defined as in ?? for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively. With ?? we can express
-$\psi_{+\pm}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\pm a_{2}\otimes b_{2}\end{pmatrix},$ (II.21a)
-$\psi_{-\pm}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}&\mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.21b)
+$\psi^{(1)}_{[\uparrow\downarrow]}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{[\uparrow\downarrow]}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.21)
+where $a_{i}$’s and $b_{i}$’s are defined as in ??
+for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively.
+The square brackets indicate that the sign on the right-hand side of each expression (+ or -) is chosen depending on the direction of the arrow on the left-hand side ($\uparrow$ or $\downarrow$, respectively).
+With ??
+we can express
+$\psi_{\uparrow,[\uparrow\downarrow]}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\pm a_{2}\otimes b_{2}\end{pmatrix},$ (II.22a)
+$\psi_{\downarrow,[\uparrow\downarrow]}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}&\mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.22b)
 We can finally construct maximally entangled Bell states via the linear combinations
-$\Phi^{\mathrm{Bell}}_{\pm}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{++}\pm\psi_{--}\right),$ (II.22a)
-$\Psi^{\mathrm{Bell}}_{\pm}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{+-}\pm\psi_{-+}\right),$ (II.22b)
+$\Phi^{\mathrm{Bell}}_{[\uparrow\downarrow]}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.23a)
+$\Psi^{\mathrm{Bell}}_{[\uparrow\downarrow]}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.23b)
 that is,
-$\Phi^{\mathrm{Bell}}_{+}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&a_{2}\otimes b_{2}\end{pmatrix},$ (II.23a)
-$\Phi^{\mathrm{Bell}}_{-}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix},$ (II.23b)
-$\Psi^{\mathrm{Bell}}_{+}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&-a_{2}\otimes b_{2}\end{pmatrix},$ (II.23c)
-$\Psi^{\mathrm{Bell}}_{-}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&-a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix}.$ (II.23d)
+$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&a_{2}\otimes b_{2}\end{pmatrix},$ (II.24a)
+$\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix},$ (II.24b)
+$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&-a_{2}\otimes b_{2}\end{pmatrix},$ (II.24c)
+$\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&-a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix}.$ (II.24d)
 By applying the mapping from emergent states to the Hilbert space of QL bits, as discussed here, we can introduce the concept of QL gates for information processing.
@@
 Specifically, we demonstrate how the emergent states of QL resources can be transformed consistently to any arbitrary quantum gate.
-We then outline a perspectives on how to translate these transformations onto operations on the edges of the graph resources.
+We then study how to translate these transformations into operations on the edges of the graph resources.
 This is particularly relevant for the future computational and experimental implementation of QL algorithms.
@@
 $V_{\mathrm{H}}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (III.2)
-denotes the standard Hadamard (H) gate. ?? acts on the factorized states as
-$U_{\mathrm{cb}}\psi_{\bm{\sigma}}$ $=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\begin{pmatrix}\frac{1}{2}[1+\sigma^{(q)}]1_{N_{G}}\\ \frac{1}{2}[1-\sigma^{(q)}]1_{N_{G}}\end{pmatrix},$ (III.3)
+denotes the standard Hadamard (H) gate. The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$. Here,
+$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (III.3)
 where
 $x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (III.4)
-The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$, where
-$U_{\mathrm{cb}}\psi_{+}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{-}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix}.$ (III.5)
 We show in ??
@@
 The QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ is defined by the matrix transformation
-$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (III.6)
+$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (III.5)
 Let us consider a few examples of ??. The QL maps of the identity operator and the three Pauli gates are, in the limit of exactly regular graphs,
-$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.7a)
-$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.7b)
-$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.7c)
-$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.7d)
+$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.6a)
+$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.6b)
+$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.6c)
+$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.6d)
 respectively. As expected, the transformations in ?? act as the standard Pauli gates:
-$U_{\mathrm{X}}\psi_{\pm}=\psi_{\mp},\hskip 14.22636ptU_{\mathrm{Y}}\psi_{\pm}=\pm i\psi_{\mp},\hskip 14.22636ptU_{\mathrm{Z}}\psi_{\pm}=\pm\psi_{\pm}.$ (III.8)
+[…]
 As another example, the QL map of the Hadamard gate entangles the states of the factorized basis:
-$U_{\mathrm{H}}\psi_{\pm}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\pm}=\frac{1}{\sqrt{2}}(\psi_{+}\pm\psi_{-}).$ (III.9)
+$U_{\mathrm{H}}\psi_{[\uparrow\downarrow]}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{[\uparrow\downarrow]}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (III.8)
 […]
-$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}},$ (III.10)
+$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}},$ (III.9)
 transforming elements of the factorized basis as follows:
-$U_{\mathrm{CN}}\psi_{+\pm}=\psi_{+\pm},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{-\pm}=\psi_{-\mp}.$ (III.11)
+$U_{\mathrm{CN}}\psi_{\uparrow[\uparrow,\downarrow]}=\psi_{\uparrow[\uparrow,\downarrow]},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow[\uparrow,\downarrow]}=\psi_{\downarrow[\downarrow,\uparrow]}.$ (III.10)
+Note that, since a universal set of gates can be constructed by combining one-QL-bit gates and the CNOT, we can limit our analysis to this set of operations. Barenco et al. (1995)
 Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis.
 We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??.
-In future work, we aim to design a bijective map $\psi\mapsto\mathcal{R}_{\psi}$ between QL states and resources.
-This approach enables the mapping of any gate transformation on states, $\psi\mapsto U_{\mathrm{g}}\psi=\psi_{\mathrm{g}}$, onto operations on cluster of edges of the corresponding resources, $\mathcal{R}_{\psi}\mapsto\mathcal{R}_{\psi_{\mathrm{g}}}$.
-This provides a clear numerical and experimental framework for QL computing.
+From an operational standpoint, it is however preferable to translate these state operations onto transformations applicable directed to the graph resources.
+This feature is key to translate our gate formalism into practical algorithms for classical computing resources and for experimental platforms of synchronizing electrical circuits.
+Ochs, Michaelis, and Roggendorf (2019) In this regard, we observe that the network transformation
+$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}$ (III.11)
+transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??. This can be inferred from the spectral decomposition of ??, given by
+$\mathcal{R}=\sum_{i=1}^{\tilde{N}}\lambda_{i}\psi_{i}\otimes\psi_{i}^{\dagger}.$ (III.12)
+Here, $\tilde{N}=(2{N_{G}})^{N_{\mathrm{QL}}}$ denotes the total number of eigenstates of the resource $\mathcal{R}$, including both the emergent and the random states.
 In the next ??
-we shift the discussion towards the benefits provided by a suitable measure theory designed to add “quantumness” to an otherwise classical network.
+we shift the discussion towards the benefits provided by a suitable measure theory designed to add “quantumness” to classical synchronizing networks.
 This framework fulfills premises which can be key for a future quantification of QL advantage, as originally proposed in Ref.
@@
 IV Quantum-like measurements
-In this section, we discuss criteria that allows to introduce a notion of QL interference of probabilities in our approach.
-This highlights that genuine quantum mechanical systems are not strictly required to build QL stochastic models provided the notions of states, observables and measurement are defined according to a given set of prescriptions.
+In this section, we discuss criteria that allow to introduce a notion of QL interference of probabilities in our approach.
+This analysis highlights that genuine quantum mechanical systems are not strictly required to build QL stochastic models provided the notions of states, observables and measurement are defined according to a given set of prescriptions.
 IV.1 Axiomatic definition
@@
 QL states are represented by normalized vectors in a Hilbert space $\mathcal{H}$.
-States can be expressed by linear combinations of elements of the factorized basis ??:
-$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (IV.1)
+States can be expressed by linear combinations of elements of the factorized basis in ??:
+$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}}.$ (IV.1)
 Pure QL states are by definition normalized, and we establish the equivalence class $\beta\psi\sim\psi$ for $\beta\in\mathbb{C}$, $|\beta|=1$.
 Axiom 2.
-QL observables are represented by self-adjointed operators on $\mathcal{H}$. Their action on QL states, expressed by linear maps, defines a QL measurement.
-Observables are expressed in term of linear combinations with structure
+QL observables are represented by self-adjointed operators on $\mathcal{H}$. Their action on QL states, expressed by a linear map, defines a QL measurement.
+Observables are defined in terms of linear combinations with structure
 $\mathcal{F}$ $=\sum_{{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}}\left(\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}+\gamma^{*}_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right),\hskip 14.22636pt\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\in\mathbb{C},$ (IV.2)
-and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$, as defined in ??.
-The QL measurement of $\mathcal{F}$ on a state $\psi$ is defined by the map $\psi\mapsto\mathcal{F}\psi$.
+and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$ [??]. The QL measurement of $\mathcal{F}$ on a state $\psi$ is defined by the map $\psi\mapsto\mathcal{F}\psi$.
 Axiom 3.
@@
 $P_{\psi}\left(\lambda_{\bm{\sigma}}^{\mathcal{F}}\right)=\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}=|\braket{f_{\bm{\sigma}},\psi}|^{2},$ (IV.3)
-the QL probability of collapsing onto the subspace spanned by the eigenvector $f_{\bm{\sigma}}$, with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$.
+the QL probability of collapsing onto the subspace spanned by the eigenvector $f_{\bm{\sigma}}$ with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$.
 ??
@@
 $P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)=\frac{\left\lVert\Pi^{\mathcal{G}}_{\bm{\sigma}^{\prime}}\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}{\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}$ (IV.6)
-the QL conditional probabilities associated to the consecutive measures of $\mathcal{F}$ and $\mathcal{G}$ on a state $\psi$.
+the QL conditional probabilities associated to the consecutive measures of $\mathcal{F}$ and $\mathcal{G}$ on $\psi$.
 We show in the next ??
@@
 Khrennikov (2021)
-As discussed in Ref.
-Khrennikov, 2021, an important feature of the QL probability formula ??
-is that it does not require the calculation of exponentially time-consuming joint probability distributions.
-Instead, these are necessary for the classical updates described by ??.
-This observation is connected to the foundation of quantum advantage.
-We aim to explore this relevant property in future test of QL algorithms.
-V Conclusions
+As noted in Ref.
+Khrennikov, 2021, a significant feature of the QL probability formula ??
+is that it bypasses the need for calculating joint probability distributions, which are typically exponential in computational complexity.
+In contrast, such calculations are required for the classical updates outlined in ??.
+This distinction relates directly to the foundational principles of quantum advantage.
+We intend to further investigate this property in future testing of QL algorithms.
+V End-to-End quantum-like information processing and measurement
+For genuine quantum resources, the primary source of error during information processing and measurement arises from the presence of uncontrolled noise.
+This can result from a variety of sources, including external electromagnetic fields, undesired interactions between qubits, fabrication defects, and imperfections in control protocols.
+Hou et al.
+(2016); Bengtsson et al.
+(2024); Funcke et al.
+(2022); Gupta and Nene (2021); Resch and Karpuzcu (2021) To account for these sources of error the final state of a quantum register after computation requires multiple measurements, which can lead to a significant overhead in the computing protocols.
+Likewise, error mitigation techniques and decoding protocols can significantly affect scalability and performance.
+Bravyi et al.
+(2021); Ali et al.
+(2024).
+In the following, we analyze both the accuracy and computational complexity of information processing and measurement within the QL formalism.
+Given that QL resources are intrinsically classical, noise from external sources of dissipation does not appear to be a fundamental limitation in our approach.
+Instead, the sources of errors are intrinsic to the formalism itself, impacting on an equal footing information processing and final measurement.
+Here, we present an analysis of these errors by examining the overall workflow of QL information protocols from end to end.
+Let us consider a generic QL algorithm consisting of a set of $M$ gates $\mathrm{g}=\{g_{1},\cdots,g_{M}\}$.
+This corresponds to the chain of transformations $U_{\bm{\mathrm{g}}}=\prod_{j=M}^{1}U_{g_{j}}$, as defined in ??.
+An exact approach to QL computing would involve applying this sequence of gates to transform via ??
+the initial register $\mathcal{R}$ [??].
+This consists of the Cartesian product of ${N_{\mathrm{QL}}}$ factors, as illustrated in panel a) of ??.
+Hence, storing and processing information following this exact strategy involves an exponential overhead with the number of edges in the total resource.
+A key insight we aim to leverage to reduce the complexity of this approach is as follows: if we can develop approximate QL protocols that maintain the structure of the Cartesian product, the exponential operations required for information processing and measurement could be circumvented entirely.
+We refer to this approach as the Cartesian-product approximation (CPA).
+This strategy enables to implement gate operations by transforming only a small subset of the edges of the total resources.
+In fact, within the CPA, information can be conveniently stored and processed with linear complexity in ${N_{\mathrm{QL}}}$ [see panel b) of ??].
+Firstly, we assess the degree to which the CPA holds within our approach and quantify the associated errors relative to exact QL computing.
+Conveniently, the Cartesian-product structure is exactly preserved by all one-QL-bit gates , defined by
+$W^{(p)}_{\mathrm{g}}=\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes U^{(p)}_{\mathrm{g}}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)},$ (V.1)
+acting locally on a single resource $p\in\{1,\cdots,{N_{\mathrm{QL}}}\}$. ?? transforms the initial graph resource ?? as
+$\displaylimits_{q=1}^{N_{\mathrm{QL}}}\mathcal{G}^{(q)}\;\begin{subarray}{c}U^{(p)}_{\mathrm{g}}\\ \longmapsto\end{subarray}\;\mathcal{G}^{(1)}\square\cdots\square\mathcal{G}^{(p)}_{\mathrm{g}}\square\cdots\square\mathcal{G}^{({N_{\mathrm{QL}}})},$ (V.2)
+where $\mathcal{G}^{(p)}_{\mathrm{g}}$ is the gate transformation of the graph $\mathcal{G}^{(p)}$.
+To extend the approach to a universal set of gates, we analyze how effectively the CPA approximation holds for the CNOT operation $U_{\mathrm{CNOT}}^{(q,p)}$, applying the Pauli-X gate to the $p$-th QL bit depending on the state of the control $q$.
+As shown in ??, this gate conserves the Cartesian product only approximately, namely
+$\displaylimits_{q=1}^{N_{\mathrm{QL}}}\mathcal{G}^{(q)}\;$ $\begin{subarray}{c}U_{\mathrm{CNOT}}^{(q,p)}\\ \longmapsto\end{subarray}\;\mathcal{G}^{(1)}\square\cdots\square\tilde{\mathcal{G}}^{(p)}\square\cdots\square\tilde{\mathcal{G}}^{(q)}\square\cdots\square\mathcal{G}^{({N_{\mathrm{QL}}})}$
+$\quad+\mathcal{O}(l/k).$ (V.3)
+??
+introduces an error which depends on the ratio between the valencies $l$ and $k$ of the elementary subgraphs in ??, respectively $C$ and $A_{i}$.
+This result indicates that the optimal choice for the parameters defining the subgraphs are
+$l=1,\hskip 14.22636ptk={N_{G}}-1,\hskip 14.22636pt{N_{G}}\gg 1.$ (V.4)
+In future work, we aim to investigate the propagation of error in ?? as several gates are concatenated to construct QL algorithms.
+An illustration of the final measurement process after information processing is shown in panel c) of ??. Let us firstly consider factorized measurement operators, with structure
+$\mathcal{F}_{{\bm{g}}^{\prime}}=\bigotimes_{q=1}^{N_{\mathrm{QL}}}U^{(q)}_{g^{\prime(q)}}.$ (V.5)
+Conveniently, even at the measurement state we can avoid to explicitly calculate the exponentially expensive Cartesian product for this common class of observable operators.
+In fact, as shown in panel c) of ??, local unitary transformations applied to one-QL-bit resources provide complete information on the measurement probabilities ??.
+For example, a local measurement of the population difference of the $q$-th QL bit acts on the final local emergent states as
+$\psi^{(q)}_{\sigma^{(q)},\bm{\mathrm{g}}}$ $\mapsto U_{\mathrm{Z}}^{(q)}\psi^{(q)}_{\sigma^{(q)},\bm{\mathrm{g}}},$ (V.6a)
+[…]
+The more general case of non-factorized observables involves the sum over $M$ terms ?? expressed by
+$\mathcal{F}_{\bm{g}_{1}^{\prime},\cdots,\bm{g}_{M}^{\prime}}=\sum_{m=1}^{M}\mathcal{F}_{\bm{g}^{\prime}_{m}}.$ (V.7)
+The action of ??
+breaks the structure of the Cartesian product, leading to a measurement cost that scales exponentially with the number of QL bits.
+It is worth noting that in a broad class of quantum algorithms, including Grover’s search Giri and Korepin (2017) and variational quantum eigensolvers, Tilly et al.
+(2022) the measurement process involves exclusively factorized operators, with structure as in ??.
+In these cases, this local measurement process can be applied with no limitations.
+We aim to compile a comprehensive list of the potential applications of our approach in future work.
+Figure 5: a) Representation of the initial register, prior to information processing.
+The number of vertices of the total graph scales exponentially with ${N_{\mathrm{QL}}}$, the number of QL-bits.
+b) Information processing implemented within the CPA, as discussed in the main text.
+Here, we store and transform the single resources individually, with a favorable linear scaling with size.
+c) Illustration of the final measurement process.
+This involves the local diagonalization of the gate-transformed resources, to obtain ${N_{\mathrm{QL}}}$ sets of single-particle emergent states.
+Factorized measurements operators can be implemented locally on these resources, without calculating the tensor product associated to the final global emergent state.
+VI Conclusions
 In this paper, we expanded upon the general approach derived in Refs.
@@
 We introduced the Hilbert space for an arbitrary number of QL bits, and we derived explicitly the Bell states in dimension two.
-We then proposed in ??
-a notion of quantum gates applicable to QL states, and we provided concrete examples of their action.
-Finally, we discussed in ??
-a list of measurement axioms which allow to observe QL interference effects in our approach.
+In ??
+we proposed a notion of quantum gates applicable to QL states, and we provided concrete examples of their action.
+We discussed in ??
+a list of measurement axioms which allows to observe QL interference effects in our approach.
+Finally, we outlined in ??
+an efficient end-to-end workflow for building QL algorithms based on a controlled approximation for two-QL-bit gates.
 Several open questions remain, presenting diverse opportunities for future research.
-In the paper we introduced notions of entanglement, interference, and QL probabilities which avoid the expensive calculation of joint distributions.
+In the paper we introduced notions of entanglement, interference, and QL probabilities which avoid the expensive calculation of joint probability distributions in the measurement process.
 In future work we plan to investigate how each of these features contributes to achieving a “QL advantage” over strictly classical information processing.
-The ultimate test of the approach proposed in this work will be to implement benchmark algorithms, and evaluate their time efficiency and resource cost.
+The ultimate test of the approach proposed in this work will be to implement the pipeline proposed in ??
+(and related ??) to study QL algorithms, and evaluate their time efficiency and resource demand.
 Bharti et al.
 (2022)
-In the last decades, advanced error correction techniques have been developed to minimize loss of information and decoherence in quantum processors.
-Cai et al.
-(2023); Campbell (2024) Similarly, we can ask how to prevent and minimize errors in our framework, e.g.
-in presence of impurities and irregularities in the structure of the graphs.
-In this regard, we note that a foundational property of synchronizing graphs is that their robustness against uncontrolled noise and dissipation increases with size.
-Scholes (2023); Scholes and Amati (2024); Anwar et al.
-(2023) Therefore, we conjecture that error correction is achievable in our approach by simply increasing the number of classical nodes of the graphs.
-We aim to further explore this concept in future work, to provide bounds of fidelity of gate transformations as a function of the size of the system and strength of external noise.
 It can be insightful to directly relate our work to the rigorous approach to quantum non-Markovianity proposed in Refs.
@@
 beyond the first classical term.
-From an experimental standpoint, different options can be explored to implement synchronizing resources for QL information processing.
-Insulator-to-metal-transition devices and spin-torque oscillators are viable candidates, as they allow for both distributed and parallel computing, while ensuring good robustness to noise and fault tolerance.
-Raychowdhury et al.
-(2019) In future work, we aim to further analyze platforms to experimentally implement QL resources, and the properties required by different approaches for optimal network implementation and manipulation.
+In this paper, we introduced a universal set of gates comprising single QL-bit terms and the CNOT operation.
+Our controlled approximation is designed to preserve the Cartesian product structure by selectively neglecting a small contribution within the CNOT gate that would otherwise lead to symmetry breaking.
+Preserving this contribution would result in an exponential increase in complexity, posing significant challenges for both computational and experimental resources.
+An open question remains as to whether directly implementing an approximation of many-QL-bit gates, rather than decomposing them into one- and two-body elementary terms, could yield even greater accuracy.
+While any local decomposition is exact in standard unitary quantum computing, this is in general not the case within our QL approach.
+Additionally, alternative choices of universal sets may provide greater computational advantages.
+In future work, we aim to explore the impact of these different scenarios within our framework.
+From an experimental standpoint, various possibilities can be considered for implementing synchronizing resources in QL information processing.
+The research on oscillator-based computing, as reviewed in Ref.
+Csaba and Porod, 2020, offers a solid foundation for this purpose.
+Within this framework, insulator-to-metal-transition devices and spin-torque oscillators emerge as promising candidates, supporting distributed and parallel computing while ensuring resilience to noise and fault tolerance.Raychowdhury et al.
+(2019) We plan to further investigate platforms for the experimental implementation of QL resources, focusing on the properties needed by various approaches to achieve optimal network design and control.
+In recent years, memcomputing has emerged as an innovative computational framework that exploits the intrinsic time-nonlocal properties of classical dynamical systems to enable a terminal-agnostic framework to implement non-Boolean logic.
+Di Ventra and Traversa (2018) This approach, which allows the computation to occur in memory rather than through sequential logic operations, could be exploited by recasting classical synchronizing dynamics leading to emergent states into time-nonlocal equations of motion.
+Amati, Meyer, and Schilling (2019) The classical Nakajima–Mori–Zwanzig formalism, which captures the influence of past states on open classical systems, offers a strategy to incorporate memory effects directly into the QL formalism, potentially enhancing stability and coherence.
 The environmental implications of the large-scale industrial production of quantum processors are topics of ongoing debate.
@@
 Aifer, Thingna, and Deffner (2024)
-Finally, it can be a fascinating topic for future research to connect QL information processing introduced here to other information theories on human decision-making and psychology derived in the framework of contextual probability models.
-This idea has been extensively discussed in the literature of QL probabilistic models.Khrennikov (2016); Ozawa and Khrennikov (2020); Scholes (2024); Widdows, Rani, and Pothos (2023) Implementing our formalism on brain-like neuromorphic circuits Pal et al.
-(2024) can foster theoretical and experimental investigations in this framework.
-VI Acknowledgments
+Finally, it can be a fascinating topic for future research to connect QL information processing introduced here to other information theories on human decision-making and psychology derived in the framework of QL contextual probability models.
+Khrennikov (2016); Ozawa and Khrennikov (2020); Scholes (2024); Widdows, Rani, and Pothos (2023) Implementing our formalism on brain-like neuromorphic circuits Pal et al.
+(2024) can support theoretical and experimental investigations in this framework.
+VII Acknowledgments
 This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
@@
 In general, this transformation cannot be expressed in analytic form as soon as the subgraphs defining the resources are not exactly regular.
-However, we show here that the problem of determining the map can be rewritten as a matrix inversion problem, easily solvable with standard numeric algebra tools.
+However, we show here that the problem of determining the map can be rewritten as a simple matrix inversion problem.
 Firstly, it is convenient to parametrize the generalization of ?? as follows:
 $U_{\mathrm{cb}}=\bigotimes_{q=1}^{N_{\mathrm{QL}}}\begin{pmatrix}X^{(q)}&Y^{(q)}\\ W^{(q)}&Z^{(q)}\end{pmatrix},$ (A.1)
-where $X^{(q)}\in\mathbb{R}^{{N_{G}}}$ is diagonal, with components $X_{ij}^{(q)}=\delta_{ij}x_{j}^{(q)}$, and similarly for the other three blocks in ??. ?? can be rewritten as the linear problem
+where $X^{(q)}\in\mathbb{R}^{{N_{G}}}$ is diagonal, with components $X_{ij}^{(q)}=\delta_{ij}x_{j}^{(q)}$, and similarly for the other three blocks in ??. ???? can be rewritten as the linear problem
 $x_{i}^{(q)}a_{1,i}^{(q)}+y_{i}^{(q)}a_{2,i}^{(q)}$ $={N_{G}}^{-1/2},$ (A.2a)
@@
 $w_{i}^{(q)}a_{1,i}^{(q)}-z_{i}^{(q)}a_{2,i}^{(q)}$ $={N_{G}}^{-1/2},$ (A.2d)
-where $q=1,\cdots,{N_{\mathrm{QL}}}$, $i=1,\cdots,{N_{G}}$, and where we made use of the generalization of ?? to arbitrary values of ${N_{\mathrm{QL}}}$:
-$\psi_{\pm}^{(q)}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}^{(q)}\\ a_{2}^{(q)}\end{pmatrix}.$ (A.3)
-?? is expressed in matrix form as
+where $q=1,\cdots,{N_{\mathrm{QL}}}$, $i=1,\cdots,{N_{G}}$. ?? is expressed in matrix form as
 $\begin{pmatrix}D_{1}^{(q)}&D_{2}^{(q)}&\mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}&D_{1}^{(q)}&D_{2}^{(q)}\\ D_{1}^{(q)}&-D_{2}^{(q)}&\mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}&D_{1}^{(q)}&-D_{2}^{(q)}\end{pmatrix}\begin{pmatrix}x^{(q)}\\ y^{(q)}\\ w^{(q)}\\ z^{(q)}\end{pmatrix}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\\ 0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (A.3)
@@
     Let us order the eigenvalues and eigenvectors of $\mathcal{F}$ in increasing order, via the map
-    $\bm{\sigma}\mapsto n=\sum_{q=1}^{N_{\mathrm{QL}}}2^{{N_{\mathrm{QL}}}-q}\left(\frac{1-\sigma^{(q)}}{2}\right)+1.$ (C.3)
-    Let us define by $V_{1}$ the set of vectors orthogonal to the eigenvector $f_{1}$ of $\mathcal{F}$ with eigenvalue $\lambda_{1}$.
+    $\bm{\sigma}\mapsto n=\sum_{l=0}^{{N_{\mathrm{QL}}}-1}2^{l}\delta_{\sigma_{l},\uparrow}.$ (C.3)
+    We denote by $V_{1}$ the set of vectors orthogonal to the eigenvector $f_{1}$ of $\mathcal{F}$ with eigenvalue $\lambda_{1}$.
 Note that $\mathcal{F}$ maps $V_{1}$ onto itself, i.e., $\mathcal{F}\psi\in V_{1}$ for $\psi\in V_{1}$.
@@
 This process allows us to define a complete set of mutually orthogonal eigenvectors.
+Appendix D Controlled-NOT gate
+In this appendix, we analyze the action of the CNOT gate on a pair of QL bits, $p,q\in{1,\dots,{N_{\mathrm{QL}}}}$.
+Here, we designate $p$ as the control QL bit, whose state determines whether to flip the state of the target QL bit $q$.
+For simplicity, we assume $p<q$; the case $p>q$ can be derived analogously.
+The unitary map implementing the transformation is
+$U_{\mathrm{CNOT}}^{(p,q)}$ $=\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\uparrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}$
+$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\downarrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes U_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)},$ (D.1)
+where we defined the projection operators
+$\mathcal{P}_{i}^{(p)}=\tfrac{1}{2}\left[U_{\mathcal{I}}^{(p)}+(-1)^{i}U_{\mathrm{Z}}^{(p)}\right],$ (D.2)
+with $p\in\{1,\cdots,{N_{\mathrm{QL}}}\}$ and $i\in\{0,1\}$.
+These are mutually orthogonal [$\mathcal{P}_{\sigma}^{(p)}\mathcal{P}_{\sigma^{\prime}}^{(p)}=\delta_{\sigma\sigma^{\prime}}\mathcal{P}_{\sigma^{\prime}}^{(p)}$] and complementary [$\mathcal{P}_{\downarrow}^{(p)}+\mathcal{P}_{\uparrow}^{(p)}=U_{\mathcal{I}}^{(p)}$].
+The action of ??
+on the total resource ??
+is
+$\mathcal{R}_{\mathrm{CNOT}(q,p)}^{[{N_{\mathrm{QL}}}]}$ $=\sum_{\sigma\in\{\uparrow,\downarrow\}}\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{R}_{\sigma\sigma}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}+\sum_{\begin{subarray}{c}r=1\\ r\neq q\end{subarray}}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}^{(r)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}$ (D.3a)
+$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\uparrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes\mathcal{R}^{(q)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-q)}+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\downarrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes\mathcal{R}_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)}$ (D.3b)
+$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\left[\mathcal{R}_{\uparrow\downarrow}^{(p)}+\mathcal{R}_{\downarrow\uparrow}^{(p)}\right]\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes U_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)},$ (D.3c)
+where $\mathcal{R}^{(p)}_{\sigma\pi}=\mathcal{P}^{(p)}_{\sigma}\mathcal{R}^{(p)}\left[\mathcal{P}^{(p)}_{\pi}\right]^{\dagger}$.
+As introduced in ??, we can minimize the computational complexity of this gate by assessing to which extent the structure of ??
+can be approximated by a Cartesian product.
+This property enables information through transformations applied exclusively to individual one-QL-bit resources.
+Let us therefore study separately all terms which would break the Cartesian product construction.
+Starting from the two terms in ??, we can study the relative error
+$\frac{\left\lVert\mathcal{R}_{\mathrm{X}}^{(q)}-\mathcal{R}^{(q)}\right\rVert}{\left\lVert\mathcal{R}^{(q)}\right\rVert}=\sqrt{\frac{8l{N_{G}}}{2k{N_{G}}+2l{N_{G}}}}=\frac{2}{\sqrt{1+\frac{k}{2l}}}\simeq 1+\frac{2l}{k}.$ (D.4)
+In ?? we introduced the Frobenius norm, defined by $\left\lVert A\right\rVert^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}|A_{ij}|^{2}$, for each $A\in M_{m\times n}(\mathbb{C})$. Also, we noticed that
+$\mathcal{R}^{(q)}_{\mathrm{X}}=U_{\mathrm{X}}^{[q]}\mathcal{R}^{(q)}\left[U_{\mathrm{X}}^{[q]}\right]^{\dagger}=\begin{pmatrix}A_{1}^{(q)}&-C^{(q)}\\ -\left[C^{(q)}\right]^{\top}&A_{2}^{(q)}\end{pmatrix}.$ (D.5)
+?? is minimized by the parameter choice defined in ??. Finally, let us consider the cross contribution in ??. This terms involves the difference
+$\mathcal{R}^{(p)}_{\uparrow\downarrow}+\mathcal{R}^{(p)}_{\downarrow\uparrow}$ $=\frac{1}{2}(\mathcal{R}^{(p)}-\mathcal{R}^{(p)}_{Z})$
+$=\frac{1}{2}\begin{pmatrix}A_{1}^{(p)}-A_{2}^{(p)}&C^{(p)}-\left[C^{(p)}\right]^{\top}\\ \left[C^{(p)}\right]^{\top}-C^{(p)}&A_{2}^{(p)}-A_{1}^{(p)}\end{pmatrix},$ (D.6)
+where
+$\mathcal{R}^{(q)}_{\mathrm{Z}}=\begin{pmatrix}A_{2}^{(q)}&\left[C^{(q)}\right]^{\top}\\ C^{(q)}&A_{1}^{(q)}\end{pmatrix}$ (D.7)
+is calculated similarly to ??. Without loss of generality it is possible to choose here
+$A_{1}^{(p)}$ $=A_{2}^{(p)},$ (D.8a)
+$C^{(p)}$ $=\left[C^{(p)}\right]^{\top},\hskip 14.22636pt\forall\;p=1,\cdots,{N_{\mathrm{QL}}}.$ (D.8b)
+In essence, we obtain
+$\mathcal{R}_{\mathrm{CNOT}(q,p)}^{[{N_{\mathrm{QL}}}]}$ $=\sum_{\sigma\in\mathcal{S}}\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{R}_{\sigma\sigma}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}$
+$\quad+\sum_{\begin{subarray}{c}r=1\\ r\neq q\end{subarray}}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}^{(r)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}$
+$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}_{\mathrm{X}}^{(q)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}+\mathcal{O}(l/k)$ (D.9)
+provided the conditions in ?? hold.

### v5

@@
 Abstract
-In previous work we have proposed a construction of quantum-like bits that could endow a large, complex classical system, for example of oscillators, with quantum-like function that is not compromised by decoherence.
+In previous work we have proposed a construction of quantum-like bits that could endow a large synchronizing classical system, for example of oscillators, with quantum-like function that is not compromised by decoherence.
 In the present paper we investigate further this platform of quantum-like states.
-Firstly, we discuss a general protocol on how to construct synchronizing networks that allow for emergent states.
+Firstly, we discuss a general protocol on how to construct classical synchronizing networks that allow for emergent states.
 We then study how gates can be implemented on those states.
-This suggests the possibility of quantum-like computing on specially-constructed classical networks.
-Finally, we define a notion of measurement that allows for non-Kolmogorov interference, a feature that separates our model from a classical probabilistic system.
-This paper aims to explore the mathematical structure of quantum-like resources, and shows how arbitrary gates can be implemented by manipulating emergent states in those systems.
+This suggests the possibility of quantum-like information processing on a special class of many-body classical systems.
+Finally, we show that our approach allows for non-Kolmogorov interference, a feature that separates our model from a classical probabilistic system.
+This paper aims to explore the mathematical structure of quantum-like resources distilled from classical synchronizing systems, and shows how arbitrary gates can be implemented by manipulating many-body correlations.
 I Introduction
@@
 Gharehchopogh (2023)
-In the spirit of quantum-inspired computing, we showed in recent work that a large class of synchronizing classical systems can be utilized as a resource for information processing.
+A large class of regular graphs encodes the correlations required to identify robust collective states.
+This general feature can be exploited for a variety of applications.
+For example, we discussed in Ref.
+Scholes, 2025a how this principle can support the engineering of molecular delocalization.
+The same reference, along with Refs.
+Krebs and Shaheen, 2011; Brouwer and Haemers, 2011, offers the interested reader background information on the mathematical formalism of regular and expander graphs, as well as their spectral properties.
+In the spirit of quantum-inspired computing, we showed in recent work that network correlations can serve as resources for quantum information processing.
 Scholes (2020, 2023, 2024); Scholes and Amati (2025) We focused, in particular, on many body systems whose correlations obey the graph structure shown in panel a) of ??.
 Here, two densely connected regions (in blue) of a network are mutually interacting via sparse links (in red).
-A special feature of this class of networks is that they produce pairs of emergent states, obtained by the direct sum of the emergent states of the blue subgraphs, with identical and opposite phases.
-The key requirement of the blue subgraphs, then, is that they exhibit an emergent state—a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs.
-Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) In our framework, emergent states can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??).
-Their robustness against noise and stochastic disorder increases with system size, exhibiting a favorable trend in contrast to the typical behavior of genuine quantum resources.
-Scholes (2023, 2020) This intriguing feature is connected to the occurrence of classical synchronization phenomena.
+A spectral analysis of this class of graphs highlights the presence of two isolated emergent states.
+These an can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??).
+The robustness of emergent states against noise appears to be a direct consequence of synchronization.
+Scholes (2023, 2020) This intriguing feature is related to the occurrence of classical synchronization phenomena.
 Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) We showed in Ref.
-Scholes, 2024 that the emergent states of connected expander subgraphs are isomorphic to the Hilbert space of a qubit.
+Scholes, 2024 that the emergent states of connected expander graphs are isomorphic to the Hilbert space of a qubit.
 This led us to introducing the notion of quantum-like (QL) bits within the context of classically-synchronizing networks.
 In Ref.
-Scholes and Amati, 2024 we built on these results and proposed a general framework to construct resources for multiple QL bits.
+Scholes and Amati, 2025 we built on these results and proposed a general framework to construct resources for multiple QL bits.
 In particular, we showed that the Cartesian product ($\square$) of ${N_{\mathrm{QL}}}$ bipartite graphs obeys the same dimensionality of a register of ${N_{\mathrm{QL}}}$ qubits [see panel b) of ??].
 In the present paper, we build on the foundational work of Ref.
-Scholes, 2024; Scholes and Amati, 2024; Scholes, 2023 to formalize the concept of QL states and gates.
-We present a computational framework for implementing information processing within the QL formalism through appropriate manipulations of the emergent states of network resources.
-As depicted in panel c) of ??, we show that an arbitrary gate transformation $V_{\mathrm{g}}$ can be expressed both as graph transformation, and as a map of the emergent states of the graph.
-Finally, we propose a measurement approach upon information processing that aligns with the QL probabilistic framework developed by Khrennikov Khrennikov (2007a, 2019, 2010).
-This theory shows that genuine quantum systems are not strictly required to process information in a QL way.
-Instead, to observe QL phenomena it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and their measurement.
-Consistently with Khrennikov’s approach, Khrennikov (2010) we list a set of measurement axioms which allows to define a generalized notion of nonclassical interference between observable operators.
+Scholes, 2024; Scholes and Amati, 2025; Scholes, 2023, by proposing a self-contained formalization of the concepts of QL states and gates and their application to classical many-body systems.
+As sketched in ??, we present a framework to map the emergent states of a classical synchronizing network of oscillators onto a qubit Hilbert space.
+We highlight a connection between emergent states and steady-state synchronization dynamics, and we show that the latter is robust even in presence of stochastic disorder in the network.
+This makes the emergent states viable candidates for QL information processing, which we implement by transforming the matrix of the two-body correlations between the oscillators.
+Our analysis is general and applies to any classical system whose two-body correlations align with the network structures discussed later in this paper.
+However, to make the approach more concrete, we provide explicit examples in this manuscript by examining the synchronization dynamics of the Kuramoto model.
+This is a non-symplectic classical system consisting of a network of nonlinear oscillators, whose collective properties have been extensively explored in the literature.
+Kuramoto (1975); Kawamura et al.
+(2010); Scholes (2022); Rodrigues et al.
+(2016)
+Our approach leads to a natural correspondence between “QL gates” transformations and the steady states of the classical oscillatory dynamics.
+Additionally, we propose a variety of experimental resources that can be utilized to test the formalism.
+This supports the applicability of the approach for information protocols involving low-dimensional Hilbert spaces, where the exponentially large complexity of the classical phase space can be handled by available experimental resources.
+Finally, we connect our perspective to the QL probabilistic framework developed by Khrennikov.
+Khrennikov (2007a, 2019, 2010) This theory formalizes the notion that genuine quantum systems are not required to process information in a QL way.
+Instead, to observe nonclassical interference it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and projective measurements.
 The present paper is structured as follows.
 In ??
-we summarize the formalism describing how to construct a single-QL-bit Hilbert space.
+we summarize our previous work describing how to construct a single-QL-bit Hilbert space.
 In ??
@@
 In ??
+we show the fundamental role of emergent states in the synchronizing dynamics of the Kuramoto model as a paradigmatic classical system of synchronizing oscillators.
+In ??
 we study the action of a complete set of QL gates, and we provide concrete examples of their action.
@@
 In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models.
-Khrennikov (2010) Finally, in ??, we outline the complete workflow for QL information processing and discuss its efficiency and accuracy.
+Khrennikov (2010) Finally, in ??, we propose a variety of experimental resources that could be utilized to implement and verify experimentally our formalism.
 In conjunction with Refs.
-Scholes, 2023; Scholes and Amati, 2024; Scholes, 2024, this work provides the mathematical foundation for a rigorous approach to QL information processing from classical synchronizing systems.
+Scholes, 2023; Scholes and Amati, 2025; Scholes, 2024, this work provides the mathematical foundation for a rigorous approach to QL information processing from classical many-body synchronizing systems.
 Figure 1: Illustration of the main concepts presented in this paper and in prior work by the authors.
-Scholes (2024); Scholes and Amati (2024) a) Many body systems consisting of networks of synchronizing dipoles can easily exhibit classical synchronization phenomena.
+Scholes (2024); Scholes and Amati (2025) a) Many body systems consisting of networks of oscillators can easily exhibit classical synchronization phenomena.
 By structuring two-body correlations into two densely connected regions (blue) weakly interacting through sparse links (red), synchronization gives rise to pairs of emergent states.
@@
 b) In Ref.
-Scholes and Amati, 2024, we demonstrated that the emergent states of the Cartesian product of a set of synchronizing graphs scale with the same dimensionality of a qubit system.
-In this paper, we expand over these concepts by providing further mathematical details on the formalism for this many-body problem.
-c) Additionally, we elaborate on the concept of information processing within the QL framework.
-We show that operations on graphs, implemented through suitable transformations of their edges, are equivalent to gate transformations $V_{\mathrm{g}}$ applied directly to the emergent states.
+Scholes and Amati, 2025, we demonstrated that the emergent states of the Cartesian product of a set of synchronizing graphs scale with the same dimensionality of a qubit system.
+In this paper, we expand over these concepts by bridging classical synchronization dynamics and the QL formalism.
+c) Additionally, we elaborate on the concept of QL information processing, and how this connects to steady-state synchronization dynamics.
+We show that operations on graphs, implemented through suitable transformations of their edges, are equivalent to gate transformations $\hat{V}_{\mathrm{g}}$ applied directly to the emergent states.
 II Quantum-like bits
-In this section we derive a mapping from a classical network to the Hilbert space of interacting QL bits.
+In this section we derive a mapping from the emergent states of classical networks to the Hilbert space of interacting QL bits.
 While the concept has been already elucidated in Ref.
 Scholes, 2024, here we lay out a more formal description.
-Later in the paper we use this formalism to study the notion of gates to perform quantum information processing, and we provide concrete examples.
+The present analysis provides a complete description of the axioms of states and operators in quantum mechanics, within the notation of the emergent states.
+Later in the paper we show the direct connection between this abstract approach and the dynamics of many-body classical systems.
 II.1 Isolated system
 Let us consider two classical many-body systems with pairwise interactions between the elementary constituents.
-Those systems can be mapped onto two graphs $G_{1}$ and $G_{2}$.
-We assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices.
-For simplicity we focus here on undirected simple graphs, without loops.
+The pair-wise interactions can be mapped onto two graphs $G_{1}$ and $G_{2}$.
+Here, we assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices.
+Also, we consider at this stage undirected simple graphs, without loops (while in the paper we will discuss graph transformations that break these symmetries).
 The adjacency matrix for this class of graphs is symmetric with entries either 00 or $1$ and null diagonal.
@@
 Diestel (2017) In particular,
-$\sum_{i=1}^{N_{G}}[A_{n}]_{ij^{\prime}}=\sum_{j=1}^{N_{G}}[A_{n}]_{i^{\prime}j}=k,\hskip 22.76219pt\forall i^{\prime},j^{\prime}=1,\cdots,{N_{G}},$ (II.1)
-where $n=1,2$.
-We highlight that, as proven in recent work, Scholes (2023); Scholes and Amati (2024) the present formalism is robust even in the presence of disorder in the network, which may violate the condition of exact regularity.
-Consequently, the theory discussed in the paper is also applicable to quasi-regular resources, provided analytical results on matrix algebra are replaced by their more general numerical analogs.
-The direct sum of the two adjacency matrices is
+$\sum_{i=1}^{N_{G}}[A_{n}]_{ij^{\prime}}=\sum_{j=1}^{N_{G}}[A_{n}]_{i^{\prime}j}=k,\hskip 22.76219pt\forall i^{\prime},j^{\prime}=1,\ldots,{N_{G}},$ (II.1)
+and $n=1,2$.
+A key property of these subgraphs is that they exhibit an emergent state—a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs.
+Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) The direct sum of the two adjacency matrices is
 $A_{1}\oplus A_{2}=\begin{bmatrix}A_{1}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&A_{2}\end{bmatrix},$ (II.2)
@@
 As proposed in Ref.
-Scholes, 2024, to define QL bits we switch on an interaction between the two subgraphs, and define the resource
+Scholes, 2024, we define a QL bit resource by switching on an interaction between the two subgraphs, and define the resource
 $\mathcal{R}=\begin{bmatrix}A_{1}&C\\ C^{\top}&A_{2}\end{bmatrix},$ (II.3)
@@
 By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information processing that can be controlled by network manipulation.
-We will elaborate on this important concept further in the manuscript.
+We will elaborate on these important concepts further in the manuscript.
 We study the spectral properties of the resources by plotting in ?? the distribution $\rho(\lambda)$ of the eigenvalues of ?? for increasing values of the valency $l$ of $C$.
@@
 where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$.
+As both the $A_{i}$’s and $C$ are exactly regular, then $a_{i}=(1,,1\ldots,1)^{\top}/\sqrt{2{N_{G}}}$ are vectors with constant weights.
+We keep, however, the notation general, to account for possible structural disorder in the network leading to this break of symmetry.
 The vectors $\psi_{\sigma}$, $\sigma\in\{\downarrow,\uparrow\}\equiv\mathcal{S}$, are eigenvectors of ??
 with eigenvalues $\{k-l,k+l\}$, respectively.
-The current arrow notation for the computational basis helps prevent any ambiguity with additional numeric subscripts and superscripts, which are used here and in the following sections for different purposes.
-In prior work we used the intention of the QL bit construction to represent explicitly the basis of the the subgraphs, so that the emergent state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and similarly $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2024) ??
+Note that in prior work we used the intention of the QL bit construction to represent explicitly the basis of the the subgraphs, so that the emergent state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and similarly $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2025) ??
 describes another basis choice, equally valid.
@@
 we expand on an idea proposed in Ref.
-Scholes and Amati, 2024, to represent entanglement and superposition by coupling together multiple graph resources.
+Scholes and Amati, 2025, to represent entanglement and superposition by coupling together multiple graph resources.
 Formal derivations on how to implement information processing and define a QL probabilistic model in this framework, follow in the later ????, respectively.
@@
 In Ref.
-Scholes and Amati, 2024 we propose a formal representation of states applicable to an arbitrary number of QL bits.
+Scholes and Amati, 2025 we propose a formal representation of states applicable to an arbitrary number of QL bits.
 This is based on the Cartesian product of single-QL-bit graphs.
@@
 are eigenvectors of ??.
-Here, different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\cdots,\sigma^{(N_{\mathrm{QL}})}\right\}$, while the components of ??
+Here, different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\cdots,\sigma^{(N_{\mathrm{QL}})}\right\}\in\bm{\mathcal{S}}$, where $\bm{\mathcal{S}}=\mathcal{S}^{N_{\mathrm{QL}}}$.
+The components of ??
 are expressed via the multi-index $\bm{i}=\left\{i^{(1)},\cdots,i^{(N_{\mathrm{QL}})}\right\}$ as
 $[\psi_{\bm{\sigma}}]_{\bm{i}}$ $=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}.$ (II.11)
-As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\{0,1\}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$.
+As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\mathcal{S}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$.
 Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$.
@@
 To obtain the explicit expression of the eigenvectors, it is convenient to label the eigenvectors of each QL bit as
-$\psi^{(1)}_{[\uparrow\downarrow]}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{[\uparrow\downarrow]}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.21)
+$\psi^{(1)}_{\uparrow/\downarrow}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\uparrow/\downarrow}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.21)
 where $a_{i}$’s and $b_{i}$’s are defined as in ??
 for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively.
-The square brackets indicate that the sign on the right-hand side of each expression (+ or -) is chosen depending on the direction of the arrow on the left-hand side ($\uparrow$ or $\downarrow$, respectively).
+The notation in ??
+indicate that the upper arrow ($\uparrow$) on the left-hand side is associated to the positive sign on right-hand side, and similarly for $\downarrow$ and the negative sign.
 With ??
 we can express
-$\psi_{\uparrow,[\uparrow\downarrow]}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\pm a_{2}\otimes b_{2}\end{pmatrix},$ (II.22a)
-$\psi_{\downarrow,[\uparrow\downarrow]}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}&\mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.22b)
+$\psi_{\uparrow,\uparrow/\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \pm a_{2}\otimes b_{2}\end{pmatrix},\hskip 14.22636pt\psi_{\downarrow,\uparrow\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}\\ \mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.22a)
 We can finally construct maximally entangled Bell states via the linear combinations
-$\Phi^{\mathrm{Bell}}_{[\uparrow\downarrow]}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.23a)
-$\Psi^{\mathrm{Bell}}_{[\uparrow\downarrow]}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.23b)
+$\Phi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.23a)
+$\Psi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.23b)
 that is,
-$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&a_{2}\otimes b_{2}\end{pmatrix},$ (II.24a)
-$\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix},$ (II.24b)
-$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&-a_{2}\otimes b_{2}\end{pmatrix},$ (II.24c)
-$\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&-a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix}.$ (II.24d)
-By applying the mapping from emergent states to the Hilbert space of QL bits, as discussed here, we can introduce the concept of QL gates for information processing.
-We propose an approach to implement these gates in the next ??.
-III Quantum-like gates
-In this section, we discuss the fundamental principles of QL information processing.
-Specifically, we demonstrate how the emergent states of QL resources can be transformed consistently to any arbitrary quantum gate.
-We then study how to translate these transformations into operations on the edges of the graph resources.
-This is particularly relevant for the future computational and experimental implementation of QL algorithms.
-To connect the usual definition of gates acting on an abstract Hilbert space to our QL formalism, it is convenient to introduce a map $U_{\mathrm{cb}}$ transforming the emergent eigenvectors onto the standard factorized computational basis.
-This map is defined by
-$U_{\mathrm{cb}}=[V_{\mathrm{H}}\otimes\mathbb{1}_{N_{G}}]^{\otimes{N_{\mathrm{QL}}}},$ (III.1)
+$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\\ \mathbb{0}_{{N_{G}}}\\ a_{2}\otimes b_{2}\end{pmatrix},$ $\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{{N_{G}}}\\ a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\end{pmatrix},$ (II.24a)
+$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\\ \mathbb{0}_{{N_{G}}}\\ -a_{2}\otimes b_{2}\end{pmatrix},$ $\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{{N_{G}}}\\ -a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\end{pmatrix}.$ (II.24b)
+So far, we have presented general results on the QL formalism and demonstrated that the approach enables an exact mapping to the Hilbert space of a qubit system.
+In the following ??, we explore how the abstract concepts of graph resources and emergent states relate to the classical dynamics of a system of oscillators.
+We discuss the resilience of this approach in the presence of noise in the network correlations and highlight the critical role of emergent states in the classical synchronizing dynamics.
+III Classical synchronization
+Here, we investigate classical synchronization in dynamical systems of oscillators, where two-body correlations are governed by a connectivity matrix with the structure described in ??.
+Systems that align with our formalism must, as a general requirement, involve classical many-body dynamics, where the two-body couplings between degrees of freedom are governed by a connectivity matrix $\mathcal{R}$, as outlined in ??.
+Additionally, the dynamics should guarantee that synchronization leads to a globally stable state that remains robust over long time scales.
+These features are satisfied by the non-symplectic dynamics of the Kuramoto model Kuramoto (1975); Kawamura et al.
+(2010); Scholes (2022); Rodrigues et al.
+(2016), which is characterized by its nonlinear equations of motion
+$\dot{\theta}_{i}=\omega_{i}+\frac{\Gamma}{2{N_{G}}}\sum_{j=1}^{2{N_{G}}}\mathcal{R}_{ij}\sin(\theta_{j}-\theta_{i}),\hskip 19.91692pti=1,\cdots,2{N_{G}}.$ (III.1)
+Here, the $\theta_{i}$’s are angular variables corresponding to $2{N_{G}}$ oscillators, defined on the vertices of a graph.
+The matrix $\mathcal{R}$ encodes the correlations between these degrees of freedom.
+The $\omega_{i}$’s denote the intrinsic oscillation frequencies.
+The time evolution of the angular variables of the model is shown in panel a) of ??.
+For our chosen system parameters (see caption of the figure) the system reaches a synchronized steady state, where all oscillators evolve at a uniform, constant phase.
+The synchronized steady state appears robust even in presence of disorder in the network, as shown in panel b) of the same figure.
+Here, white static disorder, with variance $\sigma=3$ and a zero mean, is added to all nonzero entries of the adjacency matrix $\mathcal{R}$, originally valued at one.
+Despite this strong perturbation, a collective synchronization state persists at long times, with only minor phase shifts affecting small subgroups of oscillators.
+This observation further supports the authors’ previous findings on the resilience of synchronization dynamics to noise–a property particularly appealing for robust information encoding and processing in the steady states of these classical dynamical systems.
+Scholes and Amati (2025); Scholes (2023, 2024)
+To examine the role of the emergent states in the synchronization dynamics it is convenient to express the adjacency matrix through its spectral decomposition, as
+$\mathcal{R}=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\lambda_{\bm{\sigma}}\left(\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}}^{\dagger}\right)+\sum_{i=1}^{M-2}\tilde{\lambda}_{i}\left(\tilde{\psi}_{i}\otimes\tilde{\psi}_{i}^{\dagger}\right),$ (III.2)
+where $M=(2{N_{G}})^{N_{\mathrm{QL}}}$ denotes the total number of states.
+The first term in ??
+represents the contribution of the emergent states $\psi_{\bm{\sigma}}$’s.
+The second term corresponds to the random part of the spectrum, with eigenvalues $\tilde{\lambda}_{i}$’s and corresponding eigenvectors $\tilde{\psi}_{i}$’s.
+Panel b) of ??
+depicts the time evolution of the Kuramoto dynamics when only the random part of the spectrum is taken into account [that is, by imposing $\lambda_{\bm{\sigma}}\equiv 0$ $\forall\bm{\sigma}\in\bm{\mathcal{S}}$ in ??].
+In this scenario, synchronization fails to emerge, with all degrees of freedom maintaining their initial random phases throughout the entire time evolution.
+This highlights that the structure of two-body correlations and emergent states, encoded in the adjacency matrix $\mathcal{R}$, play a critical role in achieving classical synchronization.
+In the next ??, we build on these idea by exploring the connection between dynamical synchronization and information processing.
+This approach provides a framework for linking the steady states of classical many-body systems to quantum information protocols.
+Figure 5: a): Time evolution of a system with $602{N_{G}}=$ 60 oscillators governed by the Kuramoto model [??].
+The angular variables $\theta_{i}$ are initialized uniformly in $[0,2\pi)$, and the intrinsic frequencies $\omega_{i}$ are sampled uniformly from $(0,1)$.
+The coupling constant is set to $\Gamma=10$, and the connectivity matrix $\mathcal{R}$ is structured as a regular graph, as described in ??, with the valencies of the diagonal and off-diagonal blocks set to $k=29$ and $l=15$, respectively.
+b): Time evolution of the system under the same parameters and initial conditions as in a), but with disorder introduced into the connectivity matrix $\mathcal{R}$.
+Specifically, each nonzero entry of $\mathcal{R}$ is perturbed by random white noise with zero mean and variance $\sigma=3$.
+c): Dynamics of the Kuramoto model with the connectivity matrix restricted to its random components [see discussion following ??].
+The system parameters and initial conditions are the same as in a) and b).
+IV Quantum-like gates
+In this section, we explore the fundamental principles of QL information processing.
+In ??, we formalize how the emergent states of QL resources can be systematically transformed according to arbitrary quantum gates.
+We further investigate the correspondence between state transformations and network resources.
+In ??, we connect this abstract formalism to the dynamics of classical synchronization.
+Specifically, we demonstrate how QL gates can be practically realized in classical many-body systems of synchronizing oscillators and analyze the impact of these transformations on the long-time steady states of their dynamics.
+IV.1 Formalism
+To connect the usual definition of gates acting on an abstract Hilbert space to our QL formalism, it is convenient to introduce a map $U_{\mathrm{cb}}$ transforming the emergent eigenvectors onto the standard computational basis.
+As that all elementary subrgaphs in our resources [????] are regular, this map can be expressed analytically by
+$U_{\mathrm{cb}}=[V_{\mathrm{H}}\otimes\mathbb{1}_{N_{G}}]^{\otimes{N_{\mathrm{QL}}}},$ (IV.1)
 where
-$V_{\mathrm{H}}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (III.2)
-denotes the standard Hadamard (H) gate. The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$. Here,
-$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (III.3)
-where
-$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (III.4)
+$V_{\mathrm{H}}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (IV.2)
+denotes the matrix representation of the Hadamard (H) gate in the computational basis. The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$, where
+$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (IV.3)
+and
+$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (IV.4)
+Note that, given that all the basis subgraphs are regular, the solution to ??
+is independent of their local structure.
 We show in ??
-how to construct the transformation $U_{\mathrm{cb}}$ in the more general case of non-regular graphs.
-The QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ is defined by the matrix transformation
-$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (III.5)
-Let us consider a few examples of ??. The QL maps of the identity operator and the three Pauli gates are, in the limit of exactly regular graphs,
-$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.6a)
-$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.6b)
-$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.6c)
-$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.6d)
+that a solution for the map $U_{\mathrm{cb}}$ can be determined if the resources are not exactly regular, by solving numerically a linear algebra problem.
+We define the QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ by the matrix transformation
+$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (IV.5)
+To elucidate the action of ??, let us consider a few examples. The QL maps of the identity operator ($\mathcal{I}$) and the three Pauli gates (X,Y,Z) are
+$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.6a)
+$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.6b)
+$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.6c)
+$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.6d)
 respectively. As expected, the transformations in ?? act as the standard Pauli gates:
 […]
-As another example, the QL map of the Hadamard gate entangles the states of the factorized basis:
-$U_{\mathrm{H}}\psi_{[\uparrow\downarrow]}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{[\uparrow\downarrow]}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (III.8)
+As another example, the QL map corresponding to the Hadamard gate transforms states in the factorized basis into superposition states:
+$U_{\mathrm{H}}\psi_{\uparrow/\downarrow}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\uparrow/\downarrow}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (IV.8)
+The formal approach proposed here applies equally to any arbitrary gates, including those outside the Clifford group.
+While gates outside the Clifford group break the stabilizer structure and require exponentially more resources compared to Clifford gates—which can be efficiently simulated classically—the distinction between these two classes of gates is not relevant to our framework.
+Our formalism is designed to be computationally exact, addressing the exponentially high cost associated with scaling QL resources, regardless of the type of gate involved.
+The action of the non-Clifford T-gate on the factorized basis is
+$U_{\mathrm{T}}\psi_{\uparrow/\downarrow}=\frac{1}{2}\left[(1+\,\mathrm{e}^{i\pi/4})\psi_{\uparrow}+(1-\,\mathrm{e}^{i\pi/4})\psi_{\downarrow}\right].$ (IV.9)
 […]
-$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}},$ (III.9)
-transforming elements of the factorized basis as follows:
-$U_{\mathrm{CN}}\psi_{\uparrow[\uparrow,\downarrow]}=\psi_{\uparrow[\uparrow,\downarrow]},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow[\uparrow,\downarrow]}=\psi_{\downarrow[\downarrow,\uparrow]}.$ (III.10)
-Note that, since a universal set of gates can be constructed by combining one-QL-bit gates and the CNOT, we can limit our analysis to this set of operations. Barenco et al. (1995)
+$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}}$ (IV.10)
+transforms the elements of the factorized basis as follows:
+$U_{\mathrm{CN}}\psi_{\uparrow[\uparrow,\downarrow]}=\psi_{\uparrow[\uparrow,\downarrow]},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow[\uparrow,\downarrow]}=\psi_{\downarrow[\downarrow,\uparrow]}.$ (IV.11)
 Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis.
 We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??.
-From an operational standpoint, it is however preferable to translate these state operations onto transformations applicable directed to the graph resources.
-This feature is key to translate our gate formalism into practical algorithms for classical computing resources and for experimental platforms of synchronizing electrical circuits.
-Ochs, Michaelis, and Roggendorf (2019) In this regard, we observe that the network transformation
-$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}$ (III.11)
-transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??. This can be inferred from the spectral decomposition of ??, given by
-$\mathcal{R}=\sum_{i=1}^{\tilde{N}}\lambda_{i}\psi_{i}\otimes\psi_{i}^{\dagger}.$ (III.12)
-Here, $\tilde{N}=(2{N_{G}})^{N_{\mathrm{QL}}}$ denotes the total number of eigenstates of the resource $\mathcal{R}$, including both the emergent and the random states.
+From a practical standpoint, it is preferable to translate these state operations onto transformations applicable directed to the graph resources.
+This feature is a key requirement for a future experimental validation of our formalism with experimental resources.
+(An extended discussion on experimental resources is provided in the next ??).
+A straightforward approach in this regard is to define the unitary map
+$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}.$ (IV.12)
+??
+transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??.
+This property can be easily inferred by expanding $\mathcal{R}$ on the right-hand side of ??
+via the spectral decomposition in ??.
 In the next ??
-we shift the discussion towards the benefits provided by a suitable measure theory designed to add “quantumness” to classical synchronizing networks.
-This framework fulfills premises which can be key for a future quantification of QL advantage, as originally proposed in Ref.
-Khrennikov, 2019.
-IV Quantum-like measurements
-In this section, we discuss criteria that allow to introduce a notion of QL interference of probabilities in our approach.
+we provide concrete examples of ??, and we discuss how this network transformation affects the steady-state synchronizing solution of a system of oscillators.
+IV.2 Connection to classical synchronization
+Here, we explore the relationship between gate transformations of QL resources and the synchronization of classical many-body oscillator systems.
+Specifically, we focus on two examples of one-QL-bit gates.
+In these cases, the synchronization patterns of the classical system can be directly linked to the structure of the corresponding adjacency matrices.
+However, our approach is general and can be extended to arbitrary QL gates without any restrictions.
+The Pauli-X (NOT) gate transforms an initial resource ?? as follows:
+$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{X}}=\begin{pmatrix}A_{1}&-C\\ -C^{\top}&A_{2}\end{pmatrix}.$ (IV.13)
+We show in panel a) of ??, the synchronization pattern in the dynamics of the Kuramoto model with correlation matrix $\mathcal{R}_{\mathrm{X}}$.
+In the transformation ??, the leading emergent eigenvector, associated to the largest eigenvalue $k+l$, is mapped from $\psi_{\uparrow}$ to $\psi_{\downarrow}$ [see ??].
+The change of sign in the eigenvectors corresponds to a phase shift of $\Delta\phi=\pi$ between the first and the second half of the oscillators, as shown in the figure.
+In the case of the Hadamard gate, resources are transformed according to
+$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{H}}=\begin{pmatrix}A_{1}+A_{2}+C+C^{\top}&A_{1}-A_{2}-C+C^{\top}\\ A_{1}-A_{2}+C-C^{\top}&A_{1}+A_{2}-C-C^{\top}\end{pmatrix}.$ (IV.14)
+The sum and difference of the basis graphs $A_{1}$ and $A_{2}$ in ??
+highlight the fundamental role of this gate in creating a superposition state.
+The structure of ??
+encodes key information about the expected synchronization patterns in the dynamics of the corresponding Kuramoto model, as depicted in panel b) of ??.
+Dephasing emerges between the first and second halves of the oscillators, primarily driven by cancellation effects in the two off-diagonal blocks, where the term $A_{1}-A_{2}$ plays a dominant role.
+When $A_{1}\simeq A_{2}$ and $C\simeq C^{\top}$, the off-diagonal blocks approach zero, leading to completely uncorrelated phases between the two oscillator subgroups over extended timescales.
+Furthermore, the second half of the oscillators, labeled by indices $i\in\{{N_{G}},\cdots 2{N_{G}}\}$ exhibits weaker synchronization compared to the first one.
+This is due to the fact that the matrix norm of the lower diagonal block in ??
+is, on average, smaller than that of the upper diagonal block.
+Consequently, the second group of oscillators exhibits weaker correlations compared to the first group.
+In this section, we derived general principles for mapping network resources in a manner consistent with an arbitrary set of QL gates.
+Additionally, we presented numerical evidence supporting the connection between QL gates and the steady-state synchronization of classical oscillator resources.
+In the following ??, we explore specific examples of experimental resources that could be utilized to implement our formalism in practice.
+V Experimental perspectives
+So far, we have shown that the emergent states of synchronizing classical systems can be used as a basis to map the Hilbert space of a system of qubits.
+We discussed the primary role of emergent states in synchronization dynamics and the robustness of the latter in the presence of noise.
+In this section, we explore perspectives on implementing the proposed formalism using experimental physical resources.
+Additionally, we provide a high-level outline of computational and experimental validations that could support the further development of this framework.
+For genuine quantum resources, the primary source of errors during information processing and measurement arises from uncontrolled noise.
+This noise can result from several factors, including external electromagnetic fields, unwanted interactions between qubits, fabrication defects, and imperfections in control protocols.
+Hou et al.
+(2016); Bengtsson et al.
+(2024); Funcke et al.
+(2022); Gupta and Nene (2021); Resch and Karpuzcu (2021) Addressing such errors typically requires multiple measurements to accurately resolve the final state of a quantum register after computations.
+This can introduce significant overhead in computing protocols and hinder the scalability of the approach to larger size.
+Error mitigation and decoding techniques, although conceptually scalable, can result in further limitations in size and performance.
+Bravyi et al.
+(2021); Ali et al.
+(2024) In our approach, error correction in the presence of noise translates to enhancing the robustness of long-time classical synchronization [see the discussion on panel b) of ??].
+This ensures distinct and readily measurable transformations of the steady-state dynamics.
+Optimal synchronization can be achieved by numerically analyzing phase diagrams of synchronization indicators.
+For instance, in the case of the Kuramoto model, time-resolved order parameters offer clear insight into the synchronization strength.
+Basnarkov and Urumov (2007); Wang, Xu, and Zheng (2021) Parameter domains with the highest degree of synchronization can be identified as optimal candidates for experimental implementation.
+In terms of resource complexity, the number of degrees of freedom required to model the synchronization dynamics of a system of ${N_{\mathrm{QL}}}$ QL bits scales exponentially with system size.
+Although this scaling does not impose fundamental constraints on computational studies or experimental implementations of the approach, it does limit the practical feasibility of large-scale QL algorithms.
+Here, we discuss a few experimental platforms that are suitable candidates for implementing QL information processing within these scaling constraints.
+Additionally, we identify the classes of algorithms that can be most effectively analyzed by those resources.
+Implementing coupled LC circuits is a straightforward approach to implement macroscopic experimental resources that exhibit classical synchronization.
+Allam et al.
+(2006); Barbi (2021) Numerical analyses from Ref.
+Maffezzoni et al., 2015 show that this class of systems can be effectively approximated by Kuramoto-type models for medium-scale systems of 60 oscillators.
+The couplings between the model oscillators are tuned by adjusting the transconductance values of the circuit transistors.
+Spin-torque oscillators (STO) emerge as another promising candidate for implementing QL information processing.
+Raychowdhury et al.
+(2019) STOs are nonlinear magnetic oscillators that can be integrated into nanoscale devices operating at microwave frequencies.
+They have the ability to synchronize in both frequency and phase with external oscillatory signals as well as with other STOs.
+The full magnetization dynamics of these systems, defined by the Landau-Lifshitz-Gilbert-Slonzewski (LLGS) equation, can be simplified to the Kuramoto model for small-amplitude oscillations.
+Belanovsky et al.
+(2013); Kanao et al.
+(2019) Ref.
+Flovik, Macià, and Wahlström, 2016 demonstrates the experimental feasibility of modeling synchronizing classical systems with up to a few hundred oscillators.
+Beyond this scale, technical constraints arising from finite-size effects restrict the extent of achievable synchronization.
+Metal-to-insulator transition (MIT) devices leverage materials that can switch between insulating and metallic (conductive) states in response to external conditions such as temperature, pressure, or electrical stimuli.
+Kim et al.
+(2010) Research on small-scale MIT circuits has demonstrated their fundamental ability to produce classical synchronization behavior through voltage-triggered oscillatory dynamics.
+Kim et al.
+(2010); Parihar et al.
+(2015); Raychowdhury et al.
+(2019) One limitation of current MIT devices is their low oscillation frequency, typically below 1 MHz, which hinders the ability to achieve fine-resolution synchronization and detect subtle variations in coupled oscillator behavior.
+However, ongoing advancements in scaling down these devices are expected to increase their oscillation frequencies, enabling more precise and time-resolved synchronization measurements.
+Cotter et al.
+(2014)
+Building on the experimental perspectives outlined above, it appears feasible to validate our QL-bit theory using networks of up to a hundred oscillators.
+Such an experimental setup would enable the implementation of proof-of-concept algorithms on systems with ${N_{\mathrm{QL}}}=2$ or ${N_{\mathrm{QL}}}=3$ QL-bits, where each QL-bit comprises a small network of 555510101010 oscillators.
+A particularly intriguing class of quantum information processes for systems of this scale includes those that generate entanglement, such as the formation of Bell states (for ${N_{\mathrm{QL}}}=2$) and their higher-order generalizations, the Greenberger–Horne–Zeilinger (GHZ) states.
+Greenberger, Horne, and Zeilinger (1989); Erhard, Krenn, and Zeilinger (2020) This analysis has the potential to provide valuable insights into fundamental questions surrounding the classical-to-quantum correspondence.
+In particular, it could help clarify the role of non-emergent states in classical synchronizing systems as potential hidden variables within a QL-based framework for analyzing entanglement.
+Scholes (2025b)
+We conclude this section by presenting a practical perspective on implementing an experimental scheme to test our approach:
+  1. 1.
+    Engineer a classical synchronizing system: Design a system with two-body correlations represented by a graph, ensuring that its leading emergent state corresponds to the desired initial state of a QL algorithm.
+For example, the factorized state $\psi_{\uparrow}$ can be associated with a Cartesian-product network structured as described in ??.
+  2. 2.
+    Initialize oscillator phases: Assign arbitrary initial conditions to the oscillator phases, ensuring they adhere to the constraint that the system’s long-time steady state achieves the expected synchronization.
+  3. 3.
+    Allow the classical dynamics to equilibrate: Let the system evolve until it reaches a steady state.
+This step requires time-resolved measurements to ensure the system has sufficient time to reach a long-time synchronized state.
+  4. 4.
+    Perform gate operations: Each gate operation consists of two phases: (a) transforming the correlations between the degrees of freedom and (b) allowing the system to synchronize and relax to a new emergent state.
+  5. 5.
+    Measure the final state: Apply projection operators to extract the state of the final resource after all gate operations have been implemented.
+The proposed scheme can be tested numerically by measuring multi-time correlation functions.
+This process involves a series of time evolution operators defined by gate-transformed resources at different stages of a given QL algorithm.
+Preliminary analytical work on the Kuramoto dynamics is discussed in ??
+[see ??].
+Figure 6: a): Time evolution of a Kuramoto model where the adjacency matrix is transformed according to the mapping in ??
+for the Pauli-X gate (g = X).
+b): Similar to panel (a), but for the Hadamard gate (g = H).
+The same model parameters as defined in ??
+are used here.
+VI Quantum-like interfence
+In this final section, we discuss criteria that allow to introduce a notion of QL interference of probabilities in our approach.
 This analysis highlights that genuine quantum mechanical systems are not strictly required to build QL stochastic models provided the notions of states, observables and measurement are defined according to a given set of prescriptions.
-IV.1 Axiomatic definition
+VI.1 Axiomatic definition
 In the following we present a list of axioms, originally proposed in Ref.
@@
 States can be expressed by linear combinations of elements of the factorized basis in ??:
-$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}}.$ (IV.1)
+$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}}.$ (VI.1)
 Pure QL states are by definition normalized, and we establish the equivalence class $\beta\psi\sim\psi$ for $\beta\in\mathbb{C}$, $|\beta|=1$.
 Axiom 2.
-QL observables are represented by self-adjointed operators on $\mathcal{H}$. Their action on QL states, expressed by a linear map, defines a QL measurement.
+QL observables are represented by self-adjoint operators on $\mathcal{H}$. Their action on QL states, expressed by a linear map, defines a QL measurement.
 Observables are defined in terms of linear combinations with structure
-$\mathcal{F}$ $=\sum_{{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}}\left(\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}+\gamma^{*}_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right),\hskip 14.22636pt\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\in\mathbb{C},$ (IV.2)
+$\mathcal{F}$ $=\sum_{{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}}\left(\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}+\gamma^{*}_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right),\hskip 14.22636pt\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\in\mathbb{C},$ (VI.2)
 and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$ [??]. The QL measurement of $\mathcal{F}$ on a state $\psi$ is defined by the map $\psi\mapsto\mathcal{F}\psi$.
@@
 Specifically, let us consider a generic state $\psi\in\mathcal{H}$. We define by
-$P_{\psi}\left(\lambda_{\bm{\sigma}}^{\mathcal{F}}\right)=\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}=|\braket{f_{\bm{\sigma}},\psi}|^{2},$ (IV.3)
+$P_{\psi}\left(\lambda_{\bm{\sigma}}^{\mathcal{F}}\right)=\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}=|\braket{f_{\bm{\sigma}},\psi}|^{2},$ (VI.3)
 the QL probability of collapsing onto the subspace spanned by the eigenvector $f_{\bm{\sigma}}$ with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$.
@@
 is expressed in terms of the projector
-$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}\otimes f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (IV.4)
+$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}\otimes f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (VI.4)
 Axiom 5.
@@
 In particular, let us consider two generic QL observable operators $\mathcal{F}$ and $\mathcal{G}$ whose spectra are defined by
-$\mathcal{F}f_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}},\hskip 28.45274pt\mathcal{G}g_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}.$ (IV.5)
+$\mathcal{F}f_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}},\hskip 28.45274pt\mathcal{G}g_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}.$ (VI.5)
 For a nonzero state $\psi$ we denote by
-$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)=\frac{\left\lVert\Pi^{\mathcal{G}}_{\bm{\sigma}^{\prime}}\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}{\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}$ (IV.6)
+$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)=\frac{\left\lVert\Pi^{\mathcal{G}}_{\bm{\sigma}^{\prime}}\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}{\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}$ (VI.6)
 the QL conditional probabilities associated to the consecutive measures of $\mathcal{F}$ and $\mathcal{G}$ on $\psi$.
@@
 Khrennikov (2003, 2019)
-IV.2 Quantum probabilities in a QL model
+VI.2 Quantum probabilities in a QL model
 Before introducing a formula for the update of QL probabilities, it is instructive to briefly recall its classical correspondent, to assess analogies and differences between the two.
@@
 $P^{(\mathrm{cl})}(\mathcal{B})$ $=P^{(\mathrm{cl})}\left(\mathcal{B}\cap\cup_{j=1}^{J}A_{j}\right)$
-$=\sum_{j=1}^{J}P^{(\mathrm{cl})}\left(\mathcal{B}\cap A_{j}\right)=\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})P^{(\mathrm{cl})}(\mathcal{A}_{j}),$ (IV.7)
+$=\sum_{j=1}^{J}P^{(\mathrm{cl})}\left(\mathcal{B}\cap A_{j}\right)=\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})P^{(\mathrm{cl})}(\mathcal{A}_{j}),$ (VI.7)
 where we introduced the classical conditional probability
-$P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})=\frac{P^{(\mathrm{cl})}(\mathcal{B}\cap\mathcal{A}_{j})}{P^{(\mathrm{cl})}(\mathcal{A}_{j})}.$ (IV.8)
+$P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})=\frac{P^{(\mathrm{cl})}(\mathcal{B}\cap\mathcal{A}_{j})}{P^{(\mathrm{cl})}(\mathcal{A}_{j})}.$ (VI.8)
 We discuss now the correspondent of ??
 for the present QL framework.
-Let $\mathcal{F}$ and $\mathcal{G}$ be two nondegenerate QL observbles, whose eigenvalue problems are defined as in ??.
+We restrict the following analysis to couples of non-degenerate QL observables, $\mathcal{F}$ and $\mathcal{G}$, whose eigenvalue problems are defined as in ??.
+The eigenvalues of these operators are associated with a one-dimensional subspaces, which can be fixed by selecting any normalized eigenvector.
 In this case, the QL conditional probabilities defined in ??
-lose their dependence on the initial state $\psi$:
-$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)$ $\mapsto|\braket{g_{\bm{\sigma}^{\prime}},f_{\bm{\sigma}}}|^{2}\equiv P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right).$ (IV.9)
+lose their dependence on the initial state $\psi$, and can be simply expressed as
+$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)$ $\mapsto|\braket{g_{\bm{\sigma}^{\prime}},f_{\bm{\sigma}}}|^{2}\equiv P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right).$ (VI.9)
 Note that the probabilities in ??
@@
 The role of contextuality in a QL probabilitistic framework can be highlighted by rewriting total and conditional probabilities [????, respectively], on an equal footing:
-$P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}}|\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}\right)=\left.P_{\psi}(\lambda^{\mathcal{G}}_{\bm{\sigma}})\right|_{\psi=f_{\bm{\sigma}^{\prime}}}.$ (IV.10)
+$P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}}|\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}\right)=\left.P_{\psi}(\lambda^{\mathcal{G}}_{\bm{\sigma}})\right|_{\psi=f_{\bm{\sigma}^{\prime}}}.$ (VI.10)
 If ?? is fulfilled, Axioms 1 to 5 lead to the QL formula for probability updates Khrennikov (2007b, 2010); Bulinski and Khrennikov (2004)
@@
 $\quad+\sum_{\begin{subarray}{c}\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}\in\bm{\mathcal{S}}\\ \bm{\sigma}^{\prime}\neq\bm{\sigma}^{\prime\prime}\end{subarray}}\Lambda^{\psi}_{\bm{\sigma},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}}\Big{[}P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime}}^{\mathcal{F}}\right)P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)$
-×$\quad\times P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime\prime}}^{\mathcal{F}}\right)P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)\Big{]}^{1/2}.$ (IV.11)
+×$\quad\times P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime\prime}}^{\mathcal{F}}\right)P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)\Big{]}^{1/2}.$ (VI.11)
 An extended derivation of ??
@@
 We intend to further investigate this property in future testing of QL algorithms.
-V End-to-End quantum-like information processing and measurement
-For genuine quantum resources, the primary source of error during information processing and measurement arises from the presence of uncontrolled noise.
-This can result from a variety of sources, including external electromagnetic fields, undesired interactions between qubits, fabrication defects, and imperfections in control protocols.
-Hou et al.
-(2016); Bengtsson et al.
-(2024); Funcke et al.
-(2022); Gupta and Nene (2021); Resch and Karpuzcu (2021) To account for these sources of error the final state of a quantum register after computation requires multiple measurements, which can lead to a significant overhead in the computing protocols.
-Likewise, error mitigation techniques and decoding protocols can significantly affect scalability and performance.
-Bravyi et al.
-(2021); Ali et al.
-(2024).
-In the following, we analyze both the accuracy and computational complexity of information processing and measurement within the QL formalism.
-Given that QL resources are intrinsically classical, noise from external sources of dissipation does not appear to be a fundamental limitation in our approach.
-Instead, the sources of errors are intrinsic to the formalism itself, impacting on an equal footing information processing and final measurement.
-Here, we present an analysis of these errors by examining the overall workflow of QL information protocols from end to end.
-Let us consider a generic QL algorithm consisting of a set of $M$ gates $\mathrm{g}=\{g_{1},\cdots,g_{M}\}$.
-This corresponds to the chain of transformations $U_{\bm{\mathrm{g}}}=\prod_{j=M}^{1}U_{g_{j}}$, as defined in ??.
-An exact approach to QL computing would involve applying this sequence of gates to transform via ??
-the initial register $\mathcal{R}$ [??].
-This consists of the Cartesian product of ${N_{\mathrm{QL}}}$ factors, as illustrated in panel a) of ??.
-Hence, storing and processing information following this exact strategy involves an exponential overhead with the number of edges in the total resource.
-A key insight we aim to leverage to reduce the complexity of this approach is as follows: if we can develop approximate QL protocols that maintain the structure of the Cartesian product, the exponential operations required for information processing and measurement could be circumvented entirely.
-We refer to this approach as the Cartesian-product approximation (CPA).
-This strategy enables to implement gate operations by transforming only a small subset of the edges of the total resources.
-In fact, within the CPA, information can be conveniently stored and processed with linear complexity in ${N_{\mathrm{QL}}}$ [see panel b) of ??].
-Firstly, we assess the degree to which the CPA holds within our approach and quantify the associated errors relative to exact QL computing.
-Conveniently, the Cartesian-product structure is exactly preserved by all one-QL-bit gates , defined by
-$W^{(p)}_{\mathrm{g}}=\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes U^{(p)}_{\mathrm{g}}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)},$ (V.1)
-acting locally on a single resource $p\in\{1,\cdots,{N_{\mathrm{QL}}}\}$. ?? transforms the initial graph resource ?? as
-$\displaylimits_{q=1}^{N_{\mathrm{QL}}}\mathcal{G}^{(q)}\;\begin{subarray}{c}U^{(p)}_{\mathrm{g}}\\ \longmapsto\end{subarray}\;\mathcal{G}^{(1)}\square\cdots\square\mathcal{G}^{(p)}_{\mathrm{g}}\square\cdots\square\mathcal{G}^{({N_{\mathrm{QL}}})},$ (V.2)
-where $\mathcal{G}^{(p)}_{\mathrm{g}}$ is the gate transformation of the graph $\mathcal{G}^{(p)}$.
-To extend the approach to a universal set of gates, we analyze how effectively the CPA approximation holds for the CNOT operation $U_{\mathrm{CNOT}}^{(q,p)}$, applying the Pauli-X gate to the $p$-th QL bit depending on the state of the control $q$.
-As shown in ??, this gate conserves the Cartesian product only approximately, namely
-$\displaylimits_{q=1}^{N_{\mathrm{QL}}}\mathcal{G}^{(q)}\;$ $\begin{subarray}{c}U_{\mathrm{CNOT}}^{(q,p)}\\ \longmapsto\end{subarray}\;\mathcal{G}^{(1)}\square\cdots\square\tilde{\mathcal{G}}^{(p)}\square\cdots\square\tilde{\mathcal{G}}^{(q)}\square\cdots\square\mathcal{G}^{({N_{\mathrm{QL}}})}$
-$\quad+\mathcal{O}(l/k).$ (V.3)
+VII Conclusions
+In this paper, we built upon the general framework developed in Refs.
+Scholes, 2023, 2024; Scholes and Amati, 2025 to define and manipulate QL states.
+In ??, we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs.
+We introduced the Hilbert space for an arbitrary number of QL bits and explicitly derived the Bell states in two dimensions.
+In ??, we proposed a framework for defining quantum gates applicable to QL states and provided concrete examples of their action on many-body classical resources.
 ??
-introduces an error which depends on the ratio between the valencies $l$ and $k$ of the elementary subgraphs in ??, respectively $C$ and $A_{i}$.
-This result indicates that the optimal choice for the parameters defining the subgraphs are
-$l=1,\hskip 14.22636ptk={N_{G}}-1,\hskip 14.22636pt{N_{G}}\gg 1.$ (V.4)
-In future work, we aim to investigate the propagation of error in ?? as several gates are concatenated to construct QL algorithms.
-An illustration of the final measurement process after information processing is shown in panel c) of ??. Let us firstly consider factorized measurement operators, with structure
-$\mathcal{F}_{{\bm{g}}^{\prime}}=\bigotimes_{q=1}^{N_{\mathrm{QL}}}U^{(q)}_{g^{\prime(q)}}.$ (V.5)
-Conveniently, even at the measurement state we can avoid to explicitly calculate the exponentially expensive Cartesian product for this common class of observable operators.
-In fact, as shown in panel c) of ??, local unitary transformations applied to one-QL-bit resources provide complete information on the measurement probabilities ??.
-For example, a local measurement of the population difference of the $q$-th QL bit acts on the final local emergent states as
-$\psi^{(q)}_{\sigma^{(q)},\bm{\mathrm{g}}}$ $\mapsto U_{\mathrm{Z}}^{(q)}\psi^{(q)}_{\sigma^{(q)},\bm{\mathrm{g}}},$ (V.6a)
-[…]
-The more general case of non-factorized observables involves the sum over $M$ terms ?? expressed by
-$\mathcal{F}_{\bm{g}_{1}^{\prime},\cdots,\bm{g}_{M}^{\prime}}=\sum_{m=1}^{M}\mathcal{F}_{\bm{g}^{\prime}_{m}}.$ (V.7)
-The action of ??
-breaks the structure of the Cartesian product, leading to a measurement cost that scales exponentially with the number of QL bits.
-It is worth noting that in a broad class of quantum algorithms, including Grover’s search Giri and Korepin (2017) and variational quantum eigensolvers, Tilly et al.
-(2022) the measurement process involves exclusively factorized operators, with structure as in ??.
-In these cases, this local measurement process can be applied with no limitations.
-We aim to compile a comprehensive list of the potential applications of our approach in future work.
-Figure 5: a) Representation of the initial register, prior to information processing.
-The number of vertices of the total graph scales exponentially with ${N_{\mathrm{QL}}}$, the number of QL-bits.
-b) Information processing implemented within the CPA, as discussed in the main text.
-Here, we store and transform the single resources individually, with a favorable linear scaling with size.
-c) Illustration of the final measurement process.
-This involves the local diagonalization of the gate-transformed resources, to obtain ${N_{\mathrm{QL}}}$ sets of single-particle emergent states.
-Factorized measurements operators can be implemented locally on these resources, without calculating the tensor product associated to the final global emergent state.
-VI Conclusions
-In this paper, we expanded upon the general approach derived in Refs.
-Scholes, 2023, 2024; Scholes and Amati, 2024 to define and manipulate QL states.
-In ??
-we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs.
-We introduced the Hilbert space for an arbitrary number of QL bits, and we derived explicitly the Bell states in dimension two.
-In ??
-we proposed a notion of quantum gates applicable to QL states, and we provided concrete examples of their action.
-We discussed in ??
-a list of measurement axioms which allows to observe QL interference effects in our approach.
-Finally, we outlined in ??
-an efficient end-to-end workflow for building QL algorithms based on a controlled approximation for two-QL-bit gates.
+offers an overview of experimental platforms that could be used to implement and validate QL theory.
+Finally, in ??, we discussed that a QL notion of interference can emerge within our framework by presenting a set of axioms for interpreting QL states and observables.
 Several open questions remain, presenting diverse opportunities for future research.
-In the paper we introduced notions of entanglement, interference, and QL probabilities which avoid the expensive calculation of joint probability distributions in the measurement process.
-In future work we plan to investigate how each of these features contributes to achieving a “QL advantage” over strictly classical information processing.
-The ultimate test of the approach proposed in this work will be to implement the pipeline proposed in ??
-(and related ??) to study QL algorithms, and evaluate their time efficiency and resource demand.
-Bharti et al.
-(2022)
 It can be insightful to directly relate our work to the rigorous approach to quantum non-Markovianity proposed in Refs.
@@
 (2022); Amati et al.
-(2022); Mannouch and Richardson (2023) It can be insightful to assess whether even more accurate quasiclassical approaches can be derived from a QL stochastic theory which approximates ??
+(2022); Mannouch and Richardson (2023); Amati (2024) It can be insightful to assess whether even more accurate quasiclassical approaches can be derived from a QL stochastic theory which approximates ??
 beyond the first classical term.
-In this paper, we introduced a universal set of gates comprising single QL-bit terms and the CNOT operation.
-Our controlled approximation is designed to preserve the Cartesian product structure by selectively neglecting a small contribution within the CNOT gate that would otherwise lead to symmetry breaking.
-Preserving this contribution would result in an exponential increase in complexity, posing significant challenges for both computational and experimental resources.
-An open question remains as to whether directly implementing an approximation of many-QL-bit gates, rather than decomposing them into one- and two-body elementary terms, could yield even greater accuracy.
-While any local decomposition is exact in standard unitary quantum computing, this is in general not the case within our QL approach.
-Additionally, alternative choices of universal sets may provide greater computational advantages.
-In future work, we aim to explore the impact of these different scenarios within our framework.
-From an experimental standpoint, various possibilities can be considered for implementing synchronizing resources in QL information processing.
-The research on oscillator-based computing, as reviewed in Ref.
-Csaba and Porod, 2020, offers a solid foundation for this purpose.
-Within this framework, insulator-to-metal-transition devices and spin-torque oscillators emerge as promising candidates, supporting distributed and parallel computing while ensuring resilience to noise and fault tolerance.Raychowdhury et al.
-(2019) We plan to further investigate platforms for the experimental implementation of QL resources, focusing on the properties needed by various approaches to achieve optimal network design and control.
-In recent years, memcomputing has emerged as an innovative computational framework that exploits the intrinsic time-nonlocal properties of classical dynamical systems to enable a terminal-agnostic framework to implement non-Boolean logic.
-Di Ventra and Traversa (2018) This approach, which allows the computation to occur in memory rather than through sequential logic operations, could be exploited by recasting classical synchronizing dynamics leading to emergent states into time-nonlocal equations of motion.
-Amati, Meyer, and Schilling (2019) The classical Nakajima–Mori–Zwanzig formalism, which captures the influence of past states on open classical systems, offers a strategy to incorporate memory effects directly into the QL formalism, potentially enhancing stability and coherence.
-The environmental implications of the large-scale industrial production of quantum processors are topics of ongoing debate.
-Ezratty (2023); Ikonen, Salmilehto, and Möttönen (2017); Paler and Basmadjian (2022) For instance, superconducting devices require significant energy due to the cryogenic cooling of the constituent superconducting circuits.
-Similarly, quantum computers based on arrays of Rydberg atoms demand intensive energy resources.
-Meyer (2000); Bernien et al.
-(2017) Implementing energy-efficient classical circuits for QL computing can contribute to sustainable quantum-processing engineering.
-Additionally, our method leverages classical synchronization, which scales more efficiently with resources compared to its quantum-mechanical counterpart.
-Aifer, Thingna, and Deffner (2024)
+In this paper, we presented an early-stage analysis exploring the connection between QL gates and the many-body synchronization dynamics of oscillator systems.
+Specifically, we identified relationships between the structure of the adjacency matrix of the network and the synchronization patterns observed in the Kuramoto model.
+For future work, we aim to develop a more formal framework by deriving general expressions for multi-time correlation functions associated with network transformations induced by chains of QL gates.
+Additionally, we plan to investigate whether the structure of emergent QL states can be directly inferred from the synchronizing steady states and how this process is influenced by the microscopic details of the underlying classical dynamics.
+Future work could explore new possibilities regarding the generality of the present approach for different synchronizing dynamical systems.
+While this study focused on the nonlinear interactions of the Kuramoto model, it is crucial to assess the applicability of the results to a broader class of classical systems exhibiting synchronization.
+Extending this framework to more general models, where oscillators are represented by vectorial variables de Aguiar (2023) or elements of a unitary group, as in the Lohe model Lohe (2009); Scholes (2022), offers a promising path for further research.
+These generalizations enable the encoding of high-dimensional information within the phase relationships between graph nodes, potentially enhancing the modeling of complex dynamical systems.
+In recent years, memcomputing has emerged as a new computational paradigm that leverages the intrinsic time-nonlocal properties of nonlinear classical dynamical systems to enable non-Boolean logic computations by exploiting memory effects within physical systems.
+A key feature of memcomputing is its terminal-agnostic nature, where input and output terminals are treated equivalently.
+This allows self-organizing circuits to dynamically adjust and converge to equilibrium points that correspond to solutions, regardless of the terminal through which data is introduced or extracted.
+Di Ventra and Traversa (2018) The feasibility of the memcomputing framework has been demonstrated across various physical implementations, including oscillator-based systems.
+Traversa et al.
+(2015); Di Ventra (2022) It can be insightful to implement a variant of our approach with flavors of memcomputing techniques, by exploiting time-nonlocality dissipative nonlinear dynamical systems, e.g.
+as predicted within the Nakajima–Mori–Zwanzig formalism.
+Nakajima (1958); Mori (1965); Zwanzig (1960); Dominy and Venturi (2017); Amati et al.
+(2022); Amati, Meyer, and Schilling (2019)
+The environmental implications of large-scale industrial production of quantum processors remain a topic of ongoing debate.
+Ezratty (2023); Ikonen, Salmilehto, and Möttönen (2017); Paler and Basmadjian (2022); Sood and Chauhan (2024); Jaschke and Montangero (2023); Auffèves (2022) Conducting an early-stage assessment of the environmental impact of various resources can provide valuable metrics to identify the most energy-efficient and scalable quantum technologies for future development.
+In this context, implementing energy-efficient classical circuits for QL computing offers a promising pathway toward sustainable QL-processing engineering.
+While our perspective does not claim to provide a definitive solution, we recognize that classical systems, despite their larger size, could serve as viable energy-efficient alternatives for specific use cases, such as small-scale, proof-of-concept analyses in fundamental research.
 Finally, it can be a fascinating topic for future research to connect QL information processing introduced here to other information theories on human decision-making and psychology derived in the framework of QL contextual probability models.
@@
 (2024) can support theoretical and experimental investigations in this framework.
-VII Acknowledgments
-This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
-Appendix A Transformation of the emergent states to the computational basis
+VIII Acknowledgments
+We thank Debadrita Saha for useful discussions. This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
+Appendix A General transformation from emergent states to the computational basis
 In ??
@@
 $\psi$ $=\sum_{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}F^{\psi}_{\bm{\sigma}^{\prime}}u_{\bm{\sigma}^{\prime},\bm{\sigma}}g_{\bm{\sigma}},$ (B.2a)
-$u_{\bm{\sigma}^{\prime},\bm{\sigma}}$ $=\braket{g_{\bm{\sigma}},f_{\bm{\sigma}^{\prime}}}.$ (B.2b)
+$u_{\bm{\sigma}^{\prime},\bm{\sigma}}$ $=\braket{g_{\bm{\sigma}},f_{\bm{\sigma}^{\prime}}},$ (B.2b)
 which leads to the following mapping between the coefficients of $\psi$ in the two bases:
@@
 This process allows us to define a complete set of mutually orthogonal eigenvectors.
-Appendix D Controlled-NOT gate
-In this appendix, we analyze the action of the CNOT gate on a pair of QL bits, $p,q\in{1,\dots,{N_{\mathrm{QL}}}}$.
-Here, we designate $p$ as the control QL bit, whose state determines whether to flip the state of the target QL bit $q$.
-For simplicity, we assume $p<q$; the case $p>q$ can be derived analogously.
-The unitary map implementing the transformation is
-$U_{\mathrm{CNOT}}^{(p,q)}$ $=\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\uparrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}$
-$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\downarrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes U_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)},$ (D.1)
-where we defined the projection operators
-$\mathcal{P}_{i}^{(p)}=\tfrac{1}{2}\left[U_{\mathcal{I}}^{(p)}+(-1)^{i}U_{\mathrm{Z}}^{(p)}\right],$ (D.2)
-with $p\in\{1,\cdots,{N_{\mathrm{QL}}}\}$ and $i\in\{0,1\}$.
-These are mutually orthogonal [$\mathcal{P}_{\sigma}^{(p)}\mathcal{P}_{\sigma^{\prime}}^{(p)}=\delta_{\sigma\sigma^{\prime}}\mathcal{P}_{\sigma^{\prime}}^{(p)}$] and complementary [$\mathcal{P}_{\downarrow}^{(p)}+\mathcal{P}_{\uparrow}^{(p)}=U_{\mathcal{I}}^{(p)}$].
-The action of ??
-on the total resource ??
-is
-$\mathcal{R}_{\mathrm{CNOT}(q,p)}^{[{N_{\mathrm{QL}}}]}$ $=\sum_{\sigma\in\{\uparrow,\downarrow\}}\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{R}_{\sigma\sigma}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}+\sum_{\begin{subarray}{c}r=1\\ r\neq q\end{subarray}}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}^{(r)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}$ (D.3a)
-$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\uparrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes\mathcal{R}^{(q)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-q)}+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\downarrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes\mathcal{R}_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)}$ (D.3b)
-$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\left[\mathcal{R}_{\uparrow\downarrow}^{(p)}+\mathcal{R}_{\downarrow\uparrow}^{(p)}\right]\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes U_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)},$ (D.3c)
-where $\mathcal{R}^{(p)}_{\sigma\pi}=\mathcal{P}^{(p)}_{\sigma}\mathcal{R}^{(p)}\left[\mathcal{P}^{(p)}_{\pi}\right]^{\dagger}$.
-As introduced in ??, we can minimize the computational complexity of this gate by assessing to which extent the structure of ??
-can be approximated by a Cartesian product.
-This property enables information through transformations applied exclusively to individual one-QL-bit resources.
-Let us therefore study separately all terms which would break the Cartesian product construction.
-Starting from the two terms in ??, we can study the relative error
-$\frac{\left\lVert\mathcal{R}_{\mathrm{X}}^{(q)}-\mathcal{R}^{(q)}\right\rVert}{\left\lVert\mathcal{R}^{(q)}\right\rVert}=\sqrt{\frac{8l{N_{G}}}{2k{N_{G}}+2l{N_{G}}}}=\frac{2}{\sqrt{1+\frac{k}{2l}}}\simeq 1+\frac{2l}{k}.$ (D.4)
-In ?? we introduced the Frobenius norm, defined by $\left\lVert A\right\rVert^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}|A_{ij}|^{2}$, for each $A\in M_{m\times n}(\mathbb{C})$. Also, we noticed that
-$\mathcal{R}^{(q)}_{\mathrm{X}}=U_{\mathrm{X}}^{[q]}\mathcal{R}^{(q)}\left[U_{\mathrm{X}}^{[q]}\right]^{\dagger}=\begin{pmatrix}A_{1}^{(q)}&-C^{(q)}\\ -\left[C^{(q)}\right]^{\top}&A_{2}^{(q)}\end{pmatrix}.$ (D.5)
-?? is minimized by the parameter choice defined in ??. Finally, let us consider the cross contribution in ??. This terms involves the difference
-$\mathcal{R}^{(p)}_{\uparrow\downarrow}+\mathcal{R}^{(p)}_{\downarrow\uparrow}$ $=\frac{1}{2}(\mathcal{R}^{(p)}-\mathcal{R}^{(p)}_{Z})$
-$=\frac{1}{2}\begin{pmatrix}A_{1}^{(p)}-A_{2}^{(p)}&C^{(p)}-\left[C^{(p)}\right]^{\top}\\ \left[C^{(p)}\right]^{\top}-C^{(p)}&A_{2}^{(p)}-A_{1}^{(p)}\end{pmatrix},$ (D.6)
-where
-$\mathcal{R}^{(q)}_{\mathrm{Z}}=\begin{pmatrix}A_{2}^{(q)}&\left[C^{(q)}\right]^{\top}\\ C^{(q)}&A_{1}^{(q)}\end{pmatrix}$ (D.7)
-is calculated similarly to ??. Without loss of generality it is possible to choose here
-$A_{1}^{(p)}$ $=A_{2}^{(p)},$ (D.8a)
-$C^{(p)}$ $=\left[C^{(p)}\right]^{\top},\hskip 14.22636pt\forall\;p=1,\cdots,{N_{\mathrm{QL}}}.$ (D.8b)
-In essence, we obtain
-$\mathcal{R}_{\mathrm{CNOT}(q,p)}^{[{N_{\mathrm{QL}}}]}$ $=\sum_{\sigma\in\mathcal{S}}\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{R}_{\sigma\sigma}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}$
-$\quad+\sum_{\begin{subarray}{c}r=1\\ r\neq q\end{subarray}}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}^{(r)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}$
-$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}_{\mathrm{X}}^{(q)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}+\mathcal{O}(l/k)$ (D.9)
-provided the conditions in ?? hold.
+Appendix D Dynamics of the Kuramoto model
+Here, we present a closed-form solution for the dynamics of the Kuramoto model, as described by the nonlinear equations of motion in ??.
+Our approach follows the methods outlined in Refs.
+Muller, Mináč, and Nguyen, 2021; Kawamura et al., 2010.
+This analysis offers a practical framework for implementing the dynamics of the model efficiently, without the need for explicitly integrating the equations of motion, making it applicable to a broad range of synchronizing networks.
+Additionally, this work can serve as a starting point for future analytical studies exploring the role of emergent states in synchronization processes.
+If the coupling matrix $\mathcal{R}$ is real, it is convenient to subtract from ?? an additional complex component,
+$\dot{\theta}_{i}$ $=\omega_{i}+\frac{\Gamma}{2{N_{G}}}\sum_{j=1}^{2{N_{G}}}\mathcal{R}_{ij}\left[\sin(\theta_{j}-\theta_{i})-i\cos(\theta_{j}-\theta_{i})\right],$ (D.1)
+leading to
+$i\dot{\theta}_{i}=i\omega_{i}+\frac{\Gamma}{2{N_{G}}}\,\mathrm{e}^{-i\theta_{i}}\sum_{j=1}^{N_{\mathrm{QL}}}\mathcal{R}_{ij}\,\mathrm{e}^{i\theta_{j}}.$ (D.2)
+?? can be expressed in matrix form as
+$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathrm{e}^{i\bm{\theta}}=\left[{\mathrm{diag}}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}\right]\,\mathrm{e}^{i\bm{\theta}}$ (D.3)
+or, more compactly as, $\dot{\bm{x}}=\tilde{\mathcal{R}}\bm{x}$, where $\bm{x}=\,\mathrm{e}^{i\bm{\theta}}=(\,\mathrm{e}^{i\theta_{1}},\cdots,\,\mathrm{e}^{i\theta_{2{N_{G}}}})^{\top}$ and ~$\tilde{\mathcal{R}}=\mathrm{diag}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}$.
+This allows us to obtain as a closed form for the solution of the dynamics of the Kuramoto model
+$\bm{x}(t)=\,\mathrm{e}^{\tilde{\mathcal{R}}t}\bm{x}(0).$ (D.4)
+We can decompose the real and imaginary part of the frequency vector $\bm{\theta}=\bm{\theta}_{\mathrm{re}}+i\bm{\theta}_{\mathrm{im}}$, to obtain as a physical solution of ??
+$i\bm{\theta}_{\mathrm{re}}(t)-\bm{\theta}_{\mathrm{im}}(t)$ $=\log\left[\,\mathrm{e}^{\tilde{\mathcal{R}}t}\bm{x}(0)\right],$ (D.5)
+or
+$\bm{\theta}_{\mathrm{re}}(t)$ $=\mathrm{Im}\left\{\log\left[\,\mathrm{e}^{\tilde{\mathcal{R}}t}\bm{x}(0)\right]\right\},$ (D.6a)
+$\bm{\theta}_{\mathrm{im}}(t)$ $=-\mathrm{Re}\left\{\log\left[\,\mathrm{e}^{\tilde{\mathcal{R}}t}\bm{x}(0)\right]\right\}.$ (D.6b)

### v6

@@
 I Introduction
-The present noisy-intermediate scale quantum (NISQ) era of quantum devices is witnessing the successful implementation and test of diverse quantum algorithms on medium-size quantum hardware, from search on unsorted databases, Zhang et al.
+The present noisy-intermediate scale quantum era of quantum devices is witnessing the successful implementation and test of diverse quantum algorithms on medium-size quantum hardware, from search on unsorted databases, Zhang et al.
 (2021) to variational quantum algorithms, Cerezo et al.
@@
 For example, we discussed in Ref.
-Scholes, 2025a how this principle can support the engineering of molecular delocalization.
+Scholes, 2025 how this principle can support the engineering of molecular delocalization.
 The same reference, along with Refs.
@@
 A spectral analysis of this class of graphs highlights the presence of two isolated emergent states.
-These an can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??).
+These can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??).
 The robustness of emergent states against noise appears to be a direct consequence of synchronization.
@@
 Instead, to observe nonclassical interference it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and projective measurements.
+This paper can be read alongside other recent developments by the authors within the same framework, as presented in the complementary work in Ref.
+Amati and Scholes, 2025.
+In that study, we show that arbitrary quantum states can be encoded into the synchronizing steady state of a nonlinear oscillator system.
+We also apply the approach proposed here to construct a QL analog of entanglement.
+Although the two works are closely related, each is self-contained and can be read independently.
 The present paper is structured as follows.
@@
 In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models.
-Khrennikov (2010) Finally, in ??, we propose a variety of experimental resources that could be utilized to implement and verify experimentally our formalism.
+Khrennikov (2010) Finally, in ??, we propose a variety of experimental resources that could be utilized to experimentally implement and verify our formalism.
 In conjunction with Refs.
@@
 The present analysis provides a complete description of the axioms of states and operators in quantum mechanics, within the notation of the emergent states.
-Later in the paper we show the direct connection between this abstract approach and the dynamics of many-body classical systems.
+Later in the paper we show the direct connection between this abstract approach and the dynamics of classical nonlinear oscillators.
 II.1 Isolated system
-Let us consider two classical many-body systems with pairwise interactions between the elementary constituents.
-The pair-wise interactions can be mapped onto two graphs $G_{1}$ and $G_{2}$.
+Let us consider two classical many-body systems with pairwise interactions between the elementary constituents mapped onto two graphs, $G_{1}$ and $G_{2}$.
 Here, we assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices.
@@
 $\sum_{i=1}^{N_{G}}[A_{n}]_{ij^{\prime}}=\sum_{j=1}^{N_{G}}[A_{n}]_{i^{\prime}j}=k,\hskip 22.76219pt\forall i^{\prime},j^{\prime}=1,\ldots,{N_{G}},$ (II.1)
-and $n=1,2$.
-A key property of these subgraphs is that they exhibit an emergent state—a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs.
-Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) The direct sum of the two adjacency matrices is
+and $n=1,2$. The direct sum of the two adjacency matrices is
 $A_{1}\oplus A_{2}=\begin{bmatrix}A_{1}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&A_{2}\end{bmatrix},$ (II.2)
@@
 Conversely, ??, being overall symmetric and loop-free, describes a simple undirected graph.
-We sample $C$ by randomly assigning $1$ to $N_{\mathrm{C}}\leq{N_{G}}^{2}$ of its entries.
-This matrix is chosen to be regular with valency $l=N_{\mathrm{C}}/{N_{G}}$ [similarly to ??].
+The matrix $C$ is also chosen to be regular, with valency $l$.
 ??
@@
 In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states.
-By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information processing that can be controlled by network manipulation.
+The occurrence of emergent states is a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs.
+Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) By mapping the emergent states onto a Hilbert space, we obtain a platform for information processing that can be controlled by suitable network processing.
 We will elaborate on these important concepts further in the manuscript.
-We study the spectral properties of the resources by plotting in ?? the distribution $\rho(\lambda)$ of the eigenvalues of ?? for increasing values of the valency $l$ of $C$.
-Figure 3: Main panel: distribution of the eigenvalues $\lambda$ of the QL-resource ??, for increasing values of the order $l$ of the coupling matrix $C$, from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light).
+We study the spectral properties of the graph resources by plotting in ??
+the distribution $\rho(\lambda)$ of the eigenvalues of ??
+for increasing values of the valency $l$ of the regular off-diagonal block $C$.
+Figure 3: Main panel: distribution of the eigenvalues $\lambda$ of the QL resource ??, for increasing values of the order $l$ of the coupling matrix $C$, from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light).
 Inset: largest and second largest eigenvalues of the spectrum as a function of $l$, with the same color code as in the main panel.
 We fix here ${N_{G}}=100$ and $k=25$.
-Results are shown from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light).
-The other parameters are fixed to ${N_{G}}=100$ and $k=25$.
-The emergent states can be identified for each $l$ from the isolated sharp peaks in the spectrum.
-For $l=0$ we observe a single isolated line associated to the two-fold degenerate eigenvalue $\lambda=k=25$ (red markers).
+More specifically, we sample a set of graphs with the structure given in ??.
+Given that each block of $\mathcal{R}$ defines a regular random graph, the only constraints given in the sampling are that each of the vertices in the $A_{i}$’s and $C$ is connected to $k$ and $l$ other vertices, respectively.
+In ??
+we show a family of distributions for a fixed value of $k=25$ and increasing vaues of $l$, from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light).
+Here, the sizes of all blocks are fixed to ${N_{G}}=100$.
+The emergent states can be identified for each $l$ from the couples of isolated sharp peaks (with identical colors), displaced symmetrically around $\lambda=k=25$.
+More specifically, for $l=0$ we observe a single isolated line associated to the two-fold degenerate eigenvalue $\lambda=k=25$ (red markers).
 As the valency $l$ increases, this emergent eigenvalue split into two, and the peaks shift to $\lambda_{\pm}=k\pm l$.
@@
 where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$.
-As both the $A_{i}$’s and $C$ are exactly regular, then $a_{i}=(1,,1\ldots,1)^{\top}/\sqrt{2{N_{G}}}$ are vectors with constant weights.
-We keep, however, the notation general, to account for possible structural disorder in the network leading to this break of symmetry.
+As both the $A_{i}$’s and $C$ are exactly regular, then $a_{i}=(1,1\ldots,1)^{\top}/\sqrt{2{N_{G}}}$ are vectors with constant weights.
+Here and in the following we decided, however, to keep a more general notation, to account for possible structural disorder in the network that could break this symmetry.
 The vectors $\psi_{\sigma}$, $\sigma\in\{\downarrow,\uparrow\}\equiv\mathcal{S}$, are eigenvectors of ??
 with eigenvalues $\{k-l,k+l\}$, respectively.
-Note that in prior work we used the intention of the QL bit construction to represent explicitly the basis of the the subgraphs, so that the emergent state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and similarly $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2025) ??
-describes another basis choice, equally valid.
+In earlier work, we used the QL bit construction to explicitly associate the subgraph basis with the emergent states: for example, the state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2025) ??
+presents a different, but equally valid, choice of basis.
 The present approach establishes a direct connection between the computational basis and the emergent states of the composite graph described by ??.
-In the following we show that the mapping in ??
-leads to a well-defined Hilbert space.
-First of all, it is straightforward to show that the $\psi_{\sigma}$’s are mutually orthogonal and normalized:
-$\braket{\psi_{\sigma},\psi_{\sigma^{\prime}}}$ $=\sum_{i=1}^{2{N_{G}}}\psi_{\sigma,i}^{*}\psi_{\sigma^{\prime},i}=\frac{1}{2}\left(|a_{1}|^{2}+(2\delta_{\sigma\sigma^{\prime}}-1)|a_{2}|^{2}\right)$
-$=\delta_{\sigma\sigma^{\prime}}\mapsto\braket{{\sigma}{\sigma^{\prime}}}{,}$ (II.5)
-where $\sigma,\sigma^{\prime}\in\mathcal{S}$. Operators acting on ?? are expressed in terms of the matrices in $M_{2{N_{G}}}(\mathbb{R})$
-$\Psi_{\sigma\sigma^{\prime}}=\psi_{\sigma}\otimes\psi_{\sigma^{\prime}}^{\dagger}\mapsto\ket{\sigma}\bra{\sigma^{\prime}}.$ | . (II.6)
-As it would be expected in an abstract Hilbert space, the maps in ?? combine according to the rule
-$\Psi_{\sigma\sigma^{\prime}}\Psi_{\pi\pi^{\prime}}=\Psi_{\sigma\pi^{\prime}}\braket{\psi_{\sigma^{\prime}},\psi_{\pi}},$ (II.7)
-and they transform under Hermitian conjugation as $\Psi_{\sigma\sigma^{\prime}}^{\dagger}=\Psi_{\sigma^{\prime}\sigma}$.
 In this section we summarized a list of prescriptions for defining a mapping from a bipartite graph to the Hilbert space of a single qubit, and related properties.
@@
 II.2.1 Formalism
-Various representations of the quantum states are possible with respect to classical network structures.
+Various representations of quantum states are possible with respect to classical network structures.
 For a small number of QL bits the representation defined in Ref.
 Scholes, 2024 is useful.
-In that work the Bell states, involving two QL bits, are designed to give insight on proof-of-principle examples of entanglement.
+In that work the Bell states, involving two QL bits, are designed to give insight on proof-of-principle examples of QL entanglement.
 In Ref.
@@
 the resource for a system of ${N_{\mathrm{QL}}}$ bits
-$\mathcal{R}=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\mathbb{1}_{2{N_{G}}}\right]\otimes\mathcal{R}^{(q)}\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}\right],$ (II.8)
+$\mathcal{R}=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\mathbb{1}_{2{N_{G}}}\right]\otimes\mathcal{R}^{(q)}\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}\right],$ (II.5)
 where each $\mathcal{R}^{(q)}$ has the structure shown in ??. ?? is the adjacency matrix of the Cartesian product graph Scholes and Amati (2025)
@@
 Here, $\mathcal{G}^{(q)}$ is the single-QL-bit graph with adjacency matrix $\mathcal{R}^{(q)}$. The tensors
-$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}},$ (II.10)
+$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}}$ (II.7)
 are eigenvectors of ??.
-Here, different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\cdots,\sigma^{(N_{\mathrm{QL}})}\right\}\in\bm{\mathcal{S}}$, where $\bm{\mathcal{S}}=\mathcal{S}^{N_{\mathrm{QL}}}$.
-The components of ??
-are expressed via the multi-index $\bm{i}=\left\{i^{(1)},\cdots,i^{(N_{\mathrm{QL}})}\right\}$ as
-$[\psi_{\bm{\sigma}}]_{\bm{i}}$ $=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}.$ (II.11)
+Different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\ldots,\sigma^{(N_{\mathrm{QL}})}\right\}\in\bm{\mathcal{S}}$, where $\bm{\mathcal{S}}=\mathcal{S}^{N_{\mathrm{QL}}}$.
 As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\mathcal{S}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$.
@@
 $\mathcal{R}\psi_{\bm{\sigma}}$ $=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\psi_{\sigma^{(p)}}\right]\otimes\left[\mathcal{R}^{(q)}\psi_{\sigma^{(q)}}\right]\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(p^{\prime})}}\right]$
-$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.12)
+$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.8)
 where
-$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.13)
+$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.9)
 ????
@@
 Scholes and Amati (2025)
-In analogy to ?? for ${N_{\mathrm{QL}}}=1$, the states in ?? define an orthonormal basis for $\mathcal{H}$ according to the scalar product defined by
+The states in ?? define an orthonormal basis for $\mathcal{H}$ according to the scalar product defined by
 $\braket{\psi_{\bm{\sigma}},\psi_{\bm{\sigma}^{\prime}}}$ $=\frac{1}{2^{N_{\mathrm{QL}}}}\prod_{q=1}^{{N_{\mathrm{QL}}}}\left[|a_{1}^{(q)}|^{2}+(2\delta_{\sigma^{(q)},\sigma^{\prime(q)}}-1)|a_{2}^{(q)}|^{2}\right]$
-$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.14)
-Similarly, we extend the mapping of operators in ?? to
-$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.15)
-with components
-$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}\right]_{\bm{i,j}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}\psi^{*}_{\sigma^{\prime(q)},j^{(q)}}.$ (II.16)
-The Hermitian conjugate of ?? is defined such that
-$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.17)
-and we generalize ?? to
-$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.18)
-The mapping discussed here provides a general description of state vectors, entanglement and operators constructed from the emergent states of an arbitrary number of QL bits.
+$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.10)
+Elementary operators in our approach are defined by
+$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.11)
+The Hermitian conjugates of ?? are defined such that
+$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.12)
+and they obey the orthogonality relations
+$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.13)
+Let us note that the Cartesian product of ${N_{\mathrm{QL}}}$ QL-bits, each defined by a graph with $2{N_{G}}$ vertices, yields a graph whose adjacency matrix has dimension $(2{N_{G}})^{N_{\mathrm{QL}}}$, scaling exponentially with the number of QL-bits.
+This behavior is, to some extent, an expected trade-off when encoding quantum states into classical many-body systems.
+As discussed in ??, this issue presents a practical—though not fundamental—limitation to testing the formalism with available computational and experimental resources.
+The mapping proposed in this section provides a general description of state vectors, entanglement and operators constructed from the emergent states of an arbitrary number of QL bits.
 In the next ??
@@
 Here, we study the spectral properties of two-QL-bit resources, and we show how to construct entangled Bell states with them. ?? in this case simplifies to
-$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.19)
+$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.14)
 The spectrum $\rho(\lambda)$ of a resource with structure as in ??
@@
 via the convolution
-$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.20)
+$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.15)
 where $\rho^{(q)}(x)$ is the spectrum of a single QL bit, shown in ??.
@@
 To obtain the explicit expression of the eigenvectors, it is convenient to label the eigenvectors of each QL bit as
-$\psi^{(1)}_{\uparrow/\downarrow}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\uparrow/\downarrow}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.21)
+$\psi^{(1)}_{\uparrow/\downarrow}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\uparrow/\downarrow}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.16)
 where $a_{i}$’s and $b_{i}$’s are defined as in ??
@@
 The notation in ??
-indicate that the upper arrow ($\uparrow$) on the left-hand side is associated to the positive sign on right-hand side, and similarly for $\downarrow$ and the negative sign.
+indicate that the arrow on the left (right) of the slash is associated to the upper (lower) sign on the right-hands side.
 With ??
 we can express
-$\psi_{\uparrow,\uparrow/\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \pm a_{2}\otimes b_{2}\end{pmatrix},\hskip 14.22636pt\psi_{\downarrow,\uparrow\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}\\ \mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.22a)
+$\psi_{\uparrow,\uparrow/\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \pm a_{2}\otimes b_{2}\end{pmatrix},\hskip 14.22636pt\psi_{\downarrow,\uparrow/\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}\\ \mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.17a)
 We can finally construct maximally entangled Bell states via the linear combinations
-$\Phi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.23a)
-$\Psi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.23b)
+$\Phi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.18a)
+$\Psi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.18b)
 that is,
-$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\\ \mathbb{0}_{{N_{G}}}\\ a_{2}\otimes b_{2}\end{pmatrix},$ $\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{{N_{G}}}\\ a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\end{pmatrix},$ (II.24a)
-$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\\ \mathbb{0}_{{N_{G}}}\\ -a_{2}\otimes b_{2}\end{pmatrix},$ $\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{{N_{G}}}\\ -a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\end{pmatrix}.$ (II.24b)
+$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ 0_{N_{\mathrm{G}}^{2}}\\ 0_{N_{\mathrm{G}}^{2}}\\ a_{2}\otimes b_{2}\end{pmatrix},$ $\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0_{N_{\mathrm{G}}^{2}}\\ a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ 0_{N_{\mathrm{G}}^{2}}\end{pmatrix},$ (II.19a)
+$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ 0_{N_{\mathrm{G}}^{2}}\\ 0_{N_{\mathrm{G}}^{2}}\\ -a_{2}\otimes b_{2}\end{pmatrix},$ $\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0_{N_{\mathrm{G}}^{2}}\\ -a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ 0_{N_{\mathrm{G}}^{2}}\end{pmatrix},$ (II.19b)
+where
+$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (II.20)
 So far, we have presented general results on the QL formalism and demonstrated that the approach enables an exact mapping to the Hilbert space of a qubit system.
@@
 Additionally, the dynamics should guarantee that synchronization leads to a globally stable state that remains robust over long time scales.
-These features are satisfied by the non-symplectic dynamics of the Kuramoto model Kuramoto (1975); Kawamura et al.
+These features are satisfied by the Kuramoto model, Kuramoto (1975); Kawamura et al.
 (2010); Scholes (2022); Rodrigues et al.
-(2016), which is characterized by its nonlinear equations of motion
-$\dot{\theta}_{i}=\omega_{i}+\frac{\Gamma}{2{N_{G}}}\sum_{j=1}^{2{N_{G}}}\mathcal{R}_{ij}\sin(\theta_{j}-\theta_{i}),\hskip 19.91692pti=1,\cdots,2{N_{G}}.$ (III.1)
+(2016) described by the nonlinear and non-symplectic equations of motion
+$\dot{\theta}_{i}=\omega_{i}+\frac{\Gamma}{2{N_{G}}}\sum_{j=1}^{2{N_{G}}}\mathcal{R}_{ij}\sin(\theta_{j}-\theta_{i}),\hskip 19.91692pti=1,\ldots,2{N_{G}}.$ (III.1)
 Here, the $\theta_{i}$’s are angular variables corresponding to $2{N_{G}}$ oscillators, defined on the vertices of a graph.
@@
 Despite this strong perturbation, a collective synchronization state persists at long times, with only minor phase shifts affecting small subgroups of oscillators.
-This observation further supports the authors’ previous findings on the resilience of synchronization dynamics to noise–a property particularly appealing for robust information encoding and processing in the steady states of these classical dynamical systems.
+This observation further supports the authors’ previous findings on the resilience of synchronization dynamics to noise—a property particularly appealing for robust information encoding and processing in the steady states of these classical dynamical systems.
 Scholes and Amati (2025); Scholes (2023, 2024)
 To examine the role of the emergent states in the synchronization dynamics it is convenient to express the adjacency matrix through its spectral decomposition, as
-$\mathcal{R}=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\lambda_{\bm{\sigma}}\left(\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}}^{\dagger}\right)+\sum_{i=1}^{M-2}\tilde{\lambda}_{i}\left(\tilde{\psi}_{i}\otimes\tilde{\psi}_{i}^{\dagger}\right),$ (III.2)
-where $M=(2{N_{G}})^{N_{\mathrm{QL}}}$ denotes the total number of states.
+$\mathcal{R}=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}}\psi_{\bm{\sigma}}^{\dagger}+\sum_{i=1}^{M}\tilde{\lambda}_{i}\tilde{\psi}_{i}\tilde{\psi}_{i}^{\dagger},$ (III.2)
+where $M=(2{N_{G}})^{N_{\mathrm{QL}}}-2^{N_{\mathrm{QL}}}$.
 The first term in ??
@@
 The second term corresponds to the random part of the spectrum, with eigenvalues $\tilde{\lambda}_{i}$’s and corresponding eigenvectors $\tilde{\psi}_{i}$’s.
-Panel b) of ??
-depicts the time evolution of the Kuramoto dynamics when only the random part of the spectrum is taken into account [that is, by imposing $\lambda_{\bm{\sigma}}\equiv 0$ $\forall\bm{\sigma}\in\bm{\mathcal{S}}$ in ??].
+In panel b) of ??, we show the time evolution of a system initialized with the same conditions as in panel a) and governed by the same interaction rules, with the only difference being the addition of static random noise $n$.
+This noise has zero mean, variance $\sigma_{\mathrm{n}}=3$, and is sampled independently for each nonzero component of the adjacency matrix.
+Despite the noise being strong compared to the oscillator coupling strength (set to 1), the system still shows a clear synchronization pattern over time.
+This provides numerical evidence that, under broad conditions, the dynamics of the model can be highly resilient to static disorder.
+Such robustness is related to the collective behavior that emerges through synchronization.
+Finally, we show in panel c) of ??
+the time evolution of the Kuramoto dynamics when only the random part of the spectrum is taken into account [that is, by imposing $\lambda_{\bm{\sigma}}\equiv 0$ $\forall\bm{\sigma}\in\bm{\mathcal{S}}$ in ??].
 In this scenario, synchronization fails to emerge, with all degrees of freedom maintaining their initial random phases throughout the entire time evolution.
@@
 Figure 5: a): Time evolution of a system with $602{N_{G}}=$ 60 oscillators governed by the Kuramoto model [??].
-The angular variables $\theta_{i}$ are initialized uniformly in $[0,2\pi)$, and the intrinsic frequencies $\omega_{i}$ are sampled uniformly from $(0,1)$.
+The angular variables $\theta_{i}$ are initialized uniformly in $[0,2\pi)$, and the intrinsic frequencies $\omega_{i}$ are sampled uniformly from $[0,1]$.
 The coupling constant is set to $\Gamma=10$, and the connectivity matrix $\mathcal{R}$ is structured as a regular graph, as described in ??, with the valencies of the diagonal and off-diagonal blocks set to $k=29$ and $l=15$, respectively.
 b): Time evolution of the system under the same parameters and initial conditions as in a), but with disorder introduced into the connectivity matrix $\mathcal{R}$.
-Specifically, each nonzero entry of $\mathcal{R}$ is perturbed by random white noise with zero mean and variance $\sigma=3$.
+Specifically, each nonzero entry of $\mathcal{R}$ is perturbed by random white noise with zero mean and variance $\sigma_{\mathrm{n}}=3$.
 c): Dynamics of the Kuramoto model with the connectivity matrix restricted to its random components [see discussion following ??].
@@
 To connect the usual definition of gates acting on an abstract Hilbert space to our QL formalism, it is convenient to introduce a map $U_{\mathrm{cb}}$ transforming the emergent eigenvectors onto the standard computational basis.
-As that all elementary subrgaphs in our resources [????] are regular, this map can be expressed analytically by
+Given that all elementary subrgaphs in our resources [????] are regular, this map can be expressed analytically by
 $U_{\mathrm{cb}}=[V_{\mathrm{H}}\otimes\mathbb{1}_{N_{G}}]^{\otimes{N_{\mathrm{QL}}}},$ (IV.1)
@@
 denotes the matrix representation of the Hadamard (H) gate in the computational basis. The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$, where
-$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (IV.3)
-and
-$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (IV.4)
+$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix}.$ (IV.3)
 Note that, given that all the basis subgraphs are regular, the solution to ??
@@
 We show in ??
-that a solution for the map $U_{\mathrm{cb}}$ can be determined if the resources are not exactly regular, by solving numerically a linear algebra problem.
+that a solution for the map $U_{\mathrm{cb}}$ can be determined even if the resources are not exactly regular, by solving numerically a linear algebra problem.
 We define the QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ by the matrix transformation
-$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (IV.5)
+$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (IV.4)
 To elucidate the action of ??, let us consider a few examples. The QL maps of the identity operator ($\mathcal{I}$) and the three Pauli gates (X,Y,Z) are
-$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.6a)
-$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.6b)
-$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.6c)
-$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.6d)
-respectively. As expected, the transformations in ?? act as the standard Pauli gates:
+$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.5a)
+$U_{\mathrm{X}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.5b)
+$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.5c)
+$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.5d)
+respectively. As expected, the transformations in ?? act as the standard Pauli gates, namely,
+$U_{\mathrm{X}}\psi_{\uparrow/\downarrow}=\psi_{\downarrow/\uparrow},\hskip 14.22636ptU_{\mathrm{Y}}\psi_{\uparrow/\downarrow}=\pm i\psi_{\downarrow/\uparrow},\hskip 14.22636ptU_{\mathrm{Z}}\psi_{\uparrow/\downarrow}=\pm\psi_{\uparrow/\downarrow}.$ (IV.6)
+As another example, the QL map corresponding to the Hadamard gate transforms states in the factorized basis into superposition states:
+$U_{\mathrm{H}}\psi_{\uparrow/\downarrow}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\uparrow/\downarrow}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (IV.7)
+The approach proposed here applies equally to any kind of gate, including those outside the Clifford group.
+This class of gates break the stabilizer structure, making them harder to simulate compared to Clifford gates.
+In contrast, Clifford gates can be simulated efficiently on a classical computer.
+However, this difference between Clifford and non-Clifford gates does not matter in our framework.
+Our formalism already takes into account the exponential cost of scaling QL resources, no matter what type of gate is used.
+The action of the non-Clifford T-gate on the factorized basis is
+$U_{\mathrm{T}}\psi_{\sigma}=\left(\delta_{\sigma,\uparrow}+\,\mathrm{e}^{i\pi/4}\delta_{\sigma,\downarrow}\right)\psi_{\sigma}.$ (IV.8)
 […]
-As another example, the QL map corresponding to the Hadamard gate transforms states in the factorized basis into superposition states:
-$U_{\mathrm{H}}\psi_{\uparrow/\downarrow}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\uparrow/\downarrow}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (IV.8)
-The formal approach proposed here applies equally to any arbitrary gates, including those outside the Clifford group.
-While gates outside the Clifford group break the stabilizer structure and require exponentially more resources compared to Clifford gates—which can be efficiently simulated classically—the distinction between these two classes of gates is not relevant to our framework.
-Our formalism is designed to be computationally exact, addressing the exponentially high cost associated with scaling QL resources, regardless of the type of gate involved.
-The action of the non-Clifford T-gate on the factorized basis is
-$U_{\mathrm{T}}\psi_{\uparrow/\downarrow}=\frac{1}{2}\left[(1+\,\mathrm{e}^{i\pi/4})\psi_{\uparrow}+(1-\,\mathrm{e}^{i\pi/4})\psi_{\downarrow}\right].$ (IV.9)
-[…]
-$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}}$ (IV.10)
+$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}}$ (IV.9)
 transforms the elements of the factorized basis as follows:
-$U_{\mathrm{CN}}\psi_{\uparrow[\uparrow,\downarrow]}=\psi_{\uparrow[\uparrow,\downarrow]},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow[\uparrow,\downarrow]}=\psi_{\downarrow[\downarrow,\uparrow]}.$ (IV.11)
+$U_{\mathrm{CN}}\psi_{\uparrow,\uparrow/\downarrow}=\psi_{\uparrow,\uparrow/\downarrow},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow,\uparrow/\downarrow}=\psi_{\downarrow,\downarrow/\uparrow}.$ (IV.10)
 Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis.
-We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??.
+We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$ from the linearity of the map ??.
 From a practical standpoint, it is preferable to translate these state operations onto transformations applicable directed to the graph resources.
@@
 A straightforward approach in this regard is to define the unitary map
-$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}.$ (IV.12)
+$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}.$ (IV.11)
 ??
 transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??.
-This property can be easily inferred by expanding $\mathcal{R}$ on the right-hand side of ??
+This can be easily inferred by expanding $\mathcal{R}$ on the right-hand side of ??
 via the spectral decomposition in ??.
+This map can be straightforwardly extended to arbitrary quantum circuits, described by the sequence of $M$ gate operations $\bm{\mathrm{g}}=\{\mathrm{g}_{1},\ldots,\mathrm{g}_{M}\}$.
+In this case, we define by
+$U_{\bm{\mathrm{g}}}=U_{\mathrm{cb}}^{-1}(V_{\bm{\mathrm{g}}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}=U_{\bm{\mathrm{g}}}=U_{\mathrm{g}_{M}}\cdots U_{\mathrm{g}_{1}},$ (IV.12)
+the corresponding unitary map acting on QL states, with an analog network transformation defined by
+$\mathcal{R}_{\bm{\mathrm{g}}}=U_{\bm{\mathrm{g}}}\mathcal{R}U_{\bm{\mathrm{g}}}^{\dagger}.$ (IV.13)
 In the next ??
 we provide concrete examples of ??, and we discuss how this network transformation affects the steady-state synchronizing solution of a system of oscillators.
+A more general analysis of circuit transformations, associated to ??, can be found in Ref.
+Amati and Scholes, 2025.
 IV.2 Connection to classical synchronization
@@
 The Pauli-X (NOT) gate transforms an initial resource ?? as follows:
-$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{X}}=\begin{pmatrix}A_{1}&-C\\ -C^{\top}&A_{2}\end{pmatrix}.$ (IV.13)
-We show in panel a) of ??, the synchronization pattern in the dynamics of the Kuramoto model with correlation matrix $\mathcal{R}_{\mathrm{X}}$.
+$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{X}}=\begin{pmatrix}A_{1}&-C\\ -C^{\top}&A_{2}\end{pmatrix}.$ (IV.14)
+We show in panel a) of ??
+the synchronization pattern in the dynamics of the Kuramoto model with correlation matrix $\mathcal{R}_{\mathrm{X}}$.
 In the transformation ??, the leading emergent eigenvector, associated to the largest eigenvalue $k+l$, is mapped from $\psi_{\uparrow}$ to $\psi_{\downarrow}$ [see ??].
@@
 In the case of the Hadamard gate, resources are transformed according to
-$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{H}}=\begin{pmatrix}A_{1}+A_{2}+C+C^{\top}&A_{1}-A_{2}-C+C^{\top}\\ A_{1}-A_{2}+C-C^{\top}&A_{1}+A_{2}-C-C^{\top}\end{pmatrix}.$ (IV.14)
-The sum and difference of the basis graphs $A_{1}$ and $A_{2}$ in ??
-highlight the fundamental role of this gate in creating a superposition state.
+$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{H}}=\begin{pmatrix}A_{1}+A_{2}+C+C^{\top}&A_{1}-A_{2}-C+C^{\top}\\ A_{1}-A_{2}+C-C^{\top}&A_{1}+A_{2}-C-C^{\top}\end{pmatrix}.$ (IV.15)
 The structure of ??
-encodes key information about the expected synchronization patterns in the dynamics of the corresponding Kuramoto model, as depicted in panel b) of ??.
-Dephasing emerges between the first and second halves of the oscillators, primarily driven by cancellation effects in the two off-diagonal blocks, where the term $A_{1}-A_{2}$ plays a dominant role.
-When $A_{1}\simeq A_{2}$ and $C\simeq C^{\top}$, the off-diagonal blocks approach zero, leading to completely uncorrelated phases between the two oscillator subgroups over extended timescales.
-Furthermore, the second half of the oscillators, labeled by indices $i\in\{{N_{G}},\cdots 2{N_{G}}\}$ exhibits weaker synchronization compared to the first one.
+encodes information about the expected synchronization patterns in the dynamics of the corresponding Kuramoto model, as depicted in panel b) of ??.
+Dephasing emerges between the first and second halves of the oscillators, primarily driven by cancellation effects in the two off-diagonal blocks.
+When $A_{1}\approx A_{2}$ and $C\approx C^{\top}$, these blocks approach zero, decorrelating the phases between the two oscillator subgroups.
+Furthermore, the second half of the oscillators, labeled by indices $i\in\{{N_{G}},\ldots,2{N_{G}}\}$ exhibits weaker synchronization compared to the first half.
 This is due to the fact that the matrix norm of the lower diagonal block in ??
@@
 Consequently, the second group of oscillators exhibits weaker correlations compared to the first group.
+The interested reader is referred to Ref.
+Amati and Scholes, 2025.
+There, we show with a straightforward generalization of the present Kuramoto model that this kind classical synchronization patterns can accurately represent arbitrary quantum states.
 In this section, we derived general principles for mapping network resources in a manner consistent with an arbitrary set of QL gates.
 Additionally, we presented numerical evidence supporting the connection between QL gates and the steady-state synchronization of classical oscillator resources.
-In the following ??, we explore specific examples of experimental resources that could be utilized to implement our formalism in practice.
+In the following ??, we explore specific examples of experimental resources that could be utilized to implement this approach in practice.
 V Experimental perspectives
@@
 We discussed the primary role of emergent states in synchronization dynamics and the robustness of the latter in the presence of noise.
+Additionally, we designed a general approach to implement gate transformations on the network resources.
 In this section, we explore perspectives on implementing the proposed formalism using experimental physical resources.
@@
 (2021); Ali et al.
-(2024) In our approach, error correction in the presence of noise translates to enhancing the robustness of long-time classical synchronization [see the discussion on panel b) of ??].
-This ensures distinct and readily measurable transformations of the steady-state dynamics.
-Optimal synchronization can be achieved by numerically analyzing phase diagrams of synchronization indicators.
-For instance, in the case of the Kuramoto model, time-resolved order parameters offer clear insight into the synchronization strength.
-Basnarkov and Urumov (2007); Wang, Xu, and Zheng (2021) Parameter domains with the highest degree of synchronization can be identified as optimal candidates for experimental implementation.
+(2024) In our approach, error correction in the presence of noise could be achieved by encoding information into a larger number of oscillators.
+Given that those are defined on a classical phase space, the related cost of the strategy scales polynomially with the number of resources.
+This suggests a simple strategy to implement a QL version of repetition codes in our approach.
 In terms of resource complexity, the number of degrees of freedom required to model the synchronization dynamics of a system of ${N_{\mathrm{QL}}}$ QL bits scales exponentially with system size.
@@
 Additionally, we identify the classes of algorithms that can be most effectively analyzed by those resources.
-Implementing coupled LC circuits is a straightforward approach to implement macroscopic experimental resources that exhibit classical synchronization.
+Implementing coupled LC circuits is a straightforward approach to build macroscopic resources that exhibit classical synchronization.
 Allam et al.
@@
 Maffezzoni et al., 2015 show that this class of systems can be effectively approximated by Kuramoto-type models for medium-scale systems of 60 oscillators.
-The couplings between the model oscillators are tuned by adjusting the transconductance values of the circuit transistors.
+The couplings between the model oscillators are tuned by adjusting the transconductance of the differential pair transistors.
 Spin-torque oscillators (STO) emerge as another promising candidate for implementing QL information processing.
@@
 They have the ability to synchronize in both frequency and phase with external oscillatory signals as well as with other STOs.
-The full magnetization dynamics of these systems, defined by the Landau-Lifshitz-Gilbert-Slonzewski (LLGS) equation, can be simplified to the Kuramoto model for small-amplitude oscillations.
+The full magnetization dynamics of these systems, defined by the Landau-Lifshitz-Gilbert-Slonzewski equation, can be simplified to the Kuramoto model for small-amplitude oscillations.
 Belanovsky et al.
@@
 (2019) Ref.
-Flovik, Macià, and Wahlström, 2016 demonstrates the experimental feasibility of modeling synchronizing classical systems with up to a few hundred oscillators.
+Flovik, Macià, and Wahlström, 2016 demonstrates the experimental feasibility of modeling synchronizing classical systems with up to a few hundred STOs.
 Beyond this scale, technical constraints arising from finite-size effects restrict the extent of achievable synchronization.
@@
 (2019) One limitation of current MIT devices is their low oscillation frequency, typically below 1 MHz, which hinders the ability to achieve fine-resolution synchronization and detect subtle variations in coupled oscillator behavior.
-However, ongoing advancements in scaling down these devices are expected to increase their oscillation frequencies, enabling more precise and time-resolved synchronization measurements.
+However, ongoing advancements in scaling down these devices are expected to increase their oscillation frequencies, enabling more accurate time-resolved synchronization measurements.
 Cotter et al.
 (2014)
-Building on the experimental perspectives outlined above, it appears feasible to validate our QL-bit theory using networks of up to a hundred oscillators.
-Such an experimental setup would enable the implementation of proof-of-concept algorithms on systems with ${N_{\mathrm{QL}}}=2$ or ${N_{\mathrm{QL}}}=3$ QL-bits, where each QL-bit comprises a small network of 555510101010 oscillators.
-A particularly intriguing class of quantum information processes for systems of this scale includes those that generate entanglement, such as the formation of Bell states (for ${N_{\mathrm{QL}}}=2$) and their higher-order generalizations, the Greenberger–Horne–Zeilinger (GHZ) states.
-Greenberger, Horne, and Zeilinger (1989); Erhard, Krenn, and Zeilinger (2020) This analysis has the potential to provide valuable insights into fundamental questions surrounding the classical-to-quantum correspondence.
-In particular, it could help clarify the role of non-emergent states in classical synchronizing systems as potential hidden variables within a QL-based framework for analyzing entanglement.
-Scholes (2025b)
-We conclude this section by presenting a practical perspective on implementing an experimental scheme to test our approach:
+Building on the perspectives discussed above, experimental validation of our theory using networks of up to a hundred oscillators appears feasible.
+In such a setup, each QL bit could consist of a small network of 5–10 oscillators, allowing for proof-of-concept implementations with ${N_{\mathrm{QL}}}=2$ or ${N_{\mathrm{QL}}}=3$ QL bits.
+At this scale, a particularly interesting class of quantum information processes includes entanglement generation, such as the formation of Bell states (for ${N_{\mathrm{QL}}}=2$) and their higher-order generalizations, the Greenberger–Horne–Zeilinger (GHZ) states.
+Greenberger, Horne, and Zeilinger (1989); Erhard, Krenn, and Zeilinger (2020) Advancements in this framework are given in Ref.
+Amati and Scholes, 2025.
+Exploring small-scale quantum processes might offer a possible route to address some foundational questions.
+For example: To what extent can one construct a classical analogue of entanglement in a classical many-body system?
+Is there a fundamental limit beyond which even infinite classical resources fail to replicate essential features of entanglement?
+Such a limitation could signal an intrinsic boundary between which kind of quantum behaviors can be reproduced by classical dynamics.
+We argue that our approach can serve as a platform to investigate these general problems.
+We end this section with a practical outlook on how our approach could be tested experimentally.
   1. 1.
-    Engineer a classical synchronizing system: Design a system with two-body correlations represented by a graph, ensuring that its leading emergent state corresponds to the desired initial state of a QL algorithm.
-For example, the factorized state $\psi_{\uparrow}$ can be associated with a Cartesian-product network structured as described in ??.
+    Engineer a classical synchronizing system: Design a classical system of nonlinear oscillators with two-body interactions, defined by a graph.
+This graph should be constructed so that its leading emergent state matches the desired initial state of a quantum circuit.
+For example, the factorized state $\psi_{\uparrow}$ corresponds to the eigenvector with the largest eigenvalue of a Cartesian-product network, as given in ??.
+The other parameters of the classical system should be tuned to ensure the emergence of synchronization at long times.
   2. 2.
-    Initialize oscillator phases: Assign arbitrary initial conditions to the oscillator phases, ensuring they adhere to the constraint that the system’s long-time steady state achieves the expected synchronization.
+    Transform the network according to arbitrary quantum logic: Apply to the network a unitary transformation corresponding to a given quantum circuit, following the protocol proposed in ??.
   3. 3.
-    Allow the classical dynamics to equilibrate: Let the system evolve until it reaches a steady state.
-This step requires time-resolved measurements to ensure the system has sufficient time to reach a long-time synchronized state.
+    Allow the classical dynamics to equilibrate: Assign random initial phases to the classical oscillators and let the system evolve until it reaches a synchronized steady state.
+We discuss formally in Ref.
+Amati and Scholes, 2025 how this steady states encodes information on the gate initially applied to the classical network.
   4. 4.
-    Perform gate operations: Each gate operation consists of two phases: (a) transforming the correlations between the degrees of freedom and (b) allowing the system to synchronize and relax to a new emergent state.
-  5. 5.
-    Measure the final state: Apply projection operators to extract the state of the final resource after all gate operations have been implemented.
-The proposed scheme can be tested numerically by measuring multi-time correlation functions.
-This process involves a series of time evolution operators defined by gate-transformed resources at different stages of a given QL algorithm.
-Preliminary analytical work on the Kuramoto dynamics is discussed in ??
-[see ??].
+    Measure the final state: Calculate the overlap between the oscillators’ steady state and the eigenstates of arbitrary observable operators.
+The resulting projection coefficients can then be interpreted as nonclassical probability amplitudes within the framework of QL theories, as discussed in detail in the next ??.
+In the next and final ??
+we expand over this last point, to show that our approach allows to mimic non-Kolmogorov statistics, that violates the Bayes theorem due to the occurrence of nonclassical interference.
 Figure 6: a): Time evolution of a Kuramoto model where the adjacency matrix is transformed according to the mapping in ??
@@
 $\mathcal{F}$ $=\sum_{{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}}\left(\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}+\gamma^{*}_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right),\hskip 14.22636pt\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\in\mathbb{C},$ (VI.2)
-and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$ [??]. The QL measurement of $\mathcal{F}$ on a state $\psi$ is defined by the map $\psi\mapsto\mathcal{F}\psi$.
+and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$ [??].
 Axiom 3.
@@
 $P_{\psi}\left(\lambda_{\bm{\sigma}}^{\mathcal{F}}\right)=\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}=|\braket{f_{\bm{\sigma}},\psi}|^{2},$ (VI.3)
-the QL probability of collapsing onto the subspace spanned by the eigenvector $f_{\bm{\sigma}}$ with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$.
+the QL probability of collapsing onto the subspace spanned by the normalized eigenvector $f_{\bm{\sigma}}$ with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$.
 ??
 is expressed in terms of the projector
-$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}\otimes f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (VI.4)
+$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (VI.4)
 Axiom 5.
 QL conditional probabilities are defined via consecutive projections onto the eigenstates of physical observables.
-In particular, let us consider two generic QL observable operators $\mathcal{F}$ and $\mathcal{G}$ whose spectra are defined by
-$\mathcal{F}f_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}},\hskip 28.45274pt\mathcal{G}g_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}.$ (VI.5)
-For a nonzero state $\psi$ we denote by
+In particular, let us consider two generic QL observable operators $\mathcal{F}$ and $\mathcal{G}$ whose spectra decompositions are defined by
+$\mathcal{F}=\sum_{\bm{\sigma}\in{\bm{\mathcal{S}}}}\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}}f_{\bm{\sigma}}^{\dagger},\hskip 28.45274pt\mathcal{G}=\sum_{\bm{\sigma}\in{\bm{\mathcal{S}}}}\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}g_{\bm{\sigma}}^{\dagger},$ (VI.5)
+respectively. For a nonzero state $\psi$ we denote by
 $P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)=\frac{\left\lVert\Pi^{\mathcal{G}}_{\bm{\sigma}^{\prime}}\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}{\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}$ (VI.6)
@@
 We show in the next ??
-that the QL probability model defined via Axioms 1 to 5 violates by construction Kolmogorovian classical statistics.
+that the stochastic model defined via Axioms 1 to 5 violates by construction Kolmogorovian classical statistics.
 This feature allows to introduce a notion of QL interference in our approach.
@@
 VI.2 Quantum probabilities in a QL model
-Before introducing a formula for the update of QL probabilities, it is instructive to briefly recall its classical correspondent, to assess analogies and differences between the two.
+Before introducing a formula for the update of QL probabilities, it is useful to briefly recall its classical correspondent, to assess analogies and differences between the two.
 A classical probability space is defined by the triplet $\{\Omega,\Sigma,P^{(\mathrm{cl})}\}$, where $\Omega$ is a space of experimental outcomes, $\Sigma$ the $\sigma$-algebra of its subsets, and $P^{(\mathrm{cl})}$ a probability measure on $\Sigma$.Shiryaev and Wilson (2013) Be $\{\mathcal{A}_{j}\}_{j=1}^{J}$ a partition of $\Omega$, i.e.
-a collection of $J$ sets defined such that $A_{j}\neq\emptyset\;\forall j$, $\mathcal{A}_{j}\cap A_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ and $\cup_{j=1}^{J}\mathcal{A}_{j}=\Omega$.
+a collection of $J$ sets defined such that $\mathcal{A}_{j}\neq\emptyset\;\forall j$, $\mathcal{A}_{j}\cap A_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ and $\cup_{j=1}^{J}\mathcal{A}_{j}=\Omega$.
 $P^{(\mathrm{cl})}$ is a positive measure on the partition elements, that is $P^{(\mathrm{cl})}(\mathcal{A}_{j})>0\;\forall j$, and it is normalized: $\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{A}_{j})=1$.
@@
 for the present QL framework.
-We restrict the following analysis to couples of non-degenerate QL observables, $\mathcal{F}$ and $\mathcal{G}$, whose eigenvalue problems are defined as in ??.
-The eigenvalues of these operators are associated with a one-dimensional subspaces, which can be fixed by selecting any normalized eigenvector.
+We restrict the following analysis to couples of non-degenerate observables, $\mathcal{F}$ and $\mathcal{G}$, whose eigenvalue problems are defined as in ??.
+The eigenvalues of these operators are associated with one-dimensional subspaces, which can be fixed by selecting any eigenvector.
 In this case, the QL conditional probabilities defined in ??
@@
 is provided in ??, where we give the explicit expression of the nonclassical interference coefficient $\Lambda^{\psi}_{\bm{\sigma},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}}$ [??].
-Information processing can be designed in a QL way, to enhance or suppress some spectral measurements compared to classical computing protocols.
-Khrennikov (2021)
-As noted in Ref.
-Khrennikov, 2021, a significant feature of the QL probability formula ??
-is that it bypasses the need for calculating joint probability distributions, which are typically exponential in computational complexity.
-In contrast, such calculations are required for the classical updates outlined in ??.
-This distinction relates directly to the foundational principles of quantum advantage.
-We intend to further investigate this property in future testing of QL algorithms.
+We conclude this section by connecting the abstract discussion above to an operational strategy to simulate interference phenomena.
+In Ref.
+Amati and Scholes, 2025 we prove that the total phase vector of a system of classical oscillators can be expanded over the sum of all eigenvalues of the graph, including both emergent contributions (isolated peaks) and non-emergent terms.
+However, only the emergent states are associated to a quantum Hilbert space.
+To resolve the quantum information encoded in the system it is therefore required to maximize the spectral gap between emergent and random states.
+This is achieved via an ad-hoc choice of the valencies of the elementary graphs comprising the network, as discussed in ??.
+Conveniently, this preparation stage needs to be defined once for all for any circuit, as the graph spectra are preserved by unitary QL gates.
+Let consider as an example the specific case of LC circuits, already introduced in ??.
+It is shown in Ref.
+Maffezzoni et al., 2015 that the oscillatory patterns of these circuits allow to build Kuramoto-type models, where the interactions between oscillators are defined from transconductance elements.
+After building this network and allowing for its long-time synchronization, a readout of the phases of the oscillators phases is required.
+As the voltage $V_{j}(t)$ of each LC oscillator is related to its angle via $V_{j}(t)\propto\cos\left[\theta_{j}(t)\right]$, classical phase readout in this setup can be implemented by means of voltage probes, such as oscilloscopes or lock-in amplifiers.
+Finally, QL measurement processes can be simulated by numerically postprocessing the classical phase vector.
+For example, conditional probablities from ??
+can be calculated by projecting the phases of the oscillators onto the spectrum of a QL observable operator $\mathcal{F}$.
+These nonclassical probabilities obey by construction the five axioms listed above, mimicking nonclassical interference phenomena.
 VII Conclusions
@@
 ??
-offers an overview of experimental platforms that could be used to implement and validate QL theory.
-Finally, in ??, we discussed that a QL notion of interference can emerge within our framework by presenting a set of axioms for interpreting QL states and observables.
+offers an overview of experimental platforms that could be used to implement and test this theory.
+Finally, in ??, we discussed that a QL notion of interference can emerge within this framework by presenting a set of axioms for interpreting QL states and observables.
 Several open questions remain, presenting diverse opportunities for future research.
+As firstly discussed in ??, the number of edges in the Cartesian product graph defined in ??, along with the associated gate transformations in ??, increases exponentially with the number of QL bits.
+This exponential scaling imposes practical constraints on the system sizes that can be investigated with current computational and experimental capabilities.
+Nevertheless, we believe that moderately sized networks are sufficient to demonstrate the core principles of QL algorithms.
+For example, entangled states such as Bell and GHZ states can be effectively generated and analyzed using shallow-depth circuits.
+Scholes and Amati (2025) In terms of scaling, to date high-performance classical simulations allow the implementation of systems with up to approximately ${N_{\mathrm{QL}}}\lesssim 5-10$ QL bits and ${N_{G}}\lesssim 10$.
+Simulations of a Kuramoto model with this system size are given, e.g., in Ref.
+Skardal, 2018.
 It can be insightful to directly relate our work to the rigorous approach to quantum non-Markovianity proposed in Refs.
@@
 beyond the first classical term.
-In this paper, we presented an early-stage analysis exploring the connection between QL gates and the many-body synchronization dynamics of oscillator systems.
-Specifically, we identified relationships between the structure of the adjacency matrix of the network and the synchronization patterns observed in the Kuramoto model.
-For future work, we aim to develop a more formal framework by deriving general expressions for multi-time correlation functions associated with network transformations induced by chains of QL gates.
-Additionally, we plan to investigate whether the structure of emergent QL states can be directly inferred from the synchronizing steady states and how this process is influenced by the microscopic details of the underlying classical dynamics.
-Future work could explore new possibilities regarding the generality of the present approach for different synchronizing dynamical systems.
-While this study focused on the nonlinear interactions of the Kuramoto model, it is crucial to assess the applicability of the results to a broader class of classical systems exhibiting synchronization.
-Extending this framework to more general models, where oscillators are represented by vectorial variables de Aguiar (2023) or elements of a unitary group, as in the Lohe model Lohe (2009); Scholes (2022), offers a promising path for further research.
-These generalizations enable the encoding of high-dimensional information within the phase relationships between graph nodes, potentially enhancing the modeling of complex dynamical systems.
-In recent years, memcomputing has emerged as a new computational paradigm that leverages the intrinsic time-nonlocal properties of nonlinear classical dynamical systems to enable non-Boolean logic computations by exploiting memory effects within physical systems.
-A key feature of memcomputing is its terminal-agnostic nature, where input and output terminals are treated equivalently.
-This allows self-organizing circuits to dynamically adjust and converge to equilibrium points that correspond to solutions, regardless of the terminal through which data is introduced or extracted.
-Di Ventra and Traversa (2018) The feasibility of the memcomputing framework has been demonstrated across various physical implementations, including oscillator-based systems.
-Traversa et al.
-(2015); Di Ventra (2022) It can be insightful to implement a variant of our approach with flavors of memcomputing techniques, by exploiting time-nonlocality dissipative nonlinear dynamical systems, e.g.
-as predicted within the Nakajima–Mori–Zwanzig formalism.
-Nakajima (1958); Mori (1965); Zwanzig (1960); Dominy and Venturi (2017); Amati et al.
-(2022); Amati, Meyer, and Schilling (2019)
-The environmental implications of large-scale industrial production of quantum processors remain a topic of ongoing debate.
-Ezratty (2023); Ikonen, Salmilehto, and Möttönen (2017); Paler and Basmadjian (2022); Sood and Chauhan (2024); Jaschke and Montangero (2023); Auffèves (2022) Conducting an early-stage assessment of the environmental impact of various resources can provide valuable metrics to identify the most energy-efficient and scalable quantum technologies for future development.
-In this context, implementing energy-efficient classical circuits for QL computing offers a promising pathway toward sustainable QL-processing engineering.
-While our perspective does not claim to provide a definitive solution, we recognize that classical systems, despite their larger size, could serve as viable energy-efficient alternatives for specific use cases, such as small-scale, proof-of-concept analyses in fundamental research.
+The scope and applicability of our approach could be broadened by extending the treatment of quantum gates to any unitary transformation, in particular to time evolution operators.
+From this perspective, one direction for future research is to map the dynamics of a unitary quantum subsystem onto a large network of classical oscillators and, for instance, to explore the effects of decoherence arising from interactions with an external dissipative environment.
+Accurately reproducing quantum time evolution through classical dynamics remains a nontrivial challenge.
+It requires to extend our approach such that of the algebra of Hilbert space generators is preserved, at least approximately.
+Runeson and Richardson (2019); Kapral (2015) This strategy becomes particularly promising when coupled with the enforcement of quantum detailed balance at long times—a direction the authors have recently explored in the context of nonadiabatic quantum dynamics.
+Amati, Mannouch, and Richardson (2023); Amati, Runeson, and Richardson (2023)
+The extent to which our framework offers any advantage compared to quantum computing remains an open question, that we believe will be fully explored by testing further approximations of the exact mapping proposed here.
+More specifically, our approach requires an exponentially large number of oscillators to encode the correlations of the Hilbert space of a system of QL bits.
+This high resource demand does not offer per se any advantage in regards to scaling.
+In Appendix A of our new paper, Ref.
+Amati and Scholes, 2025, we outline as a potential strategy to mitigate the high cost of our approach by replacing full graph representations with their lower rank approximations.
+This can lead to a renormalization of both graph resources and corresponding gates, possible by truncating negligible parts of the spectrum.
+In general, we believe that it can be worth challenging the trade-off between implementing genuine quantum resources with polynomial complexity and easier accessible classical resources with unfavorable exponential scaling.
+We view these observations as a starting point for further investigations into the foundational limits and capabilities of our approach and quantum-inspired frameworks in general.
+Future work could explore new possibilities regarding the choice of resources used for information processing.
+For example, our framework could be extended to other flavors of Kuramoto-type models, where oscillators are represented by vectorial variables de Aguiar (2023) or elements of a unitary group.
+Lohe (2009); Scholes (2022) Other classical systems could be used as a resource for information processing in a similar fashion as what we proposed here.
+An example in this regard is the Fermi–Pasta–Ulam–Tsingau chain, where isolated solitary waves, emerging from symplectic classical nonlinear dynamics, could be used to store and process information in a QL fashion.
+Amati, Meyer, and Schilling (2019); Carati et al.
+(2015); Amati and Schilling (2020); Berman and Izrailev (2005) Since the dynamics of this classical model are Hamiltonian, the approach may allow for a well-defined classical detailed balance condition, which could be enforced to match its quantum analog and ensure correct long-time thermalization.
+Amati, Mannouch, and Richardson (2023)
 Finally, it can be a fascinating topic for future research to connect QL information processing introduced here to other information theories on human decision-making and psychology derived in the framework of QL contextual probability models.
@@
 VIII Acknowledgments
-We thank Debadrita Saha for useful discussions. This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
+The authors thank Dr.
+Debadrita Saha for useful discussions.
+This research was funded by the National Science Foundation under Grant No.
+2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
 Appendix A General transformation from emergent states to the computational basis
@@
 $w_{i}^{(q)}a_{1,i}^{(q)}-z_{i}^{(q)}a_{2,i}^{(q)}$ $={N_{G}}^{-1/2},$ (A.2d)
-where $q=1,\cdots,{N_{\mathrm{QL}}}$, $i=1,\cdots,{N_{G}}$. ?? is expressed in matrix form as
+where $q=1,\ldots,{N_{\mathrm{QL}}}$, $i=1,\ldots,{N_{G}}$. ?? is expressed in matrix form as
 $\begin{pmatrix}D_{1}^{(q)}&D_{2}^{(q)}&\mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}&D_{1}^{(q)}&D_{2}^{(q)}\\ D_{1}^{(q)}&-D_{2}^{(q)}&\mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{0}_{N_{G}}&D_{1}^{(q)}&-D_{2}^{(q)}\end{pmatrix}\begin{pmatrix}x^{(q)}\\ y^{(q)}\\ w^{(q)}\\ z^{(q)}\end{pmatrix}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\\ 0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (A.3)
@@
 ??
-can be inverted to determine a solution for $\left(x^{(q)},y^{(q)},w^{(q)},z^{(q)}\right)$ $\forall q=1,\cdots{N_{\mathrm{QL}}}$, corresponding to a unique expression for $U_{\mathrm{cb}}$.
+can be inverted to determine a solution for $\left(x^{(q)},y^{(q)},w^{(q)},z^{(q)}\right)$ $\forall q=1,\ldots{N_{\mathrm{QL}}}$, corresponding to a unique expression for $U_{\mathrm{cb}}$.
 Appendix B Derivation of the quantum-like formula of total probabilities
@@
 Again, $\mathcal{F}$ maps $V_{2}$ onto itself.
-By proceeding this way, we find a sequence $\{\lambda_{n},f_{n},V_{n}\}_{n}$, where the subspace $V_{n}$ is orthogonal to $\mathrm{span}\{f_{1},\cdots,f_{n}\}$.
+By proceeding this way, we find a sequence $\{\lambda_{n},f_{n},V_{n}\}_{n}$, where the subspace $V_{n}$ is orthogonal to $\mathrm{span}\{f_{1},\ldots,f_{n}\}$.
 The sequence ends at the step of order $2^{N_{\mathrm{QL}}}$, given that $\dim(V_{n})=2^{N_{\mathrm{QL}}}-n$.
@@
 $\frac{\mathrm{d}}{\mathrm{d}t}\,\mathrm{e}^{i\bm{\theta}}=\left[{\mathrm{diag}}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}\right]\,\mathrm{e}^{i\bm{\theta}}$ (D.3)
-or, more compactly as, $\dot{\bm{x}}=\tilde{\mathcal{R}}\bm{x}$, where $\bm{x}=\,\mathrm{e}^{i\bm{\theta}}=(\,\mathrm{e}^{i\theta_{1}},\cdots,\,\mathrm{e}^{i\theta_{2{N_{G}}}})^{\top}$ and ~$\tilde{\mathcal{R}}=\mathrm{diag}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}$.
+or, more compactly as, $\dot{\bm{x}}=\tilde{\mathcal{R}}\bm{x}$, where $\bm{x}=\,\mathrm{e}^{i\bm{\theta}}=(\,\mathrm{e}^{i\theta_{1}},\ldots,\,\mathrm{e}^{i\theta_{2{N_{G}}}})^{\top}$ and ~$\tilde{\mathcal{R}}=\mathrm{diag}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}$.
 This allows us to obtain as a closed form for the solution of the dynamics of the Kuramoto model